### v2

# On the composable security of weak coin flipping

Abstract
Weak coin flipping is a cryptographic primitive in which two mutually distrustful parties generate a shared random bit to agree on a winner via remote communication. While a stand-alone secure weak coin flipping protocol can be constructed from noiseless communication channels, its composability has not been explored. In this work, we demonstrate that no weak coin flipping protocol can be abstracted into a simple black box resource with composable security. Despite this, we also establish the overall stand-alone security of weak coin flipping protocols under sequential composition.
1 Introduction
Motivation and background
Cryptographic systems or protocols are expected to guarantee security even when composed with multiple other cryptographic systems. To illustrate this, consider a cryptographic protocol designed to perform secure communication between two parties while leaking at most the length of the communicated message. One possible way to construct such a scheme is to first use a system that generates and distributes secret keys amongst the users and then use another system that pads the key with the intended message to be sent to the receiver (the so-called one-time pad). While it is relatively straightforward to see that a buggy implementation that compromises the security of key distribution or one-time pad would result in overall insecure communication, it is at times difficult to formally argue that the final composition results in a secure communication channel when both the constituent systems (key distribution and one-time pad) are assumed secure in isolation. Indeed, it might seem natural to use small accessible information of the distributed key as the security criterion for key distribution. However, the fact that key distribution using this criterion is secure in isolation does not imply that the final communication protocol is secure [1], due to the data locking effect [2]. We thus say that composability fails for this security criterion.
Several frameworks have been proposed that are intended to analyze compositions of cryptographic systems, starting with Canetti’s seminal Universally Composable (UC) security framework [3, 4]. Since then, several extensions and variants have emerged (see, e.g., [5, 6, 7, 8]). We focus here on the model of Abstract Cryptography (AC) by Maurer and Renner [9, 10]. In contrast to the previous frameworks, the latter is a top-down axiomatic approach and is intended to provide better insights via abstraction of a given complex cryptographic system and will thus be most suitable for our purposes. The security in the AC framework is built on simulation-based security. Here, for a given system to be deemed secure, it suffices to ensure that all adversarial attacks on the system can be simulated on an ideal system. Therefore, to prove the security of any multi-party protocol in the AC framework, one needs to show the existence of such a simulator simulating all possible adversary attacks.
The cryptographic setting we consider in this work is a two-party primitive called weak coin flipping (WCF). Informally speaking, in WCF two mutually distrustful parties, say Alice and Bob, need to agree on a binary outcome even though the two parties have opposite preferences. The protocol is deemed secure if it produces a fair coin toss in case both parties are honest and a dishonest party can at most bias the honest party’s outcome towards their preference. The WCF task is fundamentally different from the other commonly studied tasks such as strong coin flipping (SCF), bit commitment (BC), oblivious transfer, or more generally, secure function evaluation. While the latter protocols are shown to be insecure with an unbounded adversary [11, 12, 13, 14, 15], it is possible to construct weak coin flipping protocols with quantum communication that exhibit almost perfect security with arbitrarily small bias [16]. The explicit protocols which are shown to have an arbitrarily small bias are developed in [17, 18, 19]. The security notion employed in these results, known as information-theoretic security, allows an adversary to enjoy unlimited resources (i.e. the adversary has no restrictions on its computational power or use of memory, etc.). It is well-studied for a range of different cryptographic tasks and, in most cases, is stronger than other commonly studied security definitions such as computational [20, 21, 22], bounded storage [23, 24, 25] or even game-based [26, 27], where an adversary can perform attacks in only some prescribed ways.
Previously, it was shown that WCF enjoys information-theoretic stand-alone security, i.e., it is secure as a single protocol instance in isolation [16, 17, 18, 19]. However, while WCF itself is of limited interest, it is mostly a powerful tool in constructing optimal protocols for several other primitives such as bit commitment [28] and strong coin flipping [29]. However, such constructions would require more than stand-alone security and it is still unclear if WCF protocols have composable security when combined with other cryptographic protocols.
Contributions
Our main contributions encompass two negative theorems and one positive theorem. In Theorem 5, we demonstrate the impossibility of constructing a typical black-box WCF resource analogous to the widely-used SCF resource [30] (see Definition 4) from noiseless communication channels in the AC framework. This WCF resource generates a random bit for both parties, allowing the cheating party to learn the bit in advance and bias it towards their preferred direction. It was shown that the widely-used SCF resource cannot be achieved in the AC framework [31], but the result for the similar WCF resource remains unknown until this work. We state Theorem 5 here informally.
Theorem 1 (Informal).
The typical black-box WCF resource with composable security cannot be constructed from noiseless communication channels in the AC framework.
To prove Theorem 5, we show that WCF protocols using noiseless communication channels are vulnerable to the man-in-the-middle attack when composed in parallel. As a result, the typical black-box WCF resource cannot be constructed by any WCF protocol from noiseless communication channels in the AC framework.
Theorem 5 shows the impossibility of the typical black-box WCF resource in the AC framework. However, we can still define an alternative WCF resource at the cost of increasing the interaction complexity. Nevertheless, such an alternative WCF resource that can be constructed from noiseless communication channels must be extremely complicated in terms of its interaction complexity, as is shown in Theorem 6. Replacing an actual protocol with an ideal resource intends to simplify the analysis; therefore, it is meaningless to study such an alternative WCF resource. Here is the informal statement of Theorem 6.
Theorem 2 (Informal).
Any black-box WCF resource with composable security constructed from noiseless communication channels in the AC framework must have high interaction complexity.
The proof of Theorem 6 consists of two steps. First, we show that an alternative WCF resource with composable security constructed from noiseless communication channels implies a WCF protocol with the same interaction complexity as the alternative WCF resource. Then together with the exponential lower bound on the interaction complexity of WCF protocols, we conclude that a WCF resource with a low interaction complexity can not exist.
Next, we extend the classical version of the sequential composition theorem [32, Theorem 7.4.3] to the quantum case. We find a sufficient condition called global security where WCF and other modules can be composed into a larger system with stand-alone security. A composition between WCF and other components is globally secure if there is no correlation between the outcomes of WCF and other components. We thus show in Theorem 10 that global security holds when WCF and all other components are composed in a strict causal order. That is, there is no other component running concurrently with WCF. The constructions for optimal SCF and optimal BC exactly satisfy global security.
Theorem 3 (Informal).
A primitive constructed from WCF with stand-alone security under strict sequential composition maintains its stand-alone security.
The idea to prove Theorem 10 is by contradiction. Assume that there is some correlation between outcomes of WCF with stand-alone security and other components under sequential composition. Using the information carried from previous components to WCF, we can construct an attack breaking the stand-alone security of WCF. This contradicts our assumption and thus proves Theorem 10.
Related work
Causality in the composable framework. Some early works [33, 34] developed the paradigm that only allows for sequential composition, where protocols are executed one after another. Canetti [3] relaxed “sequential" composition to modular composition in the UC framework, where the former is a special case of the latter. The AC framework [9, 10] takes a top-down hierarchy. Only the connection of abstract systems is considered at the top level, and the execution order is left at the lower level specification. Under the AC framework, the causal box model [35] provides the causal order specification that allows for quantum superposition of different orders.
Concerning concurrent composition. Literature [24, 25] addresses the difficulty of concurrent parallel composition in the UC framework and the bounded storage model. The composability of BC was discussed in [36]. They define the weak binding security of BC in a stand-alone rather than composable manner and propose a counterexample where the string commitment constructed from multiple instances of BC is insecure. In a previous model of general cryptographic protocols [32, Section 7.3], concurrent calls to an oracle (resource) are also forbidden. This restriction primarily stems from the formulation of syntax in that model rather than an explicit security concern.
Coin flipping within the AC framework. To study the composability of SCF, Demay and Maurer [30] propose an SCF resource and prove that it can be constructed from a BC resource. However, this resource is shown to be impossible to construct from noiseless communication channels, even in a relativistic setting [31].
2 Preliminaries
2.1 The abstract cryptography framework
The building blocks of the framework are resources, converters (protocols), and distinguishers. Resources are cryptographic functionalities, and converters transform one resource into another. Distinguishers distinguish resources from each other. The framework is feasible to describe multiparty functionalities, but we only focus on two-party ones here.
  • A resource, denoted by $R$, has one interface accessible to each party. Each party can send inputs, receive outputs, perform operations and give instructions via the corresponding interface. In two-party cryptography, it is necessary to consider three different behaviours $R$, $R_{A}$ and $R_{B}$ of a resource, where $R$, $R_{A}$ and $R_{B}$ correspond to the case where both are honest, where Alice is dishonest, and where Bob is dishonest, respectively. We use $\mathcal{R}=(R,R_{A},R_{B})$ to denote a resource compactly.
  • A converter, denoted by a lower-case Greek letter $\alpha$, is a cryptographic system with an inner interface connected to an interface of a resource and an outer interface to a party. The set of all converters is denoted by $\Sigma$. Each party can instruct the converter on the outer interface to execute specific tasks on the inner interface. A converter $\alpha$ can convert one resource $R$ into another resource $S$ by connecting its inner interface to one interface of the resource $R$. The interfaces of the resource $S$ consist of the unconnected interface of the resource $R$ and the outer interface of the converter $\alpha$. We denote the connection between the converter $\alpha$ and the left-side (right-side) interface of the resource $R$ by $\alpha R$ ($R\alpha$).
  • A protocol $\pi=(\pi_{A},\pi_{B})$ consists of Alice’s converter $\pi_{A}$ and Bob’s converter $\pi_{B}$. $\pi_{A}$ and $\pi_{B}$ are operations that honest Alice and Bob should follow when building one resource from another. By connecting each converter to the corresponding interface, the protocol $\pi$ can build a required resource $S$ from a given resource $R$ if $S=\pi_{A}R\pi_{B}$. For example, suppose $R$ is a binary symmetric channel resource and $(\pi_{A}$,$\pi_{B})$ is an encoder-decoder pair corresponding to a proper error-correcting code. Then one can construct a noiseless channel resource as $S=\pi_{A}R\pi_{B}$.
  • A distinguisher, denoted by $\mathcal{D}$, is an agent who can access both interfaces of a resource. Given one of two resources $R$ and $S$ uniformly at random, a distinguisher guesses whether the resource is $R$ or $S$ by interacting with it. The distinguishing advantage of the above task is defined as
    $d(R,S)=\sup_{\mathcal{D}}|\Pr[\mathcal{D}(R)=0]-\Pr[\mathcal{D}(S)=0]|.$ (1)
    The distinguishing advantage is a pseudo metric between two resources, satisfying the symmetry and triangle inequality properties. When $d(R,S)\leq\delta$, we also write $R\approx_{\delta}S$.
With these building blocks, we define composable security for two-party functionalities.
Definition 1 (Composable security).
A protocol $\pi=(\pi_{A},\pi_{B})$ constructs resource $\mathcal{S}=(S,S_{A},S_{B})$ out of resource $\mathcal{R}=(R,R_{A},R_{B})$ with distinguishing advantage $\delta$ if the following conditions are satisfied
  1. 1.
    Correctness
    $S\approx_{\delta}\pi_{A}R\pi_{B}.$ (2)
  2. 2.
    Security against dishonest Bob
    $\exists\sigma_{B}\in\Sigma,S_{B}\sigma_{B}\approx_{\delta}\pi_{A}R_{B}.$ (3)
  3. 3.
    Security against dishonest Alice
    $\exists\sigma_{A}\in\Sigma,\sigma_{A}S_{A}\approx_{\delta}R_{A}\pi_{B}.$ (4)
This construction is denoted by $\mathcal{R}\overset{\pi,\delta}{\longrightarrow}\mathcal{S}$. Alternatively, we say that $\mathcal{S}$ is an abstraction of $\pi$ based on $\mathcal{R}$, or that $\mathcal{S}$ abstracts $\pi$ based on $\mathcal{R}$.
In the previous definition, $\sigma_{A}$ and $\sigma_{B}$ are called simulators. Intuitively, the simulators here ensure that any valid cheating strategy for resource $\mathcal{R}$ could be mapped onto some cheating strategy for resource $\mathcal{S}$. For instance, if dishonest Bob deviates from the protocol $\pi_{B}$ by following $\pi^{\prime}_{B}$ instead, then, according to Eq. (3), this is equivalent to the action $\sigma_{B}\pi^{\prime}_{B}$ on resource $S_{B}$. For a detailed description of simulators and the related security notions in cryptography, we refer to the tutorial [37].
When $\mathcal{R}$ is a noiseless communication channel, the conditions can be simplified as
$S\approx_{\delta}\pi_{A}\pi_{B},$ (5)
$\exists\sigma_{B}\in\Sigma,S_{B}\sigma_{B}\approx_{\delta}\pi_{A},$ (6)
$\exists\sigma_{A}\in\Sigma,\sigma_{A}S_{A}\approx_{\delta}\pi_{B}.$ (7)
2.2 The quantum comb model
A quantum comb [38, 39] is a general quantum information processing system with internal memory that processes the incoming messages one by one. An example of quantum comb is shown in Fig. 1. It is common to model the building blocks in the AC framework as quantum comb. If a protocol can be modeled with the quantum comb model (QCM), it is called a QCM protocol. Note that all existing WCF protocols can be classified as QCM protocols.
$\mathcal{E}_{1}$$\mathcal{E}_{2}$$\mathcal{E}_{n}$
Figure 1: An example of quantum comb. $\mathcal{E}_{i}$ is a quantum operation. The lines with an arrow are quantum states. The unconnected lines are pairs of input and output messages.
The interaction complexity of a QCM protocol $\pi$, denoted by $\mathsf{Intcom}(\pi)$, is defined as its communication rounds. As an example, for a protocol $\pi$ modeled by Fig. 1, $\mathsf{Intcom}(\pi)=n$. For resources such as $S_{A}$ and $S_{B}$, their interaction complexity refers to the communication rounds at the adversary’s side. For example, the interaction complexity of WCF resource $S_{A}$ in Fig 3(b) is the rounds at the left side, which is $1$.
2.3 The causal box model
The causal box model [35] is a more general specification of building blocks in the AC framework. The causal box model presents the causal structure of the messages sent or received by these building blocks. The quantum comb model can be treated as a special case of the causal box model.
In the causal box model, the abstract blocks are modeled as causal boxes. A causal box is an information processing system with input and output wires. Each message on a wire is described by a quantum state on space $\mathcal{H}\otimes l^{2}(\mathcal{T})$, where $\mathcal{T}$ is a partially ordered set (poset) corresponding to the order of messages and $l^{2}(\mathcal{T})$ is the set of all functions $\mathcal{T}\to\mathbb{C}$ with bounded $L^{2}$ norm. Note that $\mathcal{T}$ cannot be simply interpreted as time, which is totally ordered. For generality, a wire may contain arbitrarily many messages. Therefore, the wire space is defined as the Fock space of the message space
$\mathcal{F}_{\mathcal{H}}^{\mathcal{T}}\coloneqq\bigoplus_{n=0}^{\infty}\vee^{n}\left(\mathcal{H}\otimes l^{2}(\mathcal{T})\right),$ (8)
where $\vee^{n}\mathcal{H}$ denotes the symmetric subspace of $\mathcal{H}^{\otimes n}$ and $l^{2}(\mathcal{T})$ is the space of square summable sequence on $\mathcal{T}$. Sometimes we abuse the notation by writing $\mathcal{F}_{\mathcal{H}_{A}}^{\mathcal{T}}$ as $\mathcal{F}_{A}^{\mathcal{T}}$. With this definition, wires can be merged or split, i.e., for $\mathcal{H}_{A}=\mathcal{H}_{A_{1}}\oplus\mathcal{H}_{A_{2}}$,
$\mathcal{F}_{A}^{\mathcal{T}}=\mathcal{F}_{A_{1}}^{\mathcal{T}}\otimes\mathcal{F}_{A_{2}}^{\mathcal{T}}.$ (9)
A causal box is formally defined as a completely positive and trace-preserving (CPTP) map from the input wire $X$ to the output wire $X^{\prime}$. This map should further respect the causality condition, i.e., the output message at $t^{\prime}\in\mathcal{T}$ only depends on the input messages at $t\in\mathcal{T}$ where $t\preceq t^{\prime}$. A precise definition can be found in Appendix A.
The composition of causal boxes consists of parallel composition and loops. The parallel composition of causal boxes $\Phi$ and $\Psi$, $\Phi\|\Psi$ is defined as the map $\Phi\otimes\Psi$. Due to Equation (9), this map’s input and output space are still valid wires. A loop connects segments with the same dimension of the input and output wires. An interface labelled by $x$ of a resource or a converter $\Phi$, denoted by $\mathsf{int}_{x}(\Phi)$, is specified by a set of wires, possibly including both input and output wires. For example, the inner and outer interfaces of a converter $\alpha$ are denoted by $\mathsf{int}_{in}(\alpha)$ and $\mathsf{int}_{out}(\alpha)$, respectively. Two interfaces $\mathsf{int}_{a}(\Phi),\mathsf{int}_{b}(\Psi)$ are compatible if there exists a bijection $P:\mathsf{int}_{a}(\Phi)\leftrightarrow\mathsf{int}_{b}(\Psi)$ such that each pair of wires consists of an input wire and an output wire with the same dimension.
2.4 Weak coin flipping
Here we formulate WCF protocols with causal boxes.
Definition 2 (Weak coin flipping protocol).
A weak coin flipping protocol using noiseless communication channels is a pair of converters $\pi=(\pi_{A},\pi_{B})$ with compatible inner interfaces. $\pi_{A}$ ($\pi_{B}$) has one output wire on the outer interface and only outputs a single classical bit at $t_{a}$ ($t_{b}$). The poset $\mathcal{T}$ of the protocol is bounded both from above and below, i.e., $\exists t_{0}\in\mathcal{T}$, such that $\forall t\in\mathcal{T}$, it holds that $t_{0}\preceq t$, $t\preceq t_{a}$ and $t\preceq t_{b}$.
$\pi_{B}$$\pi_{A}$($c_{B},t_{b}$)$(c_{A},t_{a})$$|\psi\rangle\in\mathcal{F}_{X}^{\mathcal{T}}$$|\psi^{\prime}\rangle\in\mathcal{F}_{X^{\prime}}^{\mathcal{T}}$$|\phi\rangle\in\mathcal{F}_{Y}^{\mathcal{T}}$$|\phi^{\prime}\rangle\in\mathcal{F}_{Y^{\prime}}^{\mathcal{T}}$
Figure 2: A general model for coin flipping protocol using noiseless communication channels. The message spaces $X,Y$ and $X^{\prime},Y^{\prime}$ are of the same dimension, respectively.
In a WCF protocol, we say Alice wins if $c_{A}=c_{B}=0$ while Bob wins if $c_{A}=c_{B}=1$, which means Alice only wants to prevent Bob from biasing $c_{A}$ towards $1$ and Bob only wants to prevent Alice from biasing $c_{B}$ towards 00.
Notably, in the definition of weak coin flipping, players are not allowed to abort, whereas in strong coin flipping, aborting is required. This distinction arises because, in the weak coin flipping scenario, if a player detects malicious behavior, they can simply declare themselves the winner.
Definition 3 (Stand-alone security of WCF).
Let $\epsilon\leq z$ and $z\in[0,\frac{1}{2}]$. A coin flipping protocol $\pi=(\pi_{A},\pi_{B})$ is a $z$-unbalanced $\epsilon$-biased WCF protocol (with stand-alone security), denoted by $\mathrm{WCF}(z,\epsilon)$, if the following conditions are satisfied:
  1. (S1)
    The composition $\pi_{A}\pi_{B}$ by connecting the inner interface of $(\pi_{A},\pi_{B})$ outputs $(c_{A},c_{B})$ with a probability of
    $\Pr[c_{A}=c_{B}=0]=z,\quad\Pr[c_{A}=c_{B}=1]=1-z.$ (10)
  2. (S2)
    For any converter $\alpha$ compatible with the inner interface with $\pi_{B}$, the composition $\alpha\pi_{B}$ outputs $c_{B}=0$ with a probability of
    $\Pr[c_{B}=0]\leq z+\epsilon.$ (11)
  3. (S3)
    For any converter $\beta$ compatible with inner interface with $\pi_{A}$, the composition $\pi_{A}\beta$ outputs $c_{A}=1$ with a probability of
    $\Pr[c_{A}=1]\leq 1-z+\epsilon.$ (12)
We call a protocol balanced if $z=\frac{1}{2}$ and unbiased if $\epsilon=0$.
3 Weak coin flipping is not universally composably secure
This section shows that WCF cannot be universally composably secure in the AC framework. The result contains two parts. First, we consider a specific WCF resource derived from a simple but widely adapted SCF resource [30] and show that this specific WCF resource cannot be constructed from noiseless communication channel. Nevertheless, this result does not exclude the existence of other types of resources that can be constructed with, say, the protocol in [18]. Then our second result shows that, for any WCF protocol with enough stand-alone security, there does not exist any useful abstraction. Usefulness here means that the resource fully expresses the functionality of WCF and its interaction complexity is low.
3.1 Impossibility for a specific resource
Definition 4 ($z$-unbalanced $\epsilon$-biased WCF resource).
Let $\epsilon\in[0,z]$ and $z\in[0,\frac{1}{2}]$. A $z$-unbalanced $\epsilon$-biased WCF recourse is characterized by a tuple $\mathcal{S}=(S,S_{A},S_{B})$, as is shown in Figure 3.
  1. 1.
    $S$ describes the case when both Alice and Bob are honest. In this case, both party obtains a random bit $c$ satisfying
    $\Pr[c=0]=z,\quad\Pr[c=1]=1-z.$ (13)
  2. 2.
    $S_{A}$ describes the case when Alice is dishonest and Bob is honest. $S_{A}$ outputs a random bit $c^{\prime}$ subject to the distribution $(z,1-z)$ at $t_{0}^{\prime}$ and receives two bits $b^{\prime},p^{\prime}$ at $t_{1}^{\prime}$ at the left interface. Then it outputs a bit $c_{A}$ at the right interface. If $p^{\prime}=1$, then $c_{B}=1$, else the probability distribution of $c_{B}$ follows
    $\Pr[c_{B}=b^{\prime}]=\frac{\epsilon}{1-z},\quad\Pr[c_{B}=c^{\prime}]=1-\frac{\epsilon}{1-z}.$ (14)
    The causality condition is $t_{0}^{\prime}\prec t_{1}^{\prime}\prec t_{b}$.
  3. 3.
    $S_{B}$ works similarly as $S_{A}$. The causality condition is $t_{0}^{\prime\prime}\prec t_{1}^{\prime\prime}\prec t_{a}$.
$S$$(c,t_{a})$$(c,t_{b})$$S_{A}$$(c^{\prime},t_{0}^{\prime})$$(c_{B},t_{b})$$(b^{\prime},t_{1}^{\prime})$$S_{B}$$(c^{\prime\prime},t_{0}^{\prime\prime})$$(c_{A},t_{a})$$(b^{\prime\prime},t_{1}^{\prime\prime})$$(p^{\prime},t_{1}^{\prime})$$(p^{\prime\prime},t_{1}^{\prime\prime})$(a)(b)(c)
Figure 3: (a), (b) and (c) respectively show the components $S,S_{A},S_{B}$ of the WCF resource $\mathcal{S}$. Every message is assigned an element of the order set to indicate the causal order.
The resource in Definition 4 is a reasonable and simple WCF resource. The following theorem indicates that it cannot be constructed from noiseless communication channel.
Theorem 5.
There does not exist a protocol $\pi=(\pi_{A},\pi_{B})$ which can construct an $z$-unbalanced $\epsilon$-biased WCF resource $\mathcal{S}=(S,S_{A},S_{B})$ with a distinguishing advantage $\delta<\frac{1}{3}z(1-z-\epsilon)$.
Proof.
Suppose that there is a protocol $\pi=(\pi_{A},\pi_{B})$ that constructs an ideal $z$-unbalanced $\epsilon$-biased WCF resource $\mathcal{S}=(S,S_{A},S_{B})$ with a distinguishing advantage $\delta$ (without loss of generality we assume $z\in[0,\frac{1}{2}]$). Combining the security conditions for Alice and Bob, respectively, followed by the triangle inequality, we obtain
$\exists\sigma\in\Sigma,S_{B}\sigma S_{A}\approx_{2\delta}\pi_{A}\pi_{B},$ (15)
where $\sigma=\sigma_{B}\sigma_{A}$. Applying the correctness followed again by the triangle inequality, we proceed to
$\exists\sigma\in\Sigma,S_{B}\sigma S_{A}\approx_{3\delta}S.$ (16)
This relation is illustrated in Figure 4.
[…]
Figure 4: The composition of two ideal WCF resources.
We now show that, however, there does not exist $\sigma$ such that $\delta<\frac{1}{12}$. Note that the best choice of causal order for $\sigma$ is $t_{0}^{\prime}\prec t_{1}^{\prime},t_{0}^{\prime}\prec t_{1}^{\prime\prime},t_{0}^{\prime\prime}\prec t_{1}^{\prime},t_{0}^{\prime\prime}\prec t_{1}^{\prime\prime}$ since it can maximize the use of information of $c^{\prime},c^{\prime\prime}$.
The most general operation of $\sigma$ is to apply a probabilistic map $\mathcal{M}_{\sigma}:(c^{\prime},c^{\prime\prime})\mapsto(b^{\prime},b^{\prime\prime},p^{\prime},p^{\prime\prime})$. Because $S$ must output $c_{A}=c_{B}=c$ while $S_{B}\sigma S_{A}$ might output $c_{A}\neq c_{B}$, a distinguisher $\mathcal{D}_{0}$ which only checks if $c_{A}=c_{B}$ to distinguish $S$ and $S_{B}\sigma S_{A}$ will provide a lower bound to the distinguishing advantage. That is, $\mathcal{D}_{0}(\cdot)$ outputs $1$ if $c_{A}=c_{B}$ and else 00. Consider the following map on three cases of the input.
  1. 1.
    $\Pr[c^{\prime}=c^{\prime\prime}=0]=z^{2}$ or $\Pr[c^{\prime}=c^{\prime\prime}=1]=(1-z)^{2}$. Because the bits are already the same, the box $\sigma$ outputs $p^{\prime}=0,b^{\prime}=c^{\prime},p^{\prime\prime}=0,b^{\prime\prime}=c^{\prime\prime}$ to keep $c_{A}$ and $c_{B}$ the same. We have $c_{A}=c_{B}=0$ or $c_{A}=c_{B}=1$.
  2. 2.
    $\Pr[c^{\prime}=0,c^{\prime\prime}=1]=z(1-z)$. The box $\sigma$ outputs $p^{\prime}=1,p^{\prime\prime}=0$ to force the outcome $c_{A}=c_{B}=1$. We have $c_{A}=c_{B}=1$.
  3. 3.
    $\Pr[c^{\prime}=1,c^{\prime\prime}=0]=z(1-z)$. The box $\sigma$ outputs $b^{\prime}=p^{\prime}=b^{\prime\prime}=p^{\prime\prime}=0$ to force the outcome $c_{A}=c_{B}=0$ with probability $\frac{\epsilon}{1-z}$.
This strategy of $\sigma$ is obviously the optimal one to maximize distinguishing probability under distinguisher $\mathcal{D}_{0}$, which is
$\Pr[c_{A}\neq c_{B}]=\Pr[c_{A}=1,c_{B}=0]+\Pr[c_{A}=0,c_{B}=1]=z(1-z-\epsilon).$ (17)
For any box $\sigma$, we have the lower bound on distinguishing advantage:
$d(S_{B}\sigma S_{A},S)\geq\left|\Pr[\mathcal{D}_{0}(S)=1]-\Pr[\mathcal{D}_{0}(S_{B}\sigma S_{A})=1]\right|\geq z(1-z-\epsilon).$ (18)
Together with the relation in Equation (16), we have
$\delta\geq\frac{1}{3}z(1-z-\epsilon).$ (19)
which shows that $\pi$ cannot construct $\mathcal{S}$ with a vanishing distinguishing advantage. ∎
By setting $\epsilon=0,z=\frac{1}{2}$, we conclude that there does not exist a protocol $\pi$ that constructs an ideal balanced unbiased WCF resource $\mathcal{S}$ with a distinguishing advantage $\delta<\frac{1}{12}$.
3.2 Impossibility for a general resource
The composable security is defined by comparing between the ideal resource and the actual protocol. Theorem 5 denies the possibility of constructing an actual protocol close to the ideal WCF resource in Definition 4. However, if we modify the definition of the ideal WCF resource into a presumably more complicated one, then there might still exist some protocol that constructs this more complicated resource.
To address this possibility, it is crucial to determine what properties a WCF resource must possess to be useful in general. For a WCF resource $\mathcal{S}=(S,S_{A},S_{B})$, we propose the following conditions:
  • Fully expressing. A resource $\mathcal{S}$ for $\mathrm{WCF}(z,\epsilon)$ is fully expressing if
    1. 1.
      $S$ outputs the same value $c$ at each side subject to distribution: $\Pr[c=0]=z,\Pr[c=1]=1-z$.
    2. 2.
      For any converter $\alpha$, the composition $\alpha S_{A}$ outputs bit 00 at the right interface with probability at most $z+\epsilon$.
    3. 3.
      For any converter $\beta$, the composition $S_{B}\beta$ outputs bit $1$ at the left interface with probability at most $1-z+\epsilon$.
  • Simple. By simplicity of resources, we typically refer to the process of abstracting away specific details of a resource that reduces its overall interaction complexity and allows to study it in a more general context. Here we take the interaction complexity $\mathsf{Intcom}(S_{A})$, $\mathsf{Intcom}(S_{B})$ (defined in Section 2.2) to quantify it.
Next, we will show that a fully expressing and simple WCF resource cannot be constructed from noiseless communication channels. Specifically, Theorem 6 states that, if a fully expressing resource for $\mathrm{WCF}(z,\epsilon)$ with small $\epsilon$ can be constructed from noiseless communication channels, then its interaction complexity cannot be low.
Theorem 6.
Let $\pi$ be a QCM $\mathrm{WCF}(\frac{1}{2},\epsilon)$ protocol, and $\pi$ constructs a fully expressing resource $\mathcal{S}=(S,S_{A},S_{B})$. Then the interaction complexity of $\mathcal{S}$ satisfies
$\min\{\mathsf{Intcom}(S_{A}),\mathsf{Intcom}(S_{B})\}\geq\exp\left(\Omega(1/\sqrt{\epsilon})\right).$ (20)
To prove Theorem 6, we need a proposition rewritten from [40, Theorem 8.2].
Proposition 7.
If $\pi$ is a QCM $\mathrm{WCF}(\frac{1}{2},\epsilon)$ protocol, then
$\mathsf{Intcom}(\pi)\geq\exp\left(\Omega(1/\sqrt{\epsilon})\right).$ (21)
Proof of Theorem 6.
[…]
$\mathsf{Intcom}(S_{B})=\mathsf{Intcom}(\pi^{\prime})\geq\exp\left(\Omega(1/\sqrt{\epsilon})\right).$ (22)
Some parts of the argument in the above proof are of more general interest. That is, if there exists a resource $\mathcal{S}$ that abstracts a two-party protocol $\pi$, it implies the existence of another protocol $\pi^{\prime}$ with the same interaction complexity as the resource $\mathcal{S}$. Furthermore, if the resource reduces the interaction complexity of $\pi$, then $\pi^{\prime}$ is a better protocol than $\pi$ in terms of interaction complexity. An example is delegated quantum computation (DQC), where the client aims to delegate computation to the server while maintaining privacy. The $S_{A}$ (against dishonest client) of DQC resource [41] abstracting the protocol in [42, 43] has no interaction in either side, because it is just trivial to talk about a dishonest client. This implies the existence of a simple DQC protocol with no interaction, wherein the client performs all the computation herself.
4 Security of weak coin flipping under composition
Section 3 highlights the difficulties when formulating the security of WCF in the AC framework. Fortunately, despite these difficulties, there is a prospect of attaining positive outcomes regarding the composability of such protocols. This section starts with the limitation of the stand-alone security definition (Definition 3) in the context of composition. Specifically, for the construction of unbalanced WCF from a balanced one, we propose a condition (Inequality (26)) which, while implicit, is essential in the security proof of the unbalanced WCF protocol depicted in [29]. Subsequently, we introduce global security as an extension of the stand-alone definition to encompass possible interdependencies between the outputs of the WCF protocol and other components in a larger system. Furthermore, we establish the condition under which global security is maintained.
4.1 Stand-alone security is not sufficient
For a cryptographic task, some security definitions seem fine from a stand-alone viewpoint but may present some loopholes when composed with other protocols. For example, the accessible information criteria for quantum key distribution falls short of providing composable security due to the data-locking effect[2, 1]. This issue is similar to WCF. In particular, a WCF protocol $\pi=(\pi_{1},\pi_{2})$ (where $\pi_{1}$ and $\pi_{2}$ are used instead of $\pi_{A}$ and $\pi_{B}$ to avoid confusion in subsequent protocol description) may find its stand-alone security insufficient for the upper-level task.
To illustrate, consider an attacking strategy for an unbalanced WCF protocol $\mu=(\mu_{A},\mu_{B})$ constructed from two instances of balanced WCF protocol $\pi=(\pi_{1},\pi_{2})$. The protocol $\mu$ aims to model the scenario where Alice and Bob are concurrently running two instances of WCF with an undetermined execution order. Assume the protocol $\pi$ follows the quantum comb model [39, 38] with $n+1$ (assume that $n$ is odd without loss of generality) rounds of communication, where Alice and Bob send messages alternately. Then $\pi_{1}$ and $\pi_{2}$ can be characterized by the ordered set $\mathcal{T}_{\pi}$ shown in Figure 5(b), where the subset $\{t_{i}\}_{i=0}^{n}$ is totally ordered. The comb model further requires that
  1. (R1)
    Messages are sent by $\pi_{1}$ and accepted by $\pi_{2}$ only at $t_{i}$ where $i$ is even.
  2. (R2)
    Messages are sent by $\pi_{2}$ and accepted by $\pi_{1}$ only at $t_{i}$ where $i$ is odd.
The details of these causal boxes are shown in Appendix A.3.
The protocol $\mu$ is constructed by connecting two independent instances of $\pi$ with the converter $\eta=(\eta_{A},\eta_{B})$, as shown in Figure 5(a). The converter $\eta_{A}$ accepts a classical bit $c_{A}$ at position $t_{n}$ from $\pi_{1}$ and a classical bit $c_{A}^{\prime}$ at position $t_{b}^{\prime}$ from $\pi_{2}^{\prime}$ and outputs $c_{A}^{\prime\prime}$ at position $t_{A}$, where $c_{A}^{\prime\prime}=c_{A}\land c_{A}^{\prime}.$
[…]
$t_{0}$$\mathcal{T}_{\pi}:$$\mathcal{T}_{\pi}^{\prime}:$$\mathcal{T}_{\eta_{A}}:$$t_{n}$$t^{\prime}_{n}$$t_{A}$$\mathcal{T}_{\eta_{B}}:$$t^{\prime}_{n}$$t_{n}$$t_{B}$$\mathcal{T}_{\mu_{A}}:$$t_{A}$$\mathcal{T}_{\mu_{B}}:$$t_{n}$$t^{\prime}_{0}$$t^{\prime}_{n}$$t_{0}$$t_{n}$$t^{\prime}_{0}$$t^{\prime}_{n}$$t_{0}$$t_{n}$$t^{\prime}_{0}$$t^{\prime}_{n}$$t_{B}$
Figure 5: (a) The protocol $\mu=(\mu_{A},\mu_{B})$ is constructed from protocol $\pi$, $\pi^{\prime}$ and $\eta$. $\eta=(\eta_{A},\eta_{B})$ is a pair of converters and the two instances $\pi^{\prime}=(\pi^{\prime}_{1},\pi^{\prime}_{2})$ and $\pi=(\pi_{1},\pi_{2})$ are the same except the ordered sets $\mathcal{T}_{\pi}^{\prime},\mathcal{T}_{\pi}$ are independent. (b) This figure shows the Hasse diagram of the posets $\mathcal{T}_{\pi},\mathcal{T}_{\pi}^{\prime},\mathcal{T}_{\eta_{A}},\mathcal{T}_{\eta_{B}},\mathcal{T}_{\mu_{A}},\mathcal{T}_{\mu_{B}}$. The Hasse diagram depicts the order relation with directed edge and omits the transitive and reflexive connections for simplicity. The poset $\mathcal{T}_{\mu_{A}}$ is obtained by union of $\mathcal{T}_{\pi},\mathcal{T}_{\pi}^{\prime},\mathcal{T}_{\eta_{A}}$, which is defined in Definition 17. The poset of $\mathcal{T}_{\mu_{b}}$ is obtained similarly.
We demonstrate that the stand-alone security of $\mu$ can be compromised even if that of $\pi$ is maintained. As per [29, Lemma 4], if $\pi$ is a $\mathrm{WCF}(\frac{1}{2},\epsilon)$ protocol and $\epsilon<1/6$, $\mu$ is expected to be a $\mathrm{WCF}(\frac{3}{4},\frac{3}{2}\epsilon)$ protocol. However, this expectation is not met as there exists a strategy for a dishonest Bob to win the game with probability $\frac{1}{2}$. Before elaborating on this strategy, we introduce the delay box, denoted by $\Theta_{f}$. This box essentially transforms a message at position $t$ to a future position $t^{\prime}$ according to some delay function $f$. For a rigorous definition of delay box, see Appendix A.3.
Consider the parallel composition of delay boxes $\theta=\theta_{1}\|\theta_{2}$ depicted in Figure 6. The delay functions are defined as
$f_{1}:\{t_{i}\in\mathcal{T}_{\pi}\mid\text{$i$ is even}\}\to\mathcal{T}_{\theta_{1}},t_{i}\mapsto t_{i}^{\prime},$ (23)
$f_{2}:\{t_{i}^{\prime}\in\mathcal{T}_{\pi}^{\prime}\mid\text{$i$is odd}\}\to\mathcal{T}_{\theta_{2}},t_{i}^{\prime}\mapsto t_{i}.$ (24)
[…]
$t_{0}$$t_{n}$$t^{\prime}_{0}$$t^{\prime}_{n}$$t_{1}$$t^{\prime}_{1}$$t_{2}$$t^{\prime}_{2}$$t_{3}$$t^{\prime}_{3}$$t_{n-1}$$t^{\prime}_{n-1}$$\mathcal{T}_{\theta_{1}}:$$\mathcal{T}_{\theta_{2}}:$$\mathcal{T}_{(\pi_{1}\|\pi^{\prime}_{2})\theta}:$$t_{0}$$t_{n}$$t^{\prime}_{1}$$t^{\prime}_{0}$$t_{1}$$t^{\prime}_{n}$
Figure 6: (a) This figure shows the connection of $\mu_{A}$ and delay box $\theta$. The box $\theta=\theta_{1}\|\theta_{2}$ is a parallel composition of two delay boxes. (b) This figure shows the Hasse diagram of the ordered sets $\mathcal{T}_{\theta_{1}}$, $\mathcal{T}_{\theta_{2}}$ and $\mathcal{T}_{(\pi_{1}\|\pi_{2}^{\prime})\theta}$.
Proposition 8 shows that Figure 6(a) emulates a WCF protocol between Alice and herself. See Appendix B.1 for the proof.
Proposition 8.
The box $(\pi_{1}\|\pi_{2}^{\prime})\theta$ has the same output distribution as $\pi_{1}\pi_{2}$.
Since $\pi=(\pi_{1},\pi_{2})$ is a balanced WCF protocol, the distribution of the final output $c_{A}^{\prime\prime}$ is $P_{c_{A}^{\prime\prime}}=(1/2,1/2)$, which contradicts our expectation that $\mu$ is a $\mathrm{WCF}(\frac{3}{4},\frac{3}{2}\epsilon)$ protocol.
This example suggests that a stronger criterion is necessary to guarantee the safe composition of WCF protocols. To achieve a $\mathrm{WCF}(\frac{3}{4},\frac{3}{2}\epsilon)$ protocol, it is essential that for any box connected to the inner interface of $\pi_{1}\|\pi_{2}^{\prime}$, the output satisfies
$\forall x\in\{0,1\},\Pr[c_{A}^{\prime}=1|c_{A}=x]\leq\frac{1}{2}+\epsilon.$ (25)
More generally, to construct an unbalanced WCF protocol $\mathrm{WCF}(z,\epsilon^{\prime})$ for arbitrary $z\in[0,\frac{1}{2}]$ with $n$ instances of protocol $\pi$, the following condition is needed:
$\forall i\in[n],\forall x^{i-1}_{1}\in\{0,1\}^{i-1},\Pr[c_{A,i}=1|c_{A,1}=x_{1},\ldots,c_{A,i-1}=x_{i-1}]\leq\frac{1}{2}+\epsilon,$ (26)
where $x_{1}^{i}=x_{1}x_{2}\dots x_{i}$ and the result of the $i$th instance is denoted by $c_{A,i}$. Notably, inequality (26) is an implicit but crucial condition for the security of unbalanced WCF protocol introduced in [29].
4.2 Global security
What is the sufficient condition for Inequality (26) to hold? More broadly speaking, what is the safe way to compose WCF protocols? Despite negative results discussed in Section 3 on the possibility of abstracting WCF as a black-box abstraction, there are still certain ways of composition, satisfying Inequality (26). To explore this, we generalize Inequality (26), introduce the concept of global secrecy and specify the sufficient condition for it.
Definition 9 (Global security of WCF in a larger system).
Let $\pi$ be a $\mathrm{WCF}(z,\epsilon)$ protocol and $\eta$ be a protocol that only outputs classical messages $K_{A},K_{A}^{\prime},K_{B},K_{B}^{\prime}\in\mathbb{N}$ at its outer interface, where $K_{A},K_{B}$ are output before $C_{A},C_{B}$ and $K_{A}^{\prime},K_{B}^{\prime}$ are output after $C_{A},C_{B}$. For a new protocol $\xi=(\xi_{A},\xi_{B})$ with $\xi_{A}=\pi_{A}\|\eta_{A},\xi_{B}=\pi_{B}\|\eta_{B}$ (denoted by $\xi=\pi\|\eta$), the WCF protocol $\pi$ is said to be globally secure in $\xi$ if:
  1. 1.
    For any box $\alpha$ compatible with the inner interface of $\xi_{A}$: $\forall k\in\mathbb{N},\Pr[C_{A}=1|K_{A}=k]\leq z+\epsilon$,
  2. 2.
    For any box $\beta$ compatible with the inner interface of $\xi_{B}$: $\forall k\in\mathbb{N},\Pr[C_{B}=0|K_{B}=k]\leq 1-z+\epsilon$.
The global security defined above is a property of the whole composed system $\pi\|\eta$ rather than the protocol $\eta$ alone. For some protocols $\eta$ and $\eta^{\prime}$, a protocol $\pi$ may be globally secure in $\pi\|\eta$ but not in $\pi\|\eta^{\prime}$.
Notably, global security in this definition is much weaker than the universal composability in the AC framework. It is not within any composable framework and only applies to the scenario where the system composed with WCF in parallel takes no input from the outer interface. This assumption is even stricter than the non-interactive assumption in the bounded storage model [24]. Therefore, global security of all components in a larger system only ensures the stand-alone security of the larger system.
The following theorem formulates the intuition that the WCF protocol is not affected by other boxes as long as we forbid any other interaction when the WCF protocol is active.
Theorem 10.
Let $\pi$ be a $\mathrm{WCF}(z,\epsilon)$ protocol and $\eta$ be any protocol with only classical output at the outer interface. $\xi=\pi\|\eta$ is their parallel composition with the ordered set $\mathcal{T}$. Then $\pi$ is globally secure in the protocol if there exists a partition $\mathcal{T}_{1},\mathcal{T}_{2},\mathcal{T}_{3}$ of $\mathcal{T}$ such that
  1. (C1)
    $\forall t_{1}\in\mathcal{T}_{1},t_{2}\in\mathcal{T}_{2},t_{3}\in\mathcal{T}_{3}$, $t_{1}\prec t_{2}\prec t_{3}$;
  2. (C2)
    $\eta$ only accepts input and produces output on set $\mathcal{T}_{1}\cup\mathcal{T}_{3}$, while $\pi$ only accepts input and produces output on set $\mathcal{T}_{2}$.
The proof of Theorem 10 is in Appendix B.2. The proof idea is by contradiction. If the adversary can break the global security of the composed protocol, then the adversary can also break the stand-alone security of the WCF protocol, which contradicts the fact and thus completes the proof. Theorem 10 establishes the sufficient condition that the WCF protocols can be composed while maintaining the overall stand-alone security.
For more complicated functionalities, it is usually easy to construct asymmetric imperfect protocols in which one party can cheat more than the other party. Unbalanced WCF protocols make it possible for two parties to fairly choose one of these asymmetric imperfect protocols at random. This results in a less asymmetric imperfect protocol. Therefore, unbalanced WCF protocols play a key role in constructing optimal protocols of more complicated functionalities such as SCF and BC [29, 28]. An unbalanced $\mathrm{WCF}(z,\epsilon^{\prime})$ protocol can be construct from balanced $\mathrm{WCF}(\frac{1}{2},\epsilon)$ protocols where $z=0.b_{1}b_{2}\dots b_{n}$ is in binary representation, as is shown in Figure 7.
[…]
While $i<n$
   If $c_{i}\neq b_{i}$
      break;
   Endif
   $i=i+1;$
Endwhile
Output
$c=c_{i}$
$\mathcal{T}_{2}$$\mathcal{T}_{1}$$\mathcal{T}_{n}$(b)
Figure 7: (a) This figure shows Alice part of the unbalanced WCF protocol $\eta_{A}=\tau_{A}(\pi_{A,1}\|\dots\|\pi_{A,n})$ constructed from balanced protocols $\pi$. It runs $n$ instances of $\pi_{A}$ and processes their result in $\tau_{A}$. The program of $\tau_{A}$ is presented in the dashed box. Bob’s part works similarly. (b) This figure shows the form of the Hasse diagram of $\eta_{A}$, where $\mathcal{T}_{i}$ is the posets for instance $\pi_{A,i}$. The instances of $\pi_{A}$ run in a time-sequential way.
As an example, we apply Theorem 10 to this construction to prove the stand-alone security of the unbalanced WCF protocol in Corollary 11. The idea is to prove the stand-alone security of $\tau_{i}(\pi_{1}\|...\|\pi_{i})$ constructed from $\tau_{i-1}(\pi_{1}\|...\|\pi_{i-1})$ and $\pi_{i}$ using Theorem 10 recursively. The detailed proof can be found in Appendix B.3.
Corollary 11.
Let $z\in[0,\frac{1}{2}]$. The unbalanced WCF protocol shown in Figure 7 is a $\mathrm{WCF}(z,\epsilon^{\prime})$ protocol with $\epsilon^{\prime}=2\epsilon+o(\epsilon)$.
5 Acknowledgement
We thank Christopher Portmann and Vilasini Venkatesh for useful discussions. This research was initiated while AB was hosted at CQT. This research is supported by the National Research Foundation, Singapore and A*STAR under its CQT Bridging Grant and its Quantum Engineering Programme (NRF2021-QEP2-01-P06). AB acknowledges the partial support provided by Commonwealth Cyber Initiative (CCI-SWVA) under the 2023 Cyber Innovation Scholars Program.
Appendix A Causal box model
In this appendix, we introduce necessary concepts related to the causal box model. Part of this section is from [35].
A.1 Causal box
Let $\mathcal{T}$ be a partially ordered set (poset). That is, $\mathcal{T}$ is a set equipped with a binary relation $\preceq$ that is reflexive, antisymmetric and transitive. For $x,y\in\mathcal{T}$, $x\prec y$ if $x\preceq y$ and $x\neq y$. Denote $\mathcal{T}^{\preceq t}=\{x\in\mathcal{T}\mid x\preceq t\}$.
Definition 12 (Cuts).
A subset $\mathcal{C}$ of poset $\mathcal{T}$ is a cut if it takes the form
$\mathcal{C}=\bigcup_{t\in\mathcal{C}}\mathcal{T}^{\preceq t}.$ (27)
A cut is bounded if $\exists t\in\mathcal{T},\mathcal{C}\in\mathcal{T}^{\preceq t}$. The set of all cuts of $\mathcal{T}$ is denoted by $\mathfrak{C}(\mathcal{T})$. The set of all bounded cuts of $\mathcal{T}$ is denoted by ¯$\bar{\mathfrak{C}}(\mathcal{T})$.
Lemma 13 (An equivalent condition for cut).
A subset $\mathcal{C}$ of a poset $\mathcal{T}$ is a cut if and only if $t\in\mathcal{C}\Rightarrow\forall x\preceq t,x\in\mathcal{C}$.
Proof.
If $\mathcal{C}$ is a cut, for $t\in\mathcal{C}$ and $x\preceq t$, we have $x\in\mathcal{T}^{\preceq t}$, thus $x\in\mathcal{C}$. If $\forall t\in\mathcal{C},\forall x\preceq t,x\in\mathcal{C}$, then $\forall x\in\mathcal{T}^{\preceq t},x\in\mathcal{T}$. We have $\forall t\in\mathcal{C},\mathcal{T}^{\preceq t}\subseteq\mathcal{C}$, thus $\mathcal{C}=\bigcup_{t\in\mathcal{C}}\mathcal{T}^{\preceq t}$. ∎
Definition 14 (Causality function).
A function $\chi:\mathfrak{C}(\mathcal{T})\to\mathfrak{C}(\mathcal{T})$ is a causality function if:
$\forall\mathcal{C},\mathcal{D}\in\mathfrak{C}(\mathcal{T}),\quad\chi(\mathcal{C}\cup\mathcal{D})=\chi(\mathcal{C})\cup\chi(\mathcal{D})\,,$ (28)
$\forall\mathcal{C},\mathcal{D}\in\mathfrak{C}(\mathcal{T}),\quad\mathcal{C}\subseteq\mathcal{D}\implies\chi(\mathcal{C})\subseteq\chi(\mathcal{D})\,,$ (29)
$\forall\mathcal{C}\in\bar{\mathfrak{C}}(\mathcal{T})\setminus\{\emptyset\},\quad\chi(\mathcal{C})\subsetneq\mathcal{C}\,,$ (30)
$\forall\mathcal{C}\in\bar{\mathfrak{C}}(\mathcal{T}),\forall t\in\mathcal{C},\exists n\in\mathbb{N},\quad t\notin\chi^{n}(\mathcal{C})\,,$ (31)
Using cuts and causality function, we can formally define causal box.
Definition 15 (Causal box).
A causal box is a system with input wire $X$ and output wire $Y$, defined as a set of CPTP maps:
$\Phi\coloneqq\{\Phi^{\mathcal{C}}:\mathfrak{T}(\mathcal{F}_{X}^{\mathcal{T}})\to\mathfrak{T}(\mathcal{F}_{Y}^{\mathcal{C}})\}_{\mathcal{C}\in\bar{\mathfrak{C}}(\mathcal{T})},$ (32)
where $\mathfrak{T}(\mathcal{H})$ denotes the set of all trace class operators on the Hilbert space $\mathcal{H}$. These maps must be mutually consistent
$\forall\mathcal{C},\mathcal{D}\in\bar{\mathfrak{C}}(\mathcal{T}),\mathcal{C}\subseteq\mathcal{D}:\Phi^{\mathcal{C}}=\mathrm{tr}_{\mathcal{D}\setminus\mathcal{C}}\circ\Phi^{\mathcal{D}}$ (33)
and respect causality, i.e., there exists a causality function $\chi:\mathfrak{C}(\mathcal{T})\to\mathfrak{C}(\mathcal{T})$ such that
$\Phi^{\mathcal{C}}=\Phi^{\mathcal{C}}\circ\mathrm{tr}_{\mathcal{T}\setminus\chi(\mathcal{C})}.$ (34)
Alternatively, the causal box $\Phi$ can be described in Choi-Jamiolkowski (CJ) representation. To be precise, the Choi operator $R_{\Phi}^{\mathcal{C}}$ for the CPTP map $\Phi^{\mathcal{C}}$ is a sesquilinear semi-definite form on $\mathcal{F}_{Y}^{\mathcal{C}}\otimes\mathcal{F}_{X}^{\chi(\mathcal{C})}$ such that
$R_{\Phi}^{\mathcal{C}}(\psi_{Y}\otimes\psi_{X},\varphi_{Y}\otimes\varphi_{X})=\langle\psi_{Y}|\Phi^{\mathcal{C}}(|\bar{\psi}_{X}\rangle\langle\bar{\varphi}_{X}|)|\varphi_{Y}\rangle.$ (35)
where $|\overline{\psi}\rangle=\sum_{i}|i\rangle\langle\psi|i\rangle$ for a fixed basis $\{|i\rangle\}_{i}$.
A.2 Connection of causal boxes
Causal boxes can be connected by parallel composition and loop.
Definition 16 (Parallel composition of causal box).
Let $\Phi=\{\Phi^{\mathcal{C}}\}_{\mathcal{C}\in\bar{\mathfrak{C}}(\mathcal{T})}$ and $\Psi=\{\Psi^{\mathcal{C}}\}_{\mathcal{C}\in\bar{\mathfrak{C}}(\mathcal{T})}$ be two causal boxes. The parallel composition of them, denoted by $\Phi\|\Psi$ is defined as
$\Gamma\coloneqq\left\{\Phi^{\mathcal{C}}\otimes\Psi^{\mathcal{C}}\right\}.$ (36)
The above definition of parallel composition implicitly requires the causal boxes $\Phi$ and $\Psi$ share the same partially ordered set $\mathcal{T}$. However, a general causal box may not contain the information of the causal structure of some other causal box, that is, there is no such common set $\mathcal{T}$. In this case, to make the composition possible, we propose a more general version of parallel composition. Before that, we introduce the union of posets that might have intersection.
A poset $\mathcal{T}$ can be treated as a tuple of a set and a partial order relation $\mathcal{T}=(\mathfrak{S},\preceq)$. We define the union of two posets as follows.
Definition 17 (Union of partial orders).
Let $\mathcal{T}_{1}=(\mathfrak{S}_{1},\preceq_{1})$ and $\mathcal{T}_{2}=(\mathfrak{S}_{2},\preceq_{2})$ be two compatible posets. Let the relation on the set $\mathfrak{S}=\mathfrak{S}_{1}\cup\mathfrak{S}_{2}$ be $\preceq=\preceq_{1}\sqcup\preceq_{2}$. Then $x\preceq y$ if and only if there exists a sequence $\{a_{i}\}_{i=0}^{n}\subseteq\mathfrak{S}_{1}\sqcup\mathfrak{S}_{2}$ such that $a_{1}=x$, $a_{n}=y$ and $\forall i\in[n]$ either $a_{i-1}\preceq_{1}a_{i}$ or $a_{i-1}\preceq_{2}a_{i}$. The union of posets, denoted by $\mathcal{T}_{1}\sqcup\mathcal{T}_{2}$, is then defined as
$\mathcal{T}_{1}\sqcup\mathcal{T}_{2}\coloneqq(\mathfrak{S}_{1}\cup\mathfrak{S}_{2},\preceq_{1}\sqcup\preceq_{2})=(\mathfrak{S},\preceq)$ (37)
In the above definition, $\mathcal{T}_{1}$ and $\mathcal{T}_{2}$ are said to be compatible if $\preceq_{1}\sqcup\preceq_{2}$ is still a partial order.
After introducing the union of posets, we are ready to define the parallel composition of two causal boxes with compatible posets.
Definition 18 (Parallel composition of causal boxes with arbitrary compatible ordered sets).
$\Phi=\{\Phi^{\mathcal{C}}\}_{\mathcal{C}\in\bar{\mathfrak{C}}(\mathcal{T}_{1})}$ and $\Psi=\{\Psi^{\mathcal{C}}\}_{\mathcal{C}\in\bar{\mathfrak{C}}(\mathcal{T}_{2})}$ are causal boxes defined on compatible partially ordered set $\mathcal{T}_{1}$ and $\mathcal{T}_{2}$ respectively, which can have arbitrary intersection. $\Phi$ has input wire $X_{1}$ and output wire $Y_{1}$. $\Psi$ has input wire $X_{2}$ and output wire $Y_{2}$. The parallel composition of $\Phi$ and $\Psi$, denoted by $\Gamma=\Phi\|\Psi$, is defined on the new poset $\mathcal{T}=\mathcal{T}_{1}\sqcup\mathcal{T}_{2}$ as
$\Gamma$$\coloneqq\{\Gamma^{\mathcal{C}}\}_{\mathcal{C}\in\bar{\mathfrak{C}}(\mathcal{T})},$ (38)
$\Gamma^{\mathcal{C}}$ $\coloneqq\Phi^{\mathcal{C}_{1}}\otimes\Psi^{\mathcal{C}_{2}}\circ\mathrm{tr}_{\mathcal{F}^{\mathcal{T}\setminus\mathcal{T}_{1}}_{X_{1}}\otimes\mathcal{F}^{\mathcal{T}\setminus\mathcal{T}_{2}}_{X_{2}}},$ (39)
where $\mathcal{C}_{1}=\mathcal{T}_{1}\cap\mathcal{C},\mathcal{C}_{2}=\mathcal{T}_{2}\cap\mathcal{C}$.
Lemma 1920 and 21 suggest that Definition 18 is well defined.
Lemma 19.
For compatible posets $\mathcal{T}_{1}$ and $\mathcal{T}_{2}$, if $\mathcal{C}$ is a cut in $\mathcal{T}_{1}\sqcup\mathcal{T}_{2}$, then $\mathcal{C}_{1}=\mathcal{C}\cap\mathcal{T}_{1}$ and $\mathcal{C}_{2}=\mathcal{C}\cap\mathcal{T}_{2}$ are cuts in $\mathcal{T}_{1}$ and $\mathcal{T}_{2}$ respectively.
Proof.
Notice that
$\mathcal{C}_{1}=\mathcal{C}\cap\mathcal{T}_{1}=\bigcup_{t\in\mathcal{C}}(\mathcal{T}^{\preceq t}\cap\mathcal{T}_{1}).$
Assume $t^{\prime}\in\mathcal{T}^{\preceq t}\cap\mathcal{T}_{1}$, then by Definition 17, $\forall x\in\mathcal{T}_{1},x\preceq_{1}t^{\prime}\Rightarrow x\preceq t\Rightarrow x\in\mathcal{T}^{\preceq t}\cap\mathcal{T}_{1}$. According to Lemma 13, $\mathcal{T}^{\preceq t}\cap\mathcal{T}_{1}$ is a cut in $\mathcal{T}_{1}$. As the union of cuts, $\mathcal{C}_{1}$ is also a cut. The same argument follows for $\mathcal{C}_{2}$. ∎
Lemma 20 (Existence of causality function).
Let $\Gamma=\Phi\|\Psi$ with the poset $\mathcal{T}=\mathcal{T}_{1}\sqcup\mathcal{T}_{2}$ be the parallel composition of causal boxes $\Phi$ and $\Psi$ with compatible posets $\mathcal{T}_{1}$ and $\mathcal{T}_{2}$. There exists a causality function $\chi$ on $\mathfrak{C}(\mathcal{T})$ such that:
$\Gamma^{\mathcal{C}}=\Gamma^{\mathcal{C}}\circ\mathrm{tr}_{\mathcal{F}_{X_{1}\oplus X_{2}}^{\chi(\mathcal{C})}}$ (40)
Proof.
Assume $\chi_{1}$ and $\chi_{2}$ are causality functions for $\Phi$ and $\Psi$, respectively. Construct the function $\chi:\mathfrak{C}(\mathcal{T})\to\mathfrak{C}(\mathcal{T})$ such that $\chi(\mathcal{C})=\chi_{1}(\mathcal{C}\cap\mathcal{T}_{1})\cup\chi_{2}(\mathcal{C}\cap\mathcal{T}_{2})$, then it is easy to verify the conditions in Definition 14. ∎
Lemma 21 (Associativity).
For boxes $\Phi,\Psi,\Theta$ with compatible posets $\mathcal{T}_{\Phi},\mathcal{T}_{\Psi},\mathcal{T}_{\Theta}$,
$(\Phi\|\Psi)\|\Theta=\Phi\|(\Psi\|\Theta).$ (41)
Proof.
This lemma follows directly from Definition 18. ∎
Definition 22 (Loop).
Let $\Phi=\{\Phi^{\mathcal{C}}:\mathfrak{T}(\mathcal{F}_{AB}^{\mathcal{T}})\to\mathfrak{T}(\mathcal{F}_{CD}^{\mathcal{C}})\}_{\mathcal{C}\in\bar{\mathfrak{C}}(\mathcal{T})}$ be a causal box such that $\dim\mathcal{H}_{B}=\dim\mathcal{H}_{C}$ is its CJ representation. Let $\{|k_{C}\rangle\}$ and $\{|l_{C}\rangle\}$ be any orthonormal bases of $\mathcal{F}_{C}^{\mathcal{C}}$ and $\{|k_{B}\rangle\}$ and $\{|l_{B}\rangle\}$ be corresponding bases of $\mathcal{F}_{C}^{\mathcal{C}}$. The new box by looping from wire $B$ to wire $C$, denoted by $\Psi=\Phi^{C\hookrightarrow B}$, is a set of maps $\{\Psi^{\mathcal{C}}:\mathfrak{T}(\mathcal{F}_{A}^{\mathcal{T}})\to\mathfrak{T}(\mathcal{F}_{D}^{\mathcal{C}})\}$ with CJ representation
$R_{\Psi}^{\mathcal{C}}(\psi_{D}\otimes\psi_{A},\varphi_{D}\otimes\varphi_{A})$ (42)
$=\sum_{k,l}R_{\Phi}^{\mathcal{C}}(k_{C}\otimes\psi_{D}\otimes\psi_{A}\otimes\bar{k}_{B},l_{C}\otimes\varphi_{D}\otimes\varphi_{A}\otimes\bar{l}_{B}),$ (43)
where the conjugate $|\overline{\psi}\rangle=\sum_{i}|i_{B}\rangle\langle\psi|i_{B}\rangle$ is taken with respect to the base $|i_{B}\rangle$ of $\mathcal{F}_{B}^{\mathcal{T}}$ on which the CJ representation is defined.
A.3 Delay box
Definition 23 (Delay function).
Let $\mathcal{T}$ be a poset and $\mathcal{G}=\{t\in\mathcal{T}\mid\exists t^{\prime}\in\mathcal{T},t\prec t^{\prime}\}$ is a subset of $\mathcal{T}$. A function $f:\mathcal{G}\to\mathcal{T}$ is a delay function on $\mathcal{T}$ if for any $t\in\mathcal{G}$, $t\prec f(t)$.
Proposition 24.
If $\mathcal{T}$ is a finite set, a delay function $f$ on $\mathcal{T}$ induces a causality function
$\chi_{f}(\mathcal{C})=\bigcup_{t:f(t)\in\mathcal{C}}\mathcal{T}^{\leq t}.$ (44)
It is easy to verify $\chi_{f}$ is a causality function by definition.
Definition 25 (Delay box).
A delay box derived from a delay function $f:\mathcal{G}\to\mathcal{T}$ with input wire $X$ and output wire $X^{\prime}$ is denoted by $\Theta_{f}=\{\Theta_{f}^{\mathcal{C}}:\mathfrak{T}(\mathcal{F}_{X}^{\chi_{f}(\mathcal{C})})\to\mathfrak{T}(\mathcal{F}_{X}^{\mathcal{C}})\}_{\mathcal{C}\in\bar{\mathfrak{C}}(\mathcal{T})}$. $\mathcal{H}_{X}$ and $\mathcal{H}_{X^{\prime}}$ are of the same dimension. $\Theta_{f}^{\mathcal{C}}$ is defined as the following isometry:
$U_{f}^{\mathcal{C}}=\bigoplus_{n=0}^{\infty}\left(\sum_{i\in I}|i\rangle_{X^{\prime}}\langle i|_{X}\otimes\sum_{t\in\chi_{f}(\mathcal{C})}|f(t)\rangle\langle t|\right)^{\otimes n},$ (45)
for some bases $\{|i\rangle_{X}\}_{i\in I}$ and $\{|i\rangle_{X^{\prime}}\}_{i\in I}$ of spaces $\mathcal{H}_{X}$ and $\mathcal{H}_{X^{\prime}}$.
$\theta_{1}$ is a delay box from input wire $C^{\prime}$ to output wire $C^{\prime\prime}$ and is derived from $f_{1}$ defined in Equation 23. Then, $U_{f_{1}}^{\mathcal{C}}$ is an isometry between $\mathcal{F}_{C^{\prime}}^{\chi_{f_{1}}(\mathcal{C})}$ and $\mathcal{F}_{C^{\prime\prime}}^{\mathcal{C}\cap\mathcal{T}_{\pi^{\prime}}}$
Appendix B Proofs
B.1 Proof of Proposition 8
Since the protocol $\pi$ is restricted to the comb model, we can assume that $\pi_{1}$ only outputs at the set $\mathcal{T}_{\pi}^{\text{even}}=\{t_{i}\in\mathcal{T}_{\pi}\mid i\text{ is even}\}$ and $\pi_{2}$ only outputs at set $\mathcal{T}_{\pi}^{\text{odd}}=\{t_{i}\in\mathcal{T}_{\pi}\mid i\text{ is odd}\}$.
For simplicity, where the cut $\mathcal{C}$ is obvious, denote for $X=B,C,C^{\prime},C^{\prime\prime},E$
$\hat{X}$ $=\mathcal{F}_{X}^{\mathcal{C}\cap\mathcal{T}_{\pi}^{\text{even}}},$ (46)
$\tilde{X}$ $=\mathcal{F}_{X}^{\mathcal{C}\cap\mathcal{T}_{\pi^{\prime}}^{\text{even}}},$ (47)
where $\mathcal{T}_{\pi}$ is the poset for $\pi=(\pi_{1},\pi_{2})$ and $\mathcal{T}_{\pi^{\prime}}$ is the partial order for $\pi^{\prime}=(\pi_{2}^{\prime},\pi_{1}^{\prime})$. And denote the wire space of $X=A,D,D^{\prime},D^{\prime\prime},F$ by
$\hat{X}$ $=\mathcal{F}_{X}^{\mathcal{C}\cap\mathcal{T}_{\pi}^{\text{odd}}},$ (48)
$\tilde{X}$ $=\mathcal{F}_{X}^{\mathcal{C}\cap\mathcal{T}_{\pi^{\prime}}^{\text{odd}}}.$ (49)
For $\mathcal{C}\in\mathfrak{C}(\mathcal{T}_{\pi})$, $\{|i\rangle_{\hat{A}}\}$ is a orthonormal base of space $\mathcal{F}_{A}^{\chi_{\pi}(\mathcal{C})}$, The Choi-Jamiolkowski (CJ) representation of the box $\pi_{1}$ is a sesquilinear form $R_{\pi_{1}}^{\mathcal{C}}$ satisfying
$R_{\pi_{1}}^{\mathcal{C}}(\psi_{\hat{C}}\otimes\psi_{\hat{E}}\otimes\psi_{\hat{A}},\varphi_{\hat{C}}\otimes\varphi_{\hat{E}}\otimes\varphi_{\hat{A}})=\langle\psi_{\hat{C}}|\langle\psi_{\hat{E}}|\Pi_{1}^{\mathcal{C}}(|\bar{\psi}_{\hat{A}}\rangle\langle\bar{\varphi}_{\hat{A}}|)|\varphi_{\hat{E}}\rangle|\varphi_{\hat{C}}\rangle,$ (50)
where $\Pi_{1}^{\mathcal{C}}$ is the CPTP map for box $\pi_{1}$ on cut $\mathcal{C}$ and the conjugate of $\bar{\psi}_{\hat{A}},\bar{\varphi}_{\hat{A}}$ is taken on basis $\{|i\rangle_{\hat{A}}\}$. We only care about the effect of $R_{\pi_{1}}^{\mathcal{C}}$ on space $\hat{A},\hat{C},\hat{E}$ because of the conditions (R1),(R2) required by the comb model.
For $\mathcal{C}\in\mathcal{T}_{\pi}$, assume that the wire $B,C$ are connected with respect to the bases $\{|k_{\hat{C}}\rangle\},\{|l_{\hat{C}}\rangle\}$ of $\hat{C}$ and the bases $\{|k_{\hat{B}}\rangle\},\{|l_{\hat{B}}\rangle\}$ of $\hat{B}$, the wire $D,A$ are connected with respect to the orthonormal bases $\{|p_{\hat{D}}\rangle\}$, $\{|q_{\hat{D}}\rangle\}$ of $\hat{D}$ and the orthonormal bases $\{|p_{\hat{A}}\rangle\}$, $\{|q_{\hat{A}}\rangle\}$ of $\hat{A}$. The CJ representation of $\pi_{1}\pi_{2}$ is
$R_{\pi_{1}\pi_{2}}^{\mathcal{C}}(\psi_{\hat{E}}\otimes\psi_{\hat{F}},\varphi_{\hat{E}}\otimes\varphi_{\hat{F}})$ (51)
$=\sum_{k,l,p,q}R_{\pi_{1}}^{\mathcal{C}}(k_{\hat{C}}\otimes\psi_{\hat{E}}\otimes\bar{p}_{A},l_{\hat{C}}\otimes\varphi_{\hat{E}}\otimes\bar{q}_{\hat{A}})R_{\pi_{2}}^{\mathcal{C}}(p_{\hat{D}}\otimes\psi_{\hat{F}}\otimes\bar{k}_{\hat{B}},q_{\hat{D}}\otimes\varphi_{\hat{F}}\otimes\bar{l}_{\hat{B}})$
For $\mathcal{C}^{\prime}\in\mathfrak{C}(\mathcal{T}_{\theta_{1}})$, $\{|i\rangle_{\hat{C}^{\prime}}\}$ is a base of space $\mathcal{F}_{C^{\prime}}^{\chi_{\theta_{1}}(\mathcal{C}^{\prime})}$, and $\{|i\rangle_{\tilde{C}^{\prime\prime}}\}$ is a base of space $\mathcal{F}_{C^{\prime\prime}}^{f_{1}(\chi_{\theta_{1}}(\mathcal{C}^{\prime}))}$ such that $U_{f_{1}}|i\rangle_{\hat{C}}=|i\rangle_{\tilde{C}^{\prime\prime}}$  [^1] This is possible because $\mathcal{F}_{C^{\prime}}^{\chi_{\theta_{1}}(\mathcal{C}^{\prime})}\simeq\mathcal{F}_{C^{\prime\prime}}^{f_{1}(\chi_{\theta_{1}}(\mathcal{C}^{\prime}))}$.. The CJ representation of the delay box $\theta_{1}$ is
$R_{\theta_{1}}^{\mathcal{C}^{\prime}}(\psi_{\hat{C}^{\prime\prime}}\otimes\psi_{\tilde{C}^{\prime\prime}}\otimes\alpha_{\hat{C}^{\prime}},\varphi_{\hat{C}^{\prime\prime}}\otimes\varphi_{\tilde{C}^{\prime\prime}}\otimes\beta_{\hat{C}^{\prime}})$
$=\langle\psi_{\tilde{C}^{\prime\prime}}|\Theta_{1}^{\mathcal{C}^{\prime}}\left(|\bar{\alpha}_{\hat{C}^{\prime}}\rangle\langle\bar{\beta}_{\hat{C}^{\prime}}|\right)|\varphi_{\tilde{C}^{\prime\prime}}\rangle\cdot\langle\psi_{\hat{C}^{\prime\prime}}|\Omega_{\hat{C}^{\prime\prime}}\rangle\langle\Omega_{\hat{C}^{\prime\prime}}|\varphi_{\hat{C}^{\prime\prime}}\rangle$
$=\langle\psi_{\tilde{C}^{\prime\prime}}|\bar{\alpha}_{\tilde{C}^{\prime\prime}}\rangle\langle\bar{\beta}_{\tilde{C}^{\prime\prime}}|\varphi_{\tilde{C}^{\prime\prime}}\rangle\cdot\langle\psi_{\hat{C}^{\prime\prime}}|\Omega_{\hat{C}^{\prime\prime}}\rangle\langle\Omega_{\hat{C}^{\prime\prime}}|\varphi_{\hat{C}^{\prime\prime}}\rangle,$ (52)
where $\bar{\alpha}_{\tilde{C}^{\prime\prime}}\in\mathcal{F}_{C^{\prime\prime}}^{\mathcal{C}^{\prime}}$ and $\bar{\alpha}_{\hat{C}^{\prime}}\in\mathcal{F}_{C^{\prime}}^{\chi(\mathcal{C}^{\prime})}$ have the same coefficients in the bases of $\{|i\rangle_{\tilde{C}^{\prime\prime}}\}$ and $\{|i\rangle_{\hat{C}^{\prime}}\}$. Since the state on space $\hat{C}^{\prime\prime}$ is always a vacuum state, we omit the arguments $\psi_{\hat{C}^{\prime\prime}},\varphi_{\hat{C}^{\prime\prime}}$ of $R_{\theta_{1}^{\mathcal{C}^{\prime}}}$ and the second term of Equation (52) in the remaining part of the paper.
For all $\mathcal{C}\in\bar{\mathfrak{C}}(\mathcal{T}_{\pi_{1}\theta_{1}})$, wire $C$ and $C^{\prime}$ are connected with respect the bases $\{|k_{\hat{C}^{\prime}}\rangle\}$ and $\{|l_{\hat{C}^{\prime}}\rangle\}$ such that $|k_{\hat{C}^{\prime}}\rangle\cong|k_{\hat{C}}\rangle,|l_{\hat{C}^{\prime}}\rangle\cong|l_{\hat{C}}\rangle$. the CJ representation of the composition box $\bar{\pi}_{1}=\pi_{1}\theta_{1}$ is,
$R_{\bar{\pi}_{1}}^{\mathcal{C}}(\gamma_{\tilde{C}^{\prime\prime}}\otimes\psi_{\hat{E}}\otimes\psi_{\hat{A}},\eta_{\tilde{C}^{\prime\prime}}\otimes\varphi_{\hat{E}}\otimes\varphi_{\hat{A}})$
$=\sum_{k,l}R_{\pi_{1}}^{\mathcal{C}_{1}}(k_{\hat{C}}\otimes\psi_{\hat{E}}\otimes\psi_{\hat{A}},l_{\hat{C}}\otimes\varphi_{\hat{E}}\otimes\varphi_{\hat{A}})R_{\theta_{1}}^{\mathcal{C}_{2}}\left(\gamma_{\tilde{C}^{\prime\prime}}\otimes\bar{k}_{\hat{C}^{\prime}},\eta_{\tilde{C}^{\prime\prime}}\otimes\bar{l}_{\hat{C}^{\prime}}\right)$ (53)
$=\sum_{k,l}R_{\pi_{1}}^{\mathcal{C}_{1}}\left(k_{\hat{C}}\otimes\psi_{\hat{E}}\otimes\psi_{\hat{A}},l_{\hat{C}}\otimes\varphi_{\hat{E}}\otimes\varphi_{\hat{A}}\right)\langle\gamma_{\tilde{C}^{\prime\prime}}|k_{\tilde{C}^{\prime\prime}}\rangle\langle l_{\tilde{C}^{\prime\prime}}|\eta_{\tilde{C}^{\prime\prime}}\rangle$ (54)
$=R_{\pi_{1}}^{\mathcal{C}_{1}}\left(\sum_{k}\langle k_{\tilde{C}^{\prime\prime}}|\gamma_{\tilde{C}^{\prime\prime}}\rangle k_{\hat{C}}\otimes\psi_{\hat{E}}\otimes\psi_{\hat{A}},\sum_{l}\langle l_{\tilde{C}^{\prime\prime}}|\eta_{\tilde{C}^{\prime\prime}}\rangle l_{\hat{C}}\otimes\varphi_{\hat{E}}\otimes\varphi_{\hat{A}}\right)$ (55)
$=R_{\pi_{1}}^{\mathcal{C}_{1}}\left(\gamma_{\hat{C}}\otimes\psi_{\hat{E}}\otimes\psi_{\hat{A}},\eta_{\hat{C}}\otimes\varphi_{\hat{E}}\otimes\varphi_{\hat{A}}\right),$ (56)
where $|\gamma_{\hat{C}}\rangle=\left(\sum_{k}|k_{\hat{C}}\rangle\langle k_{\tilde{C}^{\prime\prime}}|\right)|\gamma_{\tilde{C}^{\prime\prime}}\rangle$ and $|\eta_{\hat{C}}\rangle=\left(\sum_{l}|l_{\hat{C}}\rangle\langle l_{\tilde{C}^{\prime\prime}}|\right)|\eta_{\tilde{C}^{\prime\prime}}\rangle$. Equation (53) follows from Definition 18, Equation (54) follows from Equation (52) and Equation (55) follows from the sesquilinearity.
Similarly, for all $\mathcal{C}\in\bar{\mathfrak{C}}(\mathcal{T}_{\pi_{2}^{\prime}\theta_{2}})$, wire $D$ and $D^{\prime}$ are connected with respect the bases $\{|p_{\tilde{D}^{\prime}}\rangle\}$ and $\{|q_{\tilde{D}^{\prime}}\rangle\}$ such that $|p_{\tilde{D}^{\prime}}\rangle\cong|p_{\tilde{D}}\rangle,|q_{\tilde{D}^{\prime}}\rangle\cong|q_{\tilde{D}}\rangle$. The CJ representation of the box $\bar{\pi}^{\prime}_{2}=\pi^{\prime}_{2}\theta_{2}$ is
$R_{\bar{\pi}^{\prime}_{2}}^{\mathcal{C}}\left(\lambda_{\hat{D}^{\prime\prime}}\otimes\psi_{\tilde{F}}\otimes\psi_{\tilde{B}},\omega_{\hat{D}^{\prime\prime}}\otimes\varphi_{\tilde{F}}\otimes\varphi_{\tilde{B}}\right)$ (57)
$=R_{\pi^{\prime}_{2}}^{\mathcal{C}_{1}}\left(\lambda_{\tilde{D}}\otimes\psi_{\hat{F}}\otimes\psi_{\hat{B}},\omega_{\tilde{D}}\otimes\varphi_{\hat{F}}\otimes\varphi_{\hat{B}}\right),$ (58)
where $|\lambda_{\tilde{D}}\rangle=\left(\sum_{p}|p_{\tilde{D}}\rangle\langle p_{\hat{D}^{\prime\prime}}|\right)|\lambda_{\hat{D}^{\prime\prime}}\rangle$ and $|\omega_{\tilde{D}}\rangle=\left(\sum_{q}|q_{\tilde{D}}\rangle\langle q_{\hat{D}^{\prime\prime}}|\right)|\omega_{\hat{D}^{\prime\prime}}\rangle$.
Notice that, for all $\mathcal{C}\in\bar{\mathfrak{C}}(\mathcal{T}_{(\pi_{1}\|\pi^{\prime}_{2})})$, we have the relation chain of the base vectors
$|k_{\tilde{B}}\rangle\cong|k_{\hat{B}}\rangle\cong|k_{\hat{C}}\rangle\cong|k_{\hat{C}^{\prime}}\rangle\cong|k_{\tilde{C}^{\prime\prime}}\rangle,$ (59)
$|l_{\tilde{B}}\rangle\cong|l_{\hat{B}}\rangle\cong|l_{\hat{C}}\rangle\cong|l_{\hat{C}^{\prime}}\rangle\cong|l_{\tilde{C}^{\prime\prime}}\rangle,$
$|p_{\hat{A}}\rangle\cong|p_{\tilde{A}}\rangle\cong|p_{\tilde{D}}\rangle\cong|p_{\tilde{D}^{\prime}}\rangle\cong|p_{\hat{D}^{\prime\prime}}\rangle,$
$|q_{\hat{A}}\rangle\cong|q_{\tilde{A}}\rangle\cong|q_{\tilde{D}}\rangle\cong|q_{\tilde{D}^{\prime}}\rangle\cong|q_{\hat{D}^{\prime\prime}}\rangle.$
The first column of $\cong$ follows naturally by simply substituting $t_{i}\in\mathcal{T}_{\pi}$ with $t^{\prime}_{i}\in\mathcal{T}_{\pi^{\prime}}$, The second column is defined by the connection of $\pi_{1},\pi_{2}$. The third column is defined by the connection inside $\bar{\pi}_{1}$ and $\bar{\pi}^{\prime}_{2}$. The fourth column follows from the effect of delay boxes $\theta_{1}$ and $\theta_{2}$.
Then for $\mathcal{C}=\mathcal{T}_{(\pi_{1}\|\pi_{2}^{\prime})\theta}$, we choose the bases $\{|k\rangle_{\tilde{C}^{\prime\prime}}\},\{|l\rangle_{\tilde{C}^{\prime\prime}}\},\{|k\rangle_{\hat{B}}\},\{|l\rangle_{\hat{B}}\}$ for the connection of $C^{\prime\prime},B$ and the bases $\{|p_{\hat{D}^{\prime\prime}}\rangle\},\{|q_{\hat{D}^{\prime\prime}}\rangle\},\{|p_{\tilde{A}}\rangle\},\{|q_{\tilde{A}}\rangle\}$ for the connection of $D^{\prime\prime},A$. The CJ representation of the final box $\bar{\pi}_{1}\bar{\pi^{\prime}_{2}}$ is
$R_{\bar{\pi}_{1}\bar{\pi}^{\prime}_{2}}^{\mathcal{T}_{(\pi_{1}\|\pi_{2})\theta}}(\psi_{\hat{E}}\otimes\psi_{\tilde{F}},\varphi_{\hat{E}}\otimes\varphi_{\tilde{F}})$ (60)
$=\sum_{k,l,p,q}R_{\bar{\pi}_{1}}^{\mathcal{T}_{\bar{\pi}_{1}}}(k_{\tilde{C}^{\prime\prime}}\otimes\psi_{E}\otimes\bar{p}_{A},l_{\tilde{C}^{\prime\prime}}\otimes\varphi_{\hat{E}}\otimes\bar{q}_{\hat{A}})R_{\bar{\pi}^{\prime}_{2}}^{\mathcal{T}_{\bar{\pi}^{\prime}_{2}}}(p_{\hat{D}^{\prime\prime}}\otimes\psi_{\tilde{F}}\otimes\bar{k}_{\tilde{B}},q_{\hat{D}^{\prime\prime}}\otimes\varphi_{\tilde{F}}\otimes\bar{l}_{\tilde{B}})$ (61)
$=\sum_{k,l,p,q}R_{\pi_{1}}^{\mathcal{T}_{\pi}}(k_{\hat{C}}\otimes\psi_{E}\otimes\bar{p}_{A},l_{\hat{C}}\otimes\varphi_{\hat{E}}\otimes\bar{q}_{\hat{A}})R_{\pi^{\prime}_{2}}^{\mathcal{T}_{\pi^{\prime}}}(p_{\tilde{D}}\otimes\psi_{\tilde{F}}\otimes\bar{k}_{\tilde{B}},q_{\tilde{D}}\otimes\varphi_{\tilde{F}}\otimes\bar{l}_{\tilde{B}}).$ (62)
where $\mathcal{C}_{1}^{\prime}=\mathcal{C}\cap\mathcal{T}_{\pi}$ and $\mathcal{C}_{2}^{\prime}=\mathcal{C}\cap\mathcal{T}_{\pi^{\prime}}$. Since $R_{\pi_{2}}$ and $R_{\pi^{\prime}_{2}}$ have the same effect on their respective spaces, there exist isomorphisms from $\hat{B},\hat{F},\hat{D}$ to $\tilde{B},\tilde{F},\tilde{D}$ such that
$R_{\pi^{\prime}_{2}}^{\mathcal{T}_{\pi^{\prime}}}(p_{\tilde{D}}\otimes\psi_{\tilde{F}}\otimes\bar{k}_{\tilde{B}},q_{\tilde{D}}\otimes\varphi_{\tilde{F}}\otimes\bar{l}_{\tilde{B}})=R_{\pi_{2}}^{\mathcal{T}_{\pi}}(p_{\hat{D}}\otimes\psi_{\hat{F}}\otimes\bar{k}_{\hat{B}},q_{\hat{D}}\otimes\varphi_{\hat{F}}\otimes\bar{l}_{\hat{B}}).$ (63)
Note that the isomorphisms here are exactly the ones used in the first column of Equation (59). Then we have
$R_{\bar{\pi}_{1}\bar{\pi}^{\prime}_{2}}^{\mathcal{T}_{(\pi_{1}\|\pi_{2})\theta}}(\psi_{\hat{E}}\otimes\psi_{\tilde{F}},\varphi_{\hat{E}}\otimes\varphi_{\tilde{F}})=R_{\pi_{1}\pi_{2}}^{\mathcal{T}_{\pi}}(\psi_{\hat{E}}\otimes\psi_{\hat{F}},\varphi_{\hat{E}}\otimes\varphi_{\hat{F}})$ (64)
B.2 Proof of Theorem 10
We first simplify the adversarial box and prove this theorem by contradiction. Let $\alpha$ be a box connected with the parallel composition $\pi_{A}\|\eta_{A}$, as shown in Figure 8(a). For convenience, $c_{A},K_{A}$ are relabeled as $E,K$. Let $\mathcal{T}$ be the ordered set of $\pi_{A}\|\eta_{A}$. Without loss of generality, we can assume the final box $(\pi_{A}\|\eta_{A})\alpha$ is defined on the ordered set $\mathcal{T}$, because any extra element $t$ introduced by $\alpha$ has no effect on $\pi_{A}\|\eta_{A}$.
[…]
$\mathcal{T}_{2}$$\mathcal{T}_{1}$$\mathcal{T}:$$\mathcal{T}_{3}$
Figure 8: (a) The basic connection. (b) An equivalent form of the box. The wire $X,Y,V,W$ are split into $X_{1},X_{2},Y_{1},Y_{2},V_{1},V_{2},W_{1},W_{2}$. (c) Causal order of the box $(\pi_{A}\|\eta_{A})\alpha$.
According to the condition (C1) and (C2), the wire $(V_{1},W_{1})$, $(B,D)$ and $(V_{2},W_{2})$ are effective only on $\mathcal{T}_{1}$, $\mathcal{T}_{2}$ and $\mathcal{T}_{3}$, respectively. Let $\{\Phi_{\alpha}^{\mathcal{C}}\}_{\mathcal{C}\in\bar{\mathfrak{C}}(\mathcal{T})}$ be the set of CPTP maps corresponding to $\alpha$. Since we are only interested in the final output, we set $\mathcal{C}=\mathcal{T}$. By [35, Lemma 9], the map $\Phi_{\alpha}^{\mathcal{T}}$ can be decomposed into three maps $\Phi_{\alpha_{1}}^{\mathcal{T}}:V_{1}\mapsto W_{1}Q_{1}$, $\Phi_{\alpha_{2}}^{\mathcal{T}}:BQ_{1}\mapsto DQ_{2}$ and $\Phi_{\alpha_{3}}^{\mathcal{T}}:V_{2}Q_{2}\mapsto W_{2}$ that satisfy
$\Phi_{\alpha}^{\mathcal{T}}=(\mathrm{id}_{W_{1}}\otimes\mathrm{id}_{D}\otimes\Phi_{\alpha_{3}}^{\mathcal{T}})\circ(\mathrm{id}_{W_{1}}\otimes\Phi_{\alpha_{2}}^{\mathcal{T}}\otimes\mathrm{id}_{V_{2}})\circ(\Phi_{\alpha_{1}}^{\mathcal{T}}\otimes\mathrm{id}_{B}\otimes\mathrm{id}_{V_{2}}).$ (65)
Here $\Phi_{\alpha_{1}}^{\mathcal{T}}$, $\Phi_{\alpha_{2}}^{\mathcal{T}}$ and $\Phi_{\alpha_{3}}^{\mathcal{T}}$ can be seen as the map of some boxes $\alpha_{1}$, $\alpha_{2}$ and $\alpha_{3}$, respectively. Similarly, the box $\eta_{A}$ is split into $\eta_{A1}$ and $\eta_{A2}$. Let $\Phi_{\eta_{A}\alpha_{1}}^{\mathcal{T}}$ and $\Phi_{\pi_{A}\alpha_{2}}^{\mathcal{T}}$ be the map of connected boxes $\eta_{A}\alpha_{1}$ and $\pi_{A}\alpha_{2}$. Denote $\bar{Q}=\mathcal{F}_{Q}^{\mathcal{T}}$, then the output of $\Phi_{\eta_{A}\alpha_{1}}^{\mathcal{T}}$ is a classical-quantum state $\rho_{\bar{Q}K}=\sum_{k}P_{K}(k)|k\rangle\langle k|\otimes\rho_{\bar{Q}}^{k}$. $\Phi_{\pi_{A}\alpha_{2}}^{\mathcal{T}}$ is a map from $\mathfrak{T}(\bar{Q})$ to the classical bit $E$, which can be represented as a binary POVM measurement $\{M_{0},M_{1}\}$. Then the conditional probability is $P_{E|K}(e|k)=\mathrm{tr}(M_{e}\rho_{\bar{Q}}^{k})$.
Assume $\exists k\in\mathbb{N},\Pr[E=1|K=k]\geq z+\epsilon$, then $\mathrm{tr}(M_{e}\rho_{\bar{Q}}^{k})\geq z+\epsilon$. Consider a box $\gamma$ that outputs state $\rho_{\bar{Q}}^{z}$ and has no input. The box $\pi_{A}\alpha_{2}\gamma$ will output $E=1$ with probability larger than $z+\epsilon$, which contradicts the stand-alone security. Therefore, the first condition in Definition 9 is satisfied. The second condition follows similarly.
B.3 Proof of Corollary 11
We denote the fractional part of $z$ as a binary string $b_{1}^{n}$. We denote the set $I=\{i\in\mathbb{N}\mid b_{i}=1\}$ as a string $a_{1}^{|I|}$ and the set $I^{\prime}=\{i\in\mathbb{N}\mid b_{i}=0\}$ as a string ${a^{\prime}}_{1}^{|I^{\prime}|}$ in ascending order. The $i$th WCF protocol $\pi_{i}$ as well as the other WCF protocols $\pi_{1}\|...\|\pi_{i-1}$ and $\pi_{i+1}\|...\|\pi_{n}$ output classical bits and follow strict temporal orders. Therefore, we obtain global security of $\pi_{i}$ according to Theorem 10
$\Pr\left[C_{i}=0|C_{1}^{i-1}=c_{1}^{i-1}\right]$ $\leq\frac{1}{2}+\epsilon,$ (66)
$\Pr\left[C_{i}=1|C_{1}^{i-1}=c_{1}^{i-1}\right]$ $\leq\frac{1}{2}+\epsilon.$ (67)
We first argue that a cheating party needs to maximize the winning probability of each $\pi_{i}$ in order to maximize the final winning probability. At each step $\pi_{i}$, if $b_{i}=1$, Alice wins with certainty if $C_{i}=0$ while wins with a probability smaller than $1$ if $C_{i}=1$; if $b_{i}=0$, Alice wins with a non-zero probability if $C_{i}=0$ while loses with certainty if $C_{i}=1$. Therefore, a cheating Alice strives to win at each step $\pi_{i}$ in order to finally win with the highest probability. It is the same for Bob. In conclusion, considering global security in Inequality (67), Alice’s winning probability is no larger than the case when
$\Pr\left[C_{i}=0|C_{1}^{i-1}=b_{1}^{i-1}\right]=\frac{1}{2}+\epsilon,$ (68)
and Bob’s winning probability is no larger than the case when
$\Pr\left[C_{i}=1|C_{1}^{i-1}=b_{1}^{i-1}\right]=\frac{1}{2}+\epsilon.$ (69)
We will use the above probabilities to compute the upper bound on Alice and Bob’s optimal winning probability.
For a cheating Alice,
Pr[Alice $\Pr[\text{Alice wins}]$ $=\sum_{i=1}^{|I|}\Pr\left[C_{1}^{a_{i}}=(b_{1}^{a_{i}-1},0)\right]$ (70)
$=\sum_{i=1}^{|I|}\Pr\left[C_{a_{i}}=0|C_{1}^{a_{i}-1}=b_{1}^{a_{i}-1}\right]\prod_{j=1}^{a_{i}-1}\Pr\left[C_{j}=b_{j}|C_{1}^{j-1}=b_{1}^{j-1}\right]$ (71)
$\leq\sum_{i=1}^{|I|}\left(\frac{1}{2}+\epsilon\right)^{a_{i}-i+1}\left(\frac{1}{2}-\epsilon\right)^{i-1},$ (72)
where $C_{1}^{0}=b_{1}^{0}$ is a certain event and Inequality (72) follows from Inequality (67). Similarly, for a cheating Bob, we obtain
Pr[Bob $\Pr[\text{Bob wins}]$ $=\sum_{i=1}^{|I^{\prime}|}\Pr\left[C_{1}^{a^{\prime}_{i}}=(b_{1}^{a^{\prime}_{i}-1},1)\right]$ (73)
$=\sum_{i=1}^{|I^{\prime}|}\Pr\left[C_{a^{\prime}_{i}}=1|C_{1}^{a^{\prime}_{i}-1}=b_{1}^{a^{\prime}_{i}-1}\right]\prod_{j=1}^{a^{\prime}_{i}-1}\Pr\left[C_{j}=b_{j}|C_{1}^{j-1}=b_{1}^{j-1}\right]$ (74)
$\leq\sum_{i=1}^{|I^{\prime}|}\left(\frac{1}{2}+\epsilon\right)^{a^{\prime}_{i}-i+1}\left(\frac{1}{2}-\epsilon\right)^{i-1},$ (75)
where we again apply Inequality (67) in (75). Then the bias is
$\epsilon^{\prime}_{A}$ $=|\Pr[\text{Alice wins}]-z|\leq\sum_{i=1}^{|I|}\left(\frac{1}{2}+\epsilon\right)^{a_{i}-i+1}\left(\frac{1}{2}-\epsilon\right)^{i-1}-\frac{1}{2^{a_{i}}}\leq\sum_{i=1}^{|I|}\frac{2(a_{i}-2i+2)}{2^{a_{i}}}\epsilon+o(\epsilon).$ (76)
We have omitted the absolute value because the winning probability of a cheating Alice is larger than that of an honest Alice, Pr[Alice $\Pr[\text{Alice wins}]\geq z$. Note that
$\sum_{n=i}^{\infty}\frac{n}{2^{n}}=\frac{i+1}{2^{i-1}}.$ (77)
When $|I|=1$,
$\epsilon^{\prime}_{A}\leq\frac{2a_{1}}{2^{a_{1}}}\epsilon+o(\epsilon)\leq\epsilon+o(\epsilon).$ (78)
When $|I|\geq 2$,
$\epsilon^{\prime}_{A}\leq\frac{2a_{1}}{2^{a_{1}}}\epsilon+\frac{1}{2}\sum_{i=2}^{|I|}\frac{(a_{i}-2)}{2^{a_{i}-2}}\epsilon+o(\epsilon)\leq\frac{2a_{1}}{2^{a_{1}}}\epsilon+\frac{1}{2}\sum_{i=2}^{\infty}\frac{i-2}{2^{i-2}}\epsilon+o(\epsilon)\leq 2\epsilon+o(\epsilon).$ (79)
In conclusion, for Alice,
$\epsilon^{\prime}_{A}=|\Pr[\text{Alice wins}]-z|\leq 2\epsilon+o(\epsilon).$ (80)
Similarly for Bob,
$\epsilon^{\prime}_{B}=|\Pr[\text{Bob wins}]-z|\leq 2\epsilon+o(\epsilon).$ (81)
The final bias is
$\epsilon^{\prime}=\max\{\epsilon^{\prime}_{A},\epsilon^{\prime}_{B}\}\leq 2\epsilon+o(\epsilon)$ (82)

### v3

@@
 Weak coin flipping is a cryptographic primitive in which two mutually distrustful parties generate a shared random bit to agree on a winner via remote communication.
-While a stand-alone secure weak coin flipping protocol can be constructed from noiseless communication channels, its composability has not been explored.
-In this work, we demonstrate that no weak coin flipping protocol can be abstracted into a simple black box resource with composable security.
-Despite this, we also establish the overall stand-alone security of weak coin flipping protocols under sequential composition.
+While a stand-alone secure weak coin flipping protocol can be constructed from noiseless quantum communication channels, its composability remains unexplored.
+In this work, we demonstrate that no weak coin flipping protocol can be abstracted as a simple black-box resource with composable security.
+Despite this, we also establish the overall stand-alone security of quantum weak coin flipping protocols under composition in sequential order.
 1 Introduction
@@
 We focus here on the model of Abstract Cryptography (AC) by Maurer and Renner [9, 10].
-In contrast to the previous frameworks, the latter is a top-down axiomatic approach and is intended to provide better insights via abstraction of a given complex cryptographic system and will thus be most suitable for our purposes.
+In contrast to the previous frameworks, the latter is a top-down axiomatic approach and is intended to provide better insights via abstraction of a given complex cryptographic system and is thus most suitable for our purposes.
 The security in the AC framework is built on simulation-based security.
@@
 The explicit protocols which are shown to have an arbitrarily small bias are developed in [17, 18, 19].
-The security notion employed in these results, known as information-theoretic security, allows an adversary to enjoy unlimited resources (i.e.
-the adversary has no restrictions on its computational power or use of memory, etc.).
+The security notion employed in these results, known as information-theoretic security, allows an adversary to enjoy unlimited resources (i.e., the adversary has no restrictions on its computational power or use of memory, etc.).
 It is well-studied for a range of different cryptographic tasks and, in most cases, is stronger than other commonly studied security definitions such as computational [20, 21, 22], bounded storage [23, 24, 25] or even game-based [26, 27], where an adversary can perform attacks in only some prescribed ways.
 Previously, it was shown that WCF enjoys information-theoretic stand-alone security, i.e., it is secure as a single protocol instance in isolation [16, 17, 18, 19].
-However, while WCF itself is of limited interest, it is mostly a powerful tool in constructing optimal protocols for several other primitives such as bit commitment [28] and strong coin flipping [29].
-However, such constructions would require more than stand-alone security and it is still unclear if WCF protocols have composable security when combined with other cryptographic protocols.
+While WCF itself is of limited interest, it is a powerful tool in constructing optimal protocols for several other primitives such as bit commitment [28] and strong coin flipping [29].
+However, such constructions would require more than stand-alone security, and it is still unclear if WCF protocols have composable security when combined with other cryptographic protocols.
 Contributions
 Our main contributions encompass two negative theorems and one positive theorem.
-In Theorem 5, we demonstrate the impossibility of constructing a typical black-box WCF resource analogous to the widely-used SCF resource [30] (see Definition 4) from noiseless communication channels in the AC framework.
-This WCF resource generates a random bit for both parties, allowing the cheating party to learn the bit in advance and bias it towards their preferred direction.
+In Theorem 5, we demonstrate the impossibility of constructing a typical WCF resource (see Definition 4) analogous to the widely-used SCF resource [30] from noiseless communication channels in the AC framework.
+Generally speaking, a resource is the abstraction of certain cryptographic functionality (see Section 2.1).
+This WCF resource generates a random bit for both parties, allowing the cheating party to learn the bit in advance and bias it towards their preferred choice.
 It was shown that the widely-used SCF resource cannot be achieved in the AC framework [31], but the result for the similar WCF resource remains unknown until this work.
@@
 Theorem 1 (Informal).
-The typical black-box WCF resource with composable security cannot be constructed from noiseless communication channels in the AC framework.
+The typical WCF resource in Definition 4 with composable security cannot be constructed from noiseless communication channels in the AC framework.
 To prove Theorem 5, we show that WCF protocols using noiseless communication channels are vulnerable to the man-in-the-middle attack when composed in parallel.
-As a result, the typical black-box WCF resource cannot be constructed by any WCF protocol from noiseless communication channels in the AC framework.
-Theorem 5 shows the impossibility of the typical black-box WCF resource in the AC framework.
-However, we can still define an alternative WCF resource at the cost of increasing the interaction complexity.
+As a result, the typical WCF resource cannot be constructed by any WCF protocol from noiseless communication channels in the AC framework.
+While Theorem 5 shows the impossibility of the typical WCF resource in the AC framework, we can still define an alternative WCF resource at the cost of increasing the interaction complexity.
 Nevertheless, such an alternative WCF resource that can be constructed from noiseless communication channels must be extremely complicated in terms of its interaction complexity, as is shown in Theorem 6.
@@
 Theorem 2 (Informal).
-Any black-box WCF resource with composable security constructed from noiseless communication channels in the AC framework must have high interaction complexity.
+Any WCF resource with composable security constructed from noiseless communication channels in the AC framework must have high interaction complexity.
 The proof of Theorem 6 consists of two steps.
 First, we show that an alternative WCF resource with composable security constructed from noiseless communication channels implies a WCF protocol with the same interaction complexity as the alternative WCF resource.
-Then together with the exponential lower bound on the interaction complexity of WCF protocols, we conclude that a WCF resource with a low interaction complexity can not exist.
-Next, we extend the classical version of the sequential composition theorem [32, Theorem 7.4.3] to the quantum case.
+Then, together with the exponential lower bound on the interaction complexity of WCF protocols [32], we conclude that a WCF resource with a low interaction complexity cannot exist.
+Next, we extend the classical version of the composition theorem for sequentially ordered protocols [33, Theorem 7.4.3] to the quantum case.
 We find a sufficient condition called global security where WCF and other modules can be composed into a larger system with stand-alone security.
@@
 Theorem 3 (Informal).
-A primitive constructed from WCF with stand-alone security under strict sequential composition maintains its stand-alone security.
+A primitive constructed from WCF with stand-alone security under strictly ordered sequential composition maintains its stand-alone security.
 The idea to prove Theorem 10 is by contradiction.
-Assume that there is some correlation between outcomes of WCF with stand-alone security and other components under sequential composition.
+Assume that there is some correlation between outcomes of WCF with stand-alone security and other components under ordered sequential composition.
 Using the information carried from previous components to WCF, we can construct an attack breaking the stand-alone security of WCF.
@@
 Causality in the composable framework.
-Some early works [33, 34] developed the paradigm that only allows for sequential composition, where protocols are executed one after another.
-Canetti [3] relaxed “sequential" composition to modular composition in the UC framework, where the former is a special case of the latter.
-The AC framework [9, 10] takes a top-down hierarchy.
+Some early works [34, 35] developed the paradigm that only allows for sequential-order composition, where protocols are executed one after another.
+Canetti [3] relaxed sequential composition to modular composition in the UC framework, where the former is a special case of the latter.
+The AC framework [9, 10] takes a top-down approach.
 Only the connection of abstract systems is considered at the top level, and the execution order is left at the lower level specification.
-Under the AC framework, the causal box model [35] provides the causal order specification that allows for quantum superposition of different orders.
+Under the AC framework, the causal box model [36] provides the causal order specification that allows for a quantum superposition of different message orders.
 Concerning concurrent composition.
-Literature [24, 25] addresses the difficulty of concurrent parallel composition in the UC framework and the bounded storage model.
+Literature [24, 25] addresses the difficulty of concurrent parallel composition in the UC framework and the bounded-storage model.
 The composability of BC was discussed in [37].
 They define the weak binding security of BC in a stand-alone rather than composable manner and propose a counterexample where the string commitment constructed from multiple instances of BC is insecure.
-In a previous model of general cryptographic protocols [32, Section 7.3], concurrent calls to an oracle (resource) are also forbidden.
-This restriction primarily stems from the formulation of syntax in that model rather than an explicit security concern.
+In a previous model of general cryptographic protocols [33, Section 7.3], concurrent calls to an oracle (resource) are also forbidden.
+This restriction primarily stems from the formulation of syntax in that model, rather than an explicit security concern.
+Recent work by Batra et al.
+[38] provides device-independent protocols for oblivious transfer and bit commitment proven secure under sequential composition, where all queries to the ideal functionality must be performed sequentially, but their concurrent composable security remains unknown.
 Coin flipping within the AC framework.
@@
 The building blocks of the framework are resources, converters (protocols), and distinguishers.
-Resources are cryptographic functionalities, and converters transform one resource into another.
+Resources are interactive systems that abstract cryptographic functionalities, and converters transform one resource into another.
 Distinguishers distinguish resources from each other.
-The framework is feasible to describe multiparty functionalities, but we only focus on two-party ones here.
+The framework can describe multiparty functionalities, but in this work, we only focus on two-party ones.
   • A resource, denoted by $R$, has one interface accessible to each party.
 Each party can send inputs, receive outputs, perform operations and give instructions via the corresponding interface.
-In two-party cryptography, it is necessary to consider three different behaviours $R$, $R_{A}$ and $R_{B}$ of a resource, where $R$, $R_{A}$ and $R_{B}$ correspond to the case where both are honest, where Alice is dishonest, and where Bob is dishonest, respectively.
-We use $\mathcal{R}=(R,R_{A},R_{B})$ to denote a resource compactly.
-  • A converter, denoted by a lower-case Greek letter $\alpha$, is a cryptographic system with an inner interface connected to an interface of a resource and an outer interface to a party.
+In two-party cryptography, it is necessary to consider three different behaviors $R$, $R_{A}$ and $R_{B}$ of a resource, in which $R$, $R_{A}$ and $R_{B}$ correspond to the case where both are honest, where only Alice is dishonest, and where only Bob is dishonest, respectively.
+We use $\mathcal{R}=(R,R_{A},R_{B})$ to compactly denote a resource.
+  • A converter, denoted by $\alpha$, is a cryptographic system with an inner interface that connects to a resource and an outer interface that connects to a player.
 The set of all converters is denoted by $\Sigma$.
@@
   • A protocol $\pi=(\pi_{A},\pi_{B})$ consists of Alice’s converter $\pi_{A}$ and Bob’s converter $\pi_{B}$.
-$\pi_{A}$ and $\pi_{B}$ are operations that honest Alice and Bob should follow when building one resource from another.
-By connecting each converter to the corresponding interface, the protocol $\pi$ can build a required resource $S$ from a given resource $R$ if $S=\pi_{A}R\pi_{B}$.
+By connecting each converter to the corresponding interface of a resource $R$, the protocol $\pi$ can transform $R$ into the required resource $S$, provided that $S=\pi_{A}R\pi_{B}$.
 For example, suppose $R$ is a binary symmetric channel resource and $(\pi_{A}$,$\pi_{B})$ is an encoder-decoder pair corresponding to a proper error-correcting code.
@@
   • A distinguisher, denoted by $\mathcal{D}$, is an agent who can access both interfaces of a resource.
-Given one of two resources $R$ and $S$ uniformly at random, a distinguisher guesses whether the resource is $R$ or $S$ by interacting with it.
+Given one of two resources $R$ and $S$ uniformly at random, a distinguisher distinguishes whether the resource is $R$ or $S$ by interacting with it on both interfaces.
 The distinguishing advantage of the above task is defined as
     $d(R,S)=\sup_{\mathcal{D}}|\Pr[\mathcal{D}(R)=0]-\Pr[\mathcal{D}(S)=0]|.$ (1)
-    The distinguishing advantage is a pseudo metric between two resources, satisfying the symmetry and triangle inequality properties. When $d(R,S)\leq\delta$, we also write $R\approx_{\delta}S$.
+    The distinguishing advantage is a pseudo metric between two resources, satisfying the symmetry and triangle inequality properties. If $d(R,S)\leq\delta$, we also write $R\approx_{\delta}S$.
 With these building blocks, we define composable security for two-party functionalities.
 Definition 1 (Composable security).
-A protocol $\pi=(\pi_{A},\pi_{B})$ constructs resource $\mathcal{S}=(S,S_{A},S_{B})$ out of resource $\mathcal{R}=(R,R_{A},R_{B})$ with distinguishing advantage $\delta$ if the following conditions are satisfied
+A protocol $\pi=(\pi_{A},\pi_{B})$ constructs resource $\mathcal{S}=(S,S_{A},S_{B})$ out of resource $\mathcal{R}=(R,R_{A},R_{B})$ with distinguishing advantage $\delta$ if the following conditions are satisfied:
   1. 1.
     Correctness
-    $S\approx_{\delta}\pi_{A}R\pi_{B}.$ (2)
+    $S\approx_{\delta}\pi_{A}R\pi_{B};$ (2)
   2. 2.
     Security against dishonest Bob
-    $\exists\sigma_{B}\in\Sigma,S_{B}\sigma_{B}\approx_{\delta}\pi_{A}R_{B}.$ (3)
+    $\exists\sigma_{B}\in\Sigma,S_{B}\sigma_{B}\approx_{\delta}\pi_{A}R_{B};$ (3)
   3. 3.
@@
 Alternatively, we say that $\mathcal{S}$ is an abstraction of $\pi$ based on $\mathcal{R}$, or that $\mathcal{S}$ abstracts $\pi$ based on $\mathcal{R}$.
-In the previous definition, $\sigma_{A}$ and $\sigma_{B}$ are called simulators.
-Intuitively, the simulators here ensure that any valid cheating strategy for resource $\mathcal{R}$ could be mapped onto some cheating strategy for resource $\mathcal{S}$.
-For instance, if dishonest Bob deviates from the protocol $\pi_{B}$ by following $\pi^{\prime}_{B}$ instead, then, according to Eq.
-(3), this is equivalent to the action $\sigma_{B}\pi^{\prime}_{B}$ on resource $S_{B}$.
+In Definition 1, $\sigma_{A}$ and $\sigma_{B}$ are called simulators.
+These simulators enforce a critical security property: any cheating strategy against the real resource $\mathcal{R}$ can be translated via the simulators into a corresponding cheating strategy against the ideal resource $\mathcal{S}$.
+For instance, if dishonest Bob deviates from $\pi_{B}$ and follows $\pi^{\prime}_{B}$ instead, then, we have $S_{B}\sigma_{B}\pi^{\prime}_{B}\approx_{\delta}\pi_{A}R_{B}\pi_{B}^{\prime}$ by Equation 3, which means Bob’s strategy $\pi_{B}^{\prime}$ in the real world (interacting with $\pi_{A}R_{B}$) is simulated by the $\sigma_{B}\pi^{\prime}_{B}$ in the ideal world (interacting with $S_{B}$).
 For a detailed description of simulators and the related security notions in cryptography, we refer to the tutorial [39].
@@
 2.2 The quantum comb model
-A quantum comb [38, 39] is a general quantum information processing system with internal memory that processes the incoming messages one by one.
-An example of quantum comb is shown in Fig.
-1.
-It is common to model the building blocks in the AC framework as quantum comb.
-If a protocol can be modeled with the quantum comb model (QCM), it is called a QCM protocol.
-Note that all existing WCF protocols can be classified as QCM protocols.
+A quantum comb [40, 41] is a general quantum information processing system with internal memory that processes incoming messages sequentially.
+An example of a quantum comb is shown in Figure 1.
+At the lower level specification of the AC framework where the execution order is taken into account, building blocks are commonly modeled as quantum combs.
+A protocol which can be modeled within the quantum comb model (QCM) is called a QCM protocol.
+All existing WCF protocols can be classified as QCM protocols.
 $\mathcal{E}_{1}$$\mathcal{E}_{2}$$\mathcal{E}_{n}$
-Figure 1: An example of quantum comb. $\mathcal{E}_{i}$ is a quantum operation. The lines with an arrow are quantum states. The unconnected lines are pairs of input and output messages.
-The interaction complexity of a QCM protocol $\pi$, denoted by $\mathsf{Intcom}(\pi)$, is defined as its communication rounds.
-As an example, for a protocol $\pi$ modeled by Fig.
-1, $\mathsf{Intcom}(\pi)=n$.
-For resources such as $S_{A}$ and $S_{B}$, their interaction complexity refers to the communication rounds at the adversary’s side.
-For example, the interaction complexity of WCF resource $S_{A}$ in Fig 3(b) is the rounds at the left side, which is $1$.
+Figure 1: An example of quantum comb.
+The lines represent quantum registers, and $\mathcal{E}_{i}$ is a quantum operation.
+Each quantum operation interacts with the environment through an input-output pair, resembling the teeth of a comb.
+The interaction complexity of a QCM protocol $\pi$, denoted by $\mathsf{Intcom}(\pi)$, is defined as the communication rounds it requires.
+As an example, for a protocol $\pi$ modeled by Figure 1, $\mathsf{Intcom}(\pi)=n$.
+For resources $S_{A}$ and $S_{B}$, their interaction complexity refers to the communication rounds at the adversary’s side.
+For example, the interaction complexity of WCF resource $S_{A}$ in Figure 3(b) is the number of rounds on the left side, which is $1$.
 2.3 The causal box model
 The causal box model [36] is a more general specification of building blocks in the AC framework.
-The causal box model presents the causal structure of the messages sent or received by these building blocks.
+The causal box model describes the causal structure of the messages exchanged by these building blocks.
 The quantum comb model can be treated as a special case of the causal box model.
@@
 A causal box is an information processing system with input and output wires.
-Each message on a wire is described by a quantum state on space $\mathcal{H}\otimes l^{2}(\mathcal{T})$, where $\mathcal{T}$ is a partially ordered set (poset) corresponding to the order of messages and $l^{2}(\mathcal{T})$ is the set of all functions $\mathcal{T}\to\mathbb{C}$ with bounded $L^{2}$ norm.
-Note that $\mathcal{T}$ cannot be simply interpreted as time, which is totally ordered.
-For generality, a wire may contain arbitrarily many messages.
-Therefore, the wire space is defined as the Fock space of the message space
+Each message on a wire is described by a quantum state on space $\mathcal{H}\otimes l^{2}(\mathcal{T})$, where $\mathcal{T}$ is a partially ordered set (poset) that defines the message ordering, and $l^{2}(\mathcal{T})$ is the set of all functions $\mathcal{T}\to\mathbb{C}$ with bounded $L^{2}$-norm.
+Note that $\mathcal{T}$ is not necessarily interpreted as time, as time is totally ordered.
+In general, a wire may carry multiple messages.
+Therefore, the wire space is defined as the Fock space of the message space,
 $\mathcal{F}_{\mathcal{H}}^{\mathcal{T}}\coloneqq\bigoplus_{n=0}^{\infty}\vee^{n}\left(\mathcal{H}\otimes l^{2}(\mathcal{T})\right),$ (8)
-where $\vee^{n}\mathcal{H}$ denotes the symmetric subspace of $\mathcal{H}^{\otimes n}$ and $l^{2}(\mathcal{T})$ is the space of square summable sequence on $\mathcal{T}$.
-Sometimes we abuse the notation by writing $\mathcal{F}_{\mathcal{H}_{A}}^{\mathcal{T}}$ as $\mathcal{F}_{A}^{\mathcal{T}}$.
+where $\vee^{n}\mathcal{H}$ denotes the symmetric subspace of $\mathcal{H}^{\otimes n}$.
+Here, $\mathcal{F}_{\mathcal{H}}^{\mathcal{T}}$ is also a Hilbert space.
+For convenience, we use the shorthand notation $\mathcal{F}_{A}^{\mathcal{T}}$ for $\mathcal{F}_{\mathcal{H}_{A}}^{\mathcal{T}}$.
 With this definition, wires can be merged or split, i.e., for $\mathcal{H}_{A}=\mathcal{H}_{A_{1}}\oplus\mathcal{H}_{A_{2}}$,
 $\mathcal{F}_{A}^{\mathcal{T}}=\mathcal{F}_{A_{1}}^{\mathcal{T}}\otimes\mathcal{F}_{A_{2}}^{\mathcal{T}}.$ (9)
-A causal box is formally defined as a completely positive and trace-preserving (CPTP) map from the input wire $X$ to the output wire $X^{\prime}$.
+Formally, a causal box is defined as a completely positive and trace-preserving (CPTP) map that transforms the input wires to the output wires.
 This map should further respect the causality condition, i.e., the output message at $t^{\prime}\in\mathcal{T}$ only depends on the input messages at $t\in\mathcal{T}$ where $t\preceq t^{\prime}$.
@@
 The composition of causal boxes consists of parallel composition and loops.
-The parallel composition of causal boxes $\Phi$ and $\Psi$, $\Phi\|\Psi$ is defined as the map $\Phi\otimes\Psi$.
-Due to Equation (9), this map’s input and output space are still valid wires.
+The parallel composition of causal boxes $\Phi$ and $\Psi$, denoted by $\Phi\|\Psi$, is defined as the map $\Phi\otimes\Psi$.
+By Equation 9, the input and output spaces of $\Phi\otimes\Psi$ remain valid wires.
 A loop connects segments with the same dimension of the input and output wires.
-An interface labelled by $x$ of a resource or a converter $\Phi$, denoted by $\mathsf{int}_{x}(\Phi)$, is specified by a set of wires, possibly including both input and output wires.
+An interface labeled by $x$ for a causal box $\Phi$ is denoted by $\mathsf{int}_{x}(\Phi)$, which consists of a set of wires, possibly including both input and output wires.
 For example, the inner and outer interfaces of a converter $\alpha$ are denoted by $\mathsf{int}_{in}(\alpha)$ and $\mathsf{int}_{out}(\alpha)$, respectively.
-Two interfaces $\mathsf{int}_{a}(\Phi),\mathsf{int}_{b}(\Psi)$ are compatible if there exists a bijection $P:\mathsf{int}_{a}(\Phi)\leftrightarrow\mathsf{int}_{b}(\Psi)$ such that each pair of wires consists of an input wire and an output wire with the same dimension.
+Two interfaces $\mathsf{int}_{a}(\Phi)$ and $\mathsf{int}_{b}(\Psi)$ are compatible if there exists a bijection $P:\mathsf{int}_{a}(\Phi)\leftrightarrow\mathsf{int}_{b}(\Psi)$ such that each input wire of $\mathsf{int}_{a}(\Phi)$ corresponds to an output wire of $\mathsf{int}_{b}(\Psi)$ with the same dimension and vice versa.
 2.4 Weak coin flipping
-Here we formulate WCF protocols with causal boxes.
+Here, we formulate WCF protocols with causal boxes.
 Definition 2 (Weak coin flipping protocol).
-A weak coin flipping protocol using noiseless communication channels is a pair of converters $\pi=(\pi_{A},\pi_{B})$ with compatible inner interfaces.
-$\pi_{A}$ ($\pi_{B}$) has one output wire on the outer interface and only outputs a single classical bit at $t_{a}$ ($t_{b}$).
-The poset $\mathcal{T}$ of the protocol is bounded both from above and below, i.e., $\exists t_{0}\in\mathcal{T}$, such that $\forall t\in\mathcal{T}$, it holds that $t_{0}\preceq t$, $t\preceq t_{a}$ and $t\preceq t_{b}$.
+A weak coin flipping protocol using noiseless communication channels is a pair of converters $\pi=(\pi_{A},\pi_{B})$ with compatible inner interfaces, where $\pi_{A}$ ($\pi_{B}$) has one output wire on the outer interface and only outputs a single classical bit at $t_{a}$ ($t_{b}$).
+The poset $\mathcal{T}$ of the protocol is bounded both from above and below, i.e., $\exists t_{0},t_{a},t_{b}\in\mathcal{T}$, such that $\forall t\in\mathcal{T}$, it holds that $t_{0}\preceq t$, $t\preceq t_{a}$ and $t\preceq t_{b}$.
 $\pi_{B}$$\pi_{A}$($c_{B},t_{b}$)$(c_{A},t_{a})$$|\psi\rangle\in\mathcal{F}_{X}^{\mathcal{T}}$$|\psi^{\prime}\rangle\in\mathcal{F}_{X^{\prime}}^{\mathcal{T}}$$|\phi\rangle\in\mathcal{F}_{Y}^{\mathcal{T}}$$|\phi^{\prime}\rangle\in\mathcal{F}_{Y^{\prime}}^{\mathcal{T}}$
-Figure 2: A general model for coin flipping protocol using noiseless communication channels. The message spaces $X,Y$ and $X^{\prime},Y^{\prime}$ are of the same dimension, respectively.
-In a WCF protocol, we say Alice wins if $c_{A}=c_{B}=0$ while Bob wins if $c_{A}=c_{B}=1$, which means Alice only wants to prevent Bob from biasing $c_{A}$ towards $1$ and Bob only wants to prevent Alice from biasing $c_{B}$ towards 00.
-Notably, in the definition of weak coin flipping, players are not allowed to abort, whereas in strong coin flipping, aborting is required.
-This distinction arises because, in the weak coin flipping scenario, if a player detects malicious behavior, they can simply declare themselves the winner.
+Figure 2: A general model for a coin flipping protocol using noiseless communication channels. The message spaces $X,Y$ and $X^{\prime},Y^{\prime}$ are of the same dimension, respectively.
+In a WCF protocol, we say Alice wins if $c_{B}=0$, while Bob wins if $c_{A}=1$.
+In this case, an honest Bob only needs to prevent a cheating Alice from biasing $c_{B}$ towards 00 and an honest Alice only needs to prevent a cheating Bob from biasing $c_{A}$ towards $1$.
 Definition 3 (Stand-alone security of WCF).
-Let $\epsilon\leq z$ and $z\in[0,\frac{1}{2}]$.
+Let $z\in[0,\frac{1}{2}]$ and $\epsilon\leq z$.
 A coin flipping protocol $\pi=(\pi_{A},\pi_{B})$ is a $z$-unbalanced $\epsilon$-biased WCF protocol (with stand-alone security), denoted by $\mathrm{WCF}(z,\epsilon)$, if the following conditions are satisfied:
@@
   2. (S2)
-    For any converter $\alpha$ compatible with the inner interface with $\pi_{B}$, the composition $\alpha\pi_{B}$ outputs $c_{B}=0$ with a probability of
+    For any converter $\alpha$ compatible with the inner interface of $\pi_{B}$, the composition $\alpha\pi_{B}$ outputs $c_{B}=0$ with a probability of
     $\Pr[c_{B}=0]\leq z+\epsilon.$ (11)
   3. (S3)
-    For any converter $\beta$ compatible with inner interface with $\pi_{A}$, the composition $\pi_{A}\beta$ outputs $c_{A}=1$ with a probability of
+    For any converter $\beta$ compatible with inner interface of $\pi_{A}$, the composition $\pi_{A}\beta$ outputs $c_{A}=1$ with a probability of
     $\Pr[c_{A}=1]\leq 1-z+\epsilon.$ (12)
@@
 The result contains two parts.
-First, we consider a specific WCF resource derived from a simple but widely adapted SCF resource [30] and show that this specific WCF resource cannot be constructed from noiseless communication channel.
-Nevertheless, this result does not exclude the existence of other types of resources that can be constructed with, say, the protocol in [18].
-Then our second result shows that, for any WCF protocol with enough stand-alone security, there does not exist any useful abstraction.
-Usefulness here means that the resource fully expresses the functionality of WCF and its interaction complexity is low.
+First, we consider a specific WCF resource derived from a simple but widely adapted SCF resource [30] and show that this specific WCF resource cannot be constructed from a noiseless communication channel.
+Nevertheless, this result does not exclude the existence of other more complicated resources that can be constructed with, say, the protocol in [18].
+Then our second result shows that, for any WCF protocol with enough stand-alone security, one cannot abstract it into a useful black-box resource, where usefulness means that the resource captures the full functionality of WCF while maintains a low interaction complexity.
 3.1 Impossibility for a specific resource
+We consider a reasonable and simple WCF resource in Figure 3 which is largely motivated from the strong coin flipping resource described in [30, Section III]111 Although the resource had origins in Blum’s protocol [42] that constructs an unfair coin flip (where a party is allowed to abort on observing the outcome), it serves as an abstraction for a large family of SCF protocols.
+When using a similar protocol here, it is equivalent to a biased coin flip if the honest party declares themselves the winner (locally) whenever the dishonest party aborts [30].
+Therefore, it is unnecessary to consider the case where players abort..
+In the WCF task, a dishonest party wants to force only one of the two outcomes.
+To reflect this idea, we model the possible malicious operation as follows: (1) a cheating Alice (Bob) intercepts the random bit $c^{\prime}$ ($c^{\prime\prime}$) which ought to be the honest output in advance and (2) based on $c^{\prime}$ ($c^{\prime\prime}$), Alice (Bob) sends $b^{\prime}$ and $p^{\prime}$ ($b^{\prime\prime}$ and $p^{\prime\prime}$) to the resource to either resign by setting $p^{\prime}=1$ ($p^{\prime\prime}=1$) or replace $c^{\prime}$ with $b$ with probability $\frac{\epsilon}{1-z}$ ($\frac{\epsilon}{z}$) by setting $p^{\prime}=0$ ($p^{\prime\prime}=0$).
+This establishes Definition 4 as a tuple of ideal WCF resources that allows for a maximum bias of $\epsilon$ for different scenarios.
+$S$$(c,t_{a})$$(c,t_{b})$$S_{A}$$(c^{\prime},t_{0}^{\prime})$$(c_{B},t_{b})$$((b^{\prime},p^{\prime}),t_{1}^{\prime})$$S_{B}$$(c^{\prime\prime},t_{0}^{\prime\prime})$$(c_{A},t_{a})$$((b^{\prime\prime},p^{\prime\prime}),t_{2}^{\prime\prime})$(a)(b)(c)
+Figure 3: (a), (b) and (c) respectively show the components $S,S_{A},S_{B}$ of the WCF resource $\mathcal{S}$. Every message is assigned an element of the order set to indicate the causal order.
 Definition 4 ($z$-unbalanced $\epsilon$-biased WCF resource).
-Let $\epsilon\in[0,z]$ and $z\in[0,\frac{1}{2}]$. A $z$-unbalanced $\epsilon$-biased WCF recourse is characterized by a tuple $\mathcal{S}=(S,S_{A},S_{B})$, as is shown in Figure 3.
+Let $z\in[0,\frac{1}{2}]$ and $\epsilon\in[0,z]$. A $z$-unbalanced $\epsilon$-biased WCF resource is characterized by a tuple $\mathcal{S}=(S,S_{A},S_{B})$, as is shown in Figure 3.
   1. 1.
-    $S$ describes the case when both Alice and Bob are honest. In this case, both party obtains a random bit $c$ satisfying
+    $S$ describes the case when both Alice and Bob are honest. In this case, both parties obtains a random bit $c$ according to the probabilities
     $\Pr[c=0]=z,\quad\Pr[c=1]=1-z.$ (13)
@@
     $S_{A}$ describes the case when Alice is dishonest and Bob is honest.
-$S_{A}$ outputs a random bit $c^{\prime}$ subject to the distribution $(z,1-z)$ at $t_{0}^{\prime}$ and receives two bits $b^{\prime},p^{\prime}$ at $t_{1}^{\prime}$ at the left interface.
-Then it outputs a bit $c_{A}$ at the right interface.
-If $p^{\prime}=1$, then $c_{B}=1$, else the probability distribution of $c_{B}$ follows
-    $\Pr[c_{B}=b^{\prime}]=\frac{\epsilon}{1-z},\quad\Pr[c_{B}=c^{\prime}]=1-\frac{\epsilon}{1-z}.$ (14)
+$S_{A}$ outputs a random bit $c^{\prime}$ subject to the probability distribution $(z,1-z)$ at $t_{0}^{\prime}$ and receives two bits $b^{\prime},p^{\prime}$ at $t_{1}^{\prime}$ at the left interface.
+Then it outputs a bit $c_{B}$ at $t_{b}$ at the right interface.
+If $p^{\prime}=1$, then $c_{B}=1$, else $c_{B}=b^{\prime}$ with probability $\frac{\epsilon}{1-z}$ and $c_{B}=c^{\prime}$ with probability $1-\frac{\epsilon}{1-z}$.
 The causality condition is $t_{0}^{\prime}\prec t_{1}^{\prime}\prec t_{b}$.
   3. 3.
-    $S_{B}$ works similarly as $S_{A}$. The causality condition is $t_{0}^{\prime\prime}\prec t_{1}^{\prime\prime}\prec t_{a}$.
-$S$$(c,t_{a})$$(c,t_{b})$$S_{A}$$(c^{\prime},t_{0}^{\prime})$$(c_{B},t_{b})$$(b^{\prime},t_{1}^{\prime})$$S_{B}$$(c^{\prime\prime},t_{0}^{\prime\prime})$$(c_{A},t_{a})$$(b^{\prime\prime},t_{1}^{\prime\prime})$$(p^{\prime},t_{1}^{\prime})$$(p^{\prime\prime},t_{1}^{\prime\prime})$(a)(b)(c)
-Figure 3: (a), (b) and (c) respectively show the components $S,S_{A},S_{B}$ of the WCF resource $\mathcal{S}$. Every message is assigned an element of the order set to indicate the causal order.
-The resource in Definition 4 is a reasonable and simple WCF resource. The following theorem indicates that it cannot be constructed from noiseless communication channel.
+    $S_{B}$ describes the case when Alice is honest and Bob is dishonest.
+$S_{B}$ outputs a random bit $c^{\prime\prime}$ subject to the probability distribution $(z,1-z)$ at $t_{0}^{\prime\prime}$ and receives two bits $b^{\prime\prime},p^{\prime\prime}$ controlled by Bob at $t_{1}^{\prime\prime}$ at the right interface.
+Then it outputs a bit $c_{A}$ at $t_{a}$ at the left interface.
+If $p^{\prime\prime}=1$, then $c_{A}=0$, else $c_{A}=b^{\prime\prime}$ with probability $\frac{\epsilon}{z}$ and $c_{A}=c^{\prime\prime}$ with probability $1-\frac{\epsilon}{z}$.
+The causality condition is $t_{0}^{\prime\prime}\prec t_{1}^{\prime\prime}\prec t_{a}$.
+To elaborate on the functionality when one party is malicious, consider an example where cheating Alice tries to bias the output $c_{B}$ of $S_{A}$ towards 00.
+For this purpose, she needs to set $p^{\prime}=0$ regardless of the value of $c^{\prime}$.
+Then the probability of her successfully forcing $c_{B}=0$ is
+$\Pr[c_{B}=0]=\Pr[c^{\prime}=0]\cdot\Pr[c_{B}=0|c^{\prime}=0]+\Pr[c^{\prime}=1]\cdot\Pr[c_{B}=0|c^{\prime}=1].$
+As $p^{\prime}=0$, the optimal strategy for Alice is to set $b^{\prime}=0$ since $c_{B}$ is either $c^{\prime}$ or $b^{\prime}$.
+If $c^{\prime}=0$, which happens with probability $\Pr[c^{\prime}=0]=z$, then $c_{B}=0$ for certain, i.e., $\Pr[c_{B}=0|c^{\prime}=0]=1$.
+If $c^{\prime}=1$, which happens with probability $\Pr[c^{\prime}=1]=1-z$, Alice’s conditional winning probability equals the probability that $c_{B}$ takes the value of $b^{\prime}$, i.e., $\Pr[c_{B}=0|c^{\prime}=1]=\frac{\epsilon}{1-z}$.
+We thus have
+$\Pr[c_{B}=0]=\Pr[c^{\prime}=0]\cdot 1+\Pr[c^{\prime}=1]\cdot\frac{\epsilon}{1-z}=z+\epsilon.$ (14)
+Similar arguments also hold for cheating Bob.
+The following theorem indicates that the resource in Definition 4 cannot be constructed from noiseless communication channel.
 Theorem 5.
-There does not exist a protocol $\pi=(\pi_{A},\pi_{B})$ which can construct an $z$-unbalanced $\epsilon$-biased WCF resource $\mathcal{S}=(S,S_{A},S_{B})$ with a distinguishing advantage $\delta<\frac{1}{3}z(1-z-\epsilon)$.
+For $z\in[0,\frac{1}{2}]$ and $\epsilon\in[0,z]$, there does not exist a protocol $\pi=(\pi_{A},\pi_{B})$ which can construct an $z$-unbalanced $\epsilon$-biased WCF resource $\mathcal{S}=(S,S_{A},S_{B})$ with a distinguishing advantage $\delta<\frac{1}{3}\left(z(1-z)-\epsilon+\epsilon\min\{2\epsilon,z\}\right)$.
 Proof.
-Suppose that there is a protocol $\pi=(\pi_{A},\pi_{B})$ that constructs an ideal $z$-unbalanced $\epsilon$-biased WCF resource $\mathcal{S}=(S,S_{A},S_{B})$ with a distinguishing advantage $\delta$ (without loss of generality we assume $z\in[0,\frac{1}{2}]$).
-Combining the security conditions for Alice and Bob, respectively, followed by the triangle inequality, we obtain
+Suppose that there exists a protocol $\pi=(\pi_{A},\pi_{B})$ that constructs an ideal $z$-unbalanced $\epsilon$-biased WCF resource $\mathcal{S}=(S,S_{A},S_{B})$ with a distinguishing advantage $\delta$ (without loss of generality, we assume $z\in[0,\frac{1}{2}]$).
+Combining the security conditions for Alice (6) and Bob (7) using triangle inequality, we obtain
 $\exists\sigma\in\Sigma,S_{B}\sigma S_{A}\approx_{2\delta}\pi_{A}\pi_{B},$ (15)
-where $\sigma=\sigma_{B}\sigma_{A}$. Applying the correctness followed again by the triangle inequality, we proceed to
-$\exists\sigma\in\Sigma,S_{B}\sigma S_{A}\approx_{3\delta}S.$ (16)
+where $\sigma=\sigma_{B}\sigma_{A}$. Once again, combining the previous equation with the correctness condition (5) using triangle inequality, we get
+$\exists\sigma\in\Sigma,S_{B}\sigma S_{A}\approx_{3\delta}S,$ (16)
+or equivalently,
+$\exists\sigma\in\Sigma,d(S_{B}\sigma S_{A},S)\leq 3\delta.$ (17)
 This relation is illustrated in Figure 4.
 […]
-Figure 4: The composition of two ideal WCF resources.
-We now show that, however, there does not exist $\sigma$ such that $\delta<\frac{1}{12}$.
-Note that the best choice of causal order for $\sigma$ is $t_{0}^{\prime}\prec t_{1}^{\prime},t_{0}^{\prime}\prec t_{1}^{\prime\prime},t_{0}^{\prime\prime}\prec t_{1}^{\prime},t_{0}^{\prime\prime}\prec t_{1}^{\prime\prime}$ since it can maximize the use of information of $c^{\prime},c^{\prime\prime}$.
-The most general operation of $\sigma$ is to apply a probabilistic map $\mathcal{M}_{\sigma}:(c^{\prime},c^{\prime\prime})\mapsto(b^{\prime},b^{\prime\prime},p^{\prime},p^{\prime\prime})$.
-Because $S$ must output $c_{A}=c_{B}=c$ while $S_{B}\sigma S_{A}$ might output $c_{A}\neq c_{B}$, a distinguisher $\mathcal{D}_{0}$ which only checks if $c_{A}=c_{B}$ to distinguish $S$ and $S_{B}\sigma S_{A}$ will provide a lower bound to the distinguishing advantage.
-That is, $\mathcal{D}_{0}(\cdot)$ outputs $1$ if $c_{A}=c_{B}$ and else 00.
-Consider the following map on three cases of the input.
+Figure 4: Illustration for Inequality (17).
+We next establish a constant lower bound on such $\delta$ (constant in terms of $z$) by deriving a lower bound on the distinguishing advantage $d(S_{B}\sigma S_{A},S)$ for any $\sigma$, i.e., $\min_{\sigma}d(S_{B}\sigma S_{A},S)$.
+As the two systems $S_{B}\sigma S_{A}$ and $S$ in Figure 4 do not need any external input, $d(S_{B}\sigma S_{A},S)$ is simply the total variation distance between their output distributions, denoted by $P_{c_{A}c_{B}}$ and $P_{cc}$ respectively, i.e.,
+$d(S_{B}\sigma S_{A},S)=\frac{1}{2}\|P_{c_{A}c_{B}}-P_{cc}\|_{1},$ (18)
+with $P_{cc}(0,0)=z,P_{cc}(1,1)=1-z$.
+The best choice of causal order for $\sigma$ is $t_{0}^{\prime}\prec t_{1}^{\prime},t_{0}^{\prime}\prec t_{1}^{\prime\prime},t_{0}^{\prime\prime}\prec t_{1}^{\prime},t_{0}^{\prime\prime}\prec t_{1}^{\prime\prime}$ since it can maximize the use of information of $c^{\prime},c^{\prime\prime}$.
+Then, the minimum distinguishing advantage can be bounded as follows:
+$\min_{\sigma}d(S_{B}\sigma S_{A},S)$ $=\min_{\sigma}\frac{1}{2}\left(\Pr[c_{A}=0,c_{B}=1]+\Pr[c_{A}=1,c_{B}=0]\right.$ (19)
+$\quad\quad+\left.\big{\lvert}\Pr[c_{A}=0,c_{B}=0]-z\big{\rvert}+\big{\lvert}\Pr[c_{A}=1,c_{B}=1]-1+z\big{\rvert}\right)$
+$\geq\min_{\sigma}\left(\Pr[c_{A}=0,c_{B}=1]+\Pr[c_{A}=1,c_{B}=0]\right)$
+$=\min_{\sigma}\Pr[c_{A}\neq c_{B}]$
+$\geq\min_{\sigma}\Pr[c^{\prime}=1,c^{\prime\prime}=0]\cdot\Pr[c_{A}\neq c_{B}|c^{\prime}=1,c^{\prime\prime}=0]\enspace$
+$\quad+\min_{\sigma}\Pr[c^{\prime}=0,c^{\prime\prime}=1]\cdot\Pr[c_{A}\neq c_{B}|c^{\prime}=0,c^{\prime\prime}=1]\enspace$
+$\quad+\min_{\sigma}\Pr[c^{\prime}=c^{\prime\prime}]\cdot\Pr[c_{A}\neq c_{B}|c^{\prime}=c^{\prime\prime}].$
+In the previous equation, the optimal distribution on $(b^{\prime},p^{\prime},b^{\prime\prime},p^{\prime\prime})$ determines the optimal $\sigma$ for each received value $c^{\prime},c^{\prime\prime}$.
+The optimal distributions for different choices of $c^{\prime}$ and $c^{\prime\prime}$ are discussed next.
   1. 1.
-    $\Pr[c^{\prime}=c^{\prime\prime}=0]=z^{2}$ or $\Pr[c^{\prime}=c^{\prime\prime}=1]=(1-z)^{2}$.
-Because the bits are already the same, the box $\sigma$ outputs $p^{\prime}=0,b^{\prime}=c^{\prime},p^{\prime\prime}=0,b^{\prime\prime}=c^{\prime\prime}$ to keep $c_{A}$ and $c_{B}$ the same.
-We have $c_{A}=c_{B}=0$ or $c_{A}=c_{B}=1$.
+    If $c^{\prime}=c^{\prime\prime}=0$ or $c^{\prime}=c^{\prime\prime}=1$, then the box $\sigma$ outputs $p^{\prime}=0,b^{\prime}=c^{\prime},p^{\prime\prime}=0,b^{\prime\prime}=c^{\prime\prime}$ to keep $c_{A}=c^{\prime\prime}=c^{\prime}=c_{B}$, implying that $\min_{\sigma}\Pr[c_{A}\neq c_{B}|c^{\prime}=c^{\prime\prime}]=0$.
   2. 2.
-    $\Pr[c^{\prime}=0,c^{\prime\prime}=1]=z(1-z)$. The box $\sigma$ outputs $p^{\prime}=1,p^{\prime\prime}=0$ to force the outcome $c_{A}=c_{B}=1$. We have $c_{A}=c_{B}=1$.
+    If $c^{\prime}=0,c^{\prime\prime}=1$, then the box $\sigma$ outputs $p^{\prime}=1,p^{\prime\prime}=0$ to force the outcome $c_{A}=c_{B}=1$, implying that $\min_{\sigma}\Pr[c_{A}\neq c_{B}|c^{\prime}=0,c^{\prime\prime}=1]=0$.
   3. 3.
-    $\Pr[c^{\prime}=1,c^{\prime\prime}=0]=z(1-z)$.
-The box $\sigma$ outputs $b^{\prime}=p^{\prime}=b^{\prime\prime}=p^{\prime\prime}=0$ to force the outcome $c_{A}=c_{B}=0$ with probability $\frac{\epsilon}{1-z}$.
-This strategy of $\sigma$ is obviously the optimal one to maximize distinguishing probability under distinguisher $\mathcal{D}_{0}$, which is
-$\Pr[c_{A}\neq c_{B}]=\Pr[c_{A}=1,c_{B}=0]+\Pr[c_{A}=0,c_{B}=1]=z(1-z-\epsilon).$ (17)
-For any box $\sigma$, we have the lower bound on distinguishing advantage:
-$d(S_{B}\sigma S_{A},S)\geq\left|\Pr[\mathcal{D}_{0}(S)=1]-\Pr[\mathcal{D}_{0}(S_{B}\sigma S_{A})=1]\right|\geq z(1-z-\epsilon).$ (18)
-Together with the relation in Equation (16), we have
-$\delta\geq\frac{1}{3}z(1-z-\epsilon).$ (19)
-which shows that $\pi$ cannot construct $\mathcal{S}$ with a vanishing distinguishing advantage. ∎
-By setting $\epsilon=0,z=\frac{1}{2}$, we conclude that there does not exist a protocol $\pi$ that constructs an ideal balanced unbiased WCF resource $\mathcal{S}$ with a distinguishing advantage $\delta<\frac{1}{12}$.
+    If $c^{\prime}=1,c^{\prime\prime}=0$, then we can assume the box $\sigma$ outputs $p^{\prime}=p^{\prime\prime}=0$ because setting $p^{\prime}=1$ ($p^{\prime\prime}=1$) has the same effect as setting $p^{\prime}=0,b^{\prime}=c^{\prime}$ ($p^{\prime\prime}=0,b^{\prime\prime}=c^{\prime\prime}$).
+Consider the event $c_{A}\neq c_{B}$ conditioned on different cases of $(b^{\prime},b^{\prime\prime})$:
+    $\Pr[c_{A}\neq c_{B}|c^{\prime}=1,c^{\prime\prime}=0,b^{\prime}=0,b^{\prime\prime}=0]=1-\frac{\epsilon}{1-z};$ (20)
+    $\Pr[c_{A}\neq c_{B}|c^{\prime}=1,c^{\prime\prime}=0,b^{\prime}=1,b^{\prime\prime}=0]=1;$
+    $\Pr[c_{A}\neq c_{B}|c^{\prime}=1,c^{\prime\prime}=0,b^{\prime}=0,b^{\prime\prime}=1]=1-\frac{\epsilon(1-2\epsilon)}{z(1-z)};$
+    $\Pr[c_{A}\neq c_{B}|c^{\prime}=1,c^{\prime\prime}=0,b^{\prime}=1,b^{\prime\prime}=1]=1-\frac{\epsilon}{z}.$
+    All possible cheating strategies are in the convex hull of the above four cases, then we have
+    $\min_{\sigma}\Pr[c_{A}\neq c_{B}|c^{\prime}=1,c^{\prime\prime}=0]$ (21)
+    $=$ $\min_{P_{b^{\prime},b^{\prime\prime}}}\sum_{i^{\prime},i^{\prime\prime}\in\{0,1\}}P_{b^{\prime},b^{\prime\prime}}(i^{\prime},i^{\prime\prime})\cdot\Pr[c_{A}\neq c_{B}|c^{\prime}=1,c^{\prime\prime}=0,b^{\prime}=i^{\prime},b^{\prime\prime}=i^{\prime\prime}]$ (22)
+    $=$ $\min_{i^{\prime},i^{\prime\prime}\in\{0,1\}}\Pr[c_{A}\neq c_{B}|c^{\prime}=1,c^{\prime\prime}=0,b^{\prime}=i^{\prime},b^{\prime\prime}=i^{\prime\prime}]$ (23)
+    $=$ $\begin{cases}1-\frac{\epsilon(1-2\epsilon)}{z(1-z)},&\text{if }0\leq\epsilon<z/2\\ 1-\frac{\epsilon}{z},&\text{if }z/2<\epsilon\leq z\end{cases}.$ (24)
+    where the conditional probabilities $\Pr[c_{A}\neq c_{B}|c^{\prime}=1,c^{\prime\prime}=0,b^{\prime}=i^{\prime},b^{\prime\prime}=i^{\prime\prime}]$ are given in Equation 20, and Equation 24 holds because $20\leq\epsilon\leq z\leq 1/$ 2.
+By the above discussion, Equation 19 continues as
+$\min_{\sigma}d(S_{B}\sigma S_{A},S)$ $\geq\min_{\sigma}\Pr[c^{\prime}=1,c^{\prime\prime}=0]\cdot\Pr[c_{A}\neq c_{B}|c^{\prime}=1,c^{\prime\prime}=0]$ (25)
+$=z(1-z)-\epsilon+\epsilon\min\{2\epsilon,z\},$ (26)
+where Equation 25 follows from Equation 19 and the implications of Item 1 and Item 2. Equation 26 simply follows from Equation 24. Finally, combining Equation 17 and Equation 26, we have
+$\delta\geq\frac{1}{3}\left(z(1-z)-\epsilon+\epsilon\min\{2\epsilon,z\}\right),$ (27)
+which implies that $\pi$ cannot construct $\mathcal{S}$ with a vanishing distinguishing advantage. ∎
+Setting $\epsilon=0,z=\frac{1}{2}$, we conclude that there does not exist a protocol $\pi$ that constructs an ideal balanced unbiased WCF resource $\mathcal{S}$ with a distinguishing advantage $\delta<\frac{1}{12}$.
 3.2 Impossibility for a general resource
-The composable security is defined by comparing between the ideal resource and the actual protocol.
-Theorem 5 denies the possibility of constructing an actual protocol close to the ideal WCF resource in Definition 4.
-However, if we modify the definition of the ideal WCF resource into a presumably more complicated one, then there might still exist some protocol that constructs this more complicated resource.
+Composable security is defined by requiring that an actual protocol emulates an ideal resource within a certain accuracy.
+As established in Theorem 5, no actual protocol can approximate the ideal WCF resource defined in Definition 4.
+However, this impossibility result pertains only to this specific ideal WCF resource.
+By redefining the ideal WCF resource, potentially introducing additional structure, a protocol that securely constructs this modified ideal resource may still exist at the cost of a more complicated and less useful ideal WCF resource.
 To address this possibility, it is crucial to determine what properties a WCF resource must possess to be useful in general.
@@
   • Simple.
-By simplicity of resources, we typically refer to the process of abstracting away specific details of a resource that reduces its overall interaction complexity and allows to study it in a more general context.
+Resource simplicity refers to abstraction of non-essential details of a protocol, thereby reducing its interaction complexity.
+This simplification facilitates the analysis of the resource in a more general context.
 Here we take the interaction complexity $\mathsf{Intcom}(S_{A})$, $\mathsf{Intcom}(S_{B})$ (defined in Section 2.2) to quantify it.
+Although the idea of a simple resource is rather subjective, we state a resource to be simple if its net interaction complexity ($\mathsf{Intcom}(S_{A})+\mathsf{Intcom}(S_{B})$) is of $\mathcal{O}(1)$.
 Next, we will show that a fully expressing and simple WCF resource cannot be constructed from noiseless communication channels.
-Specifically, Theorem 6 states that, if a fully expressing resource for $\mathrm{WCF}(z,\epsilon)$ with small $\epsilon$ can be constructed from noiseless communication channels, then its interaction complexity cannot be low.
+Specifically, Theorem 6 states that, if a fully expressing resource for $\mathrm{WCF}(z,\epsilon)$ with small $\epsilon$ can be constructed from noiseless communication channels, then its interaction complexity cannot be small.
 Theorem 6.
@@
 $\min\{\mathsf{Intcom}(S_{A}),\mathsf{Intcom}(S_{B})\}\geq\exp\left(\Omega(1/\sqrt{\epsilon})\right).$ (28)
-To prove Theorem 6, we need a proposition rewritten from [40, Theorem 8.2].
+To prove Theorem 6, we need a proposition rewritten from [32, Theorem 8.2].
 Proposition 7.
@@
 An example is delegated quantum computation (DQC), where the client aims to delegate computation to the server while maintaining privacy.
-The $S_{A}$ (against dishonest client) of DQC resource [41] abstracting the protocol in [42, 43] has no interaction in either side, because it is just trivial to talk about a dishonest client.
-This implies the existence of a simple DQC protocol with no interaction, wherein the client performs all the computation herself.
+The $S_{A}$ (against dishonest client) of DQC resource [43] abstracting the protocol in [44, 45] has no interaction on either side.
+This implies the existence of a simple DQC protocol with no interaction, wherein the client performs all the computation herself, ensuring both correctness and blindness.
 4 Security of weak coin flipping under composition
@@
 This section starts with the limitation of the stand-alone security definition (Definition 3) in the context of composition.
-Specifically, for the construction of unbalanced WCF from a balanced one, we propose a condition (Inequality (26)) which, while implicit, is essential in the security proof of the unbalanced WCF protocol depicted in [29].
-Subsequently, we introduce global security as an extension of the stand-alone definition to encompass possible interdependencies between the outputs of the WCF protocol and other components in a larger system.
+Specifically, for the construction of unbalanced WCF from a balanced one, we propose a condition (Inequality (34)) which, while implicit in [29], is essential in the security proof of the unbalanced WCF protocol depicted there.
+Subsequently, we introduce global security as an extension of the stand-alone definition to encompass possible interdependency between the outputs of the WCF protocol and other components in a larger system.
 Furthermore, we establish the condition under which global security is maintained.
 4.1 Stand-alone security is not sufficient
-For a cryptographic task, some security definitions seem fine from a stand-alone viewpoint but may present some loopholes when composed with other protocols.
+For a cryptographic task, some security definitions seem fine from a stand-alone viewpoint, but may present some loopholes when composed with other protocols.
 For example, the accessible information criteria for quantum key distribution falls short of providing composable security due to the data-locking effect [2, 1].
-This issue is similar to WCF.
-In particular, a WCF protocol $\pi=(\pi_{1},\pi_{2})$ (where $\pi_{1}$ and $\pi_{2}$ are used instead of $\pi_{A}$ and $\pi_{B}$ to avoid confusion in subsequent protocol description) may find its stand-alone security insufficient for the upper-level task.
+A similar problem occurs in WCF.
+In particular, a WCF protocol $\pi=(\pi_{1},\pi_{2})$ (where $\pi_{1}$ and $\pi_{2}$ are used instead of $\pi_{A}$ and $\pi_{B}$ to avoid confusion in subsequent protocol description) may find its stand-alone security insufficient for upper-level tasks.
 To illustrate, consider an attacking strategy for an unbalanced WCF protocol $\mu=(\mu_{A},\mu_{B})$ constructed from two instances of balanced WCF protocol $\pi=(\pi_{1},\pi_{2})$.
@@
 Assume the protocol $\pi$ follows the quantum comb model [41, 40] with $n+1$ (assume that $n$ is odd without loss of generality) rounds of communication, where Alice and Bob send messages alternately.
-Then $\pi_{1}$ and $\pi_{2}$ can be characterized by the ordered set $\mathcal{T}_{\pi}$ shown in Figure 5(b), where the subset $\{t_{i}\}_{i=0}^{n}$ is totally ordered.
+Then, $\pi_{1}$ and $\pi_{2}$ can be characterized by the ordered set $\mathcal{T}_{\pi}$ shown in Figure 5(b), where the subset $\{t_{i}\}_{i=0}^{n}$ is totally ordered.
 The comb model further requires that
@@
     Messages are sent by $\pi_{2}$ and accepted by $\pi_{1}$ only at $t_{i}$ where $i$ is odd.
-The details of these causal boxes are shown in Appendix A.3.
+The above conditions allow us to consider the wires in the two directions separately–one within the even subset of $\mathcal{T}_{\pi}$ and the other within the odd subset of $\mathcal{T}_{\pi}$.
 The protocol $\mu$ is constructed by connecting two independent instances of $\pi$ with the converter $\eta=(\eta_{A},\eta_{B})$, as shown in Figure 5(a).
@@
 As per [29, Lemma 4], if $\pi$ is a $\mathrm{WCF}(\frac{1}{2},\epsilon)$ protocol and $\epsilon<1/6$, $\mu$ is expected to be a $\mathrm{WCF}(\frac{3}{4},\frac{3}{2}\epsilon)$ protocol.
-However, this expectation is not met as there exists a strategy for a dishonest Bob to win the game with probability $\frac{1}{2}$.
+However, it fails to be a $\mathrm{WCF}(\frac{3}{4},\frac{3}{2}\epsilon)$ protocol, as there exists a strategy for dishonest Bob to win the game with probability $\frac{1}{2}$.
 Before elaborating on this strategy, we introduce the delay box, denoted by $\Theta_{f}$.
 This box essentially transforms a message at position $t$ to a future position $t^{\prime}$ according to some delay function $f$.
-For a rigorous definition of delay box, see Appendix A.3.
-Consider the parallel composition of delay boxes $\theta=\theta_{1}\|\theta_{2}$ depicted in Figure 6. The delay functions are defined as
+For a rigorous definition of delay box, see Section A.3.
+For simplicity, denote $\theta_{f_{1}},\theta_{f_{2}}$ as $\theta_{1},\theta_{2}$.
+Consider the parallel composition of delay boxes $\theta=\theta_{1}\|\theta_{2}$ depicted in Figure 6.
+The delay functions are defined as
 $f_{1}:\{t_{i}\in\mathcal{T}_{\pi}\mid\text{$i$ is even}\}\to\mathcal{T}_{\theta_{1}},t_{i}\mapsto t_{i}^{\prime},$ (31)
@@
 (b) This figure shows the Hasse diagram of the ordered sets $\mathcal{T}_{\theta_{1}}$, $\mathcal{T}_{\theta_{2}}$ and $\mathcal{T}_{(\pi_{1}\|\pi_{2}^{\prime})\theta}$.
-Proposition 8 shows that Figure 6(a) emulates a WCF protocol between Alice and herself. See Appendix B.1 for the proof.
+Proposition 8 shows that Figure 6(a) emulates a WCF protocol between Alice and herself. See Section B.1 for the proof.
 Proposition 8.
@@
 $\forall i\in[n],\forall x^{i-1}_{1}\in\{0,1\}^{i-1},\Pr[c_{A,i}=1|c_{A,1}=x_{1},\ldots,c_{A,i-1}=x_{i-1}]\leq\frac{1}{2}+\epsilon,$ (34)
-where $x_{1}^{i}=x_{1}x_{2}\dots x_{i}$ and the result of the $i$th instance is denoted by $c_{A,i}$.
+where $x_{1}^{i}=x_{1}x_{2}\dots x_{i}$ and the result of the $i$-th instance is denoted by $c_{A,i}$.
 Notably, inequality (34) is an implicit but crucial condition for the security of unbalanced WCF protocol introduced in [29].
@@
 More broadly speaking, what is the safe way to compose WCF protocols?
-Despite negative results discussed in Section 3 on the possibility of abstracting WCF as a black-box abstraction, there are still certain ways of composition, satisfying Inequality (26).
+Despite negative results discussed in Section 3 on the possibility of abstracting WCF as a resource, there are still certain secure ways of composition, e.g., those satisfying Inequality (34).
 To explore this, we generalize Inequality (34), introduce the concept of global secrecy and specify the sufficient condition for it.
@@
     $\eta$ only accepts input and produces output on set $\mathcal{T}_{1}\cup\mathcal{T}_{3}$, while $\pi$ only accepts input and produces output on set $\mathcal{T}_{2}$.
-The proof of Theorem 10 is in Appendix B.2.
+The proof of Theorem 10 is in Section B.2.
 The proof idea is by contradiction.
@@
 Therefore, unbalanced WCF protocols play a key role in constructing optimal protocols of more complicated functionalities such as SCF and BC [29, 28].
-An unbalanced $\mathrm{WCF}(z,\epsilon^{\prime})$ protocol can be construct from balanced $\mathrm{WCF}(\frac{1}{2},\epsilon)$ protocols where $z=0.b_{1}b_{2}\dots b_{n}$ is in binary representation, as is shown in Figure 7.
+An unbalanced $\mathrm{WCF}(z,\epsilon^{\prime})$ protocol can be constructed from balanced $\mathrm{WCF}(\frac{1}{2},\epsilon)$ protocols where $z=0.b_{1}b_{2}\dots b_{n}$ is in binary representation, as is shown in Figure 7.
 […]
-While $i<n$
-   If $c_{i}\neq b_{i}$
-      break;
+While
+$i<n$
+If $c_{A,i}\neq b_{i}$ Break;
 Endif
@@
 Output
-$c=c_{i}$
+$c_{A}=c_{A,i}$
 $\mathcal{T}_{2}$$\mathcal{T}_{1}$$\mathcal{T}_{n}$(b)
@@
 The idea is to prove the stand-alone security of $\tau_{i}(\pi_{1}\|...\|\pi_{i})$ constructed from $\tau_{i-1}(\pi_{1}\|...\|\pi_{i-1})$ and $\pi_{i}$ using Theorem 10 recursively.
-The detailed proof can be found in Appendix B.3.
+The detailed proof can be found in Section B.3.
 Corollary 11.
 Let $z\in[0,\frac{1}{2}]$. The unbalanced WCF protocol shown in Figure 7 is a $\mathrm{WCF}(z,\epsilon^{\prime})$ protocol with $\epsilon^{\prime}=2\epsilon+o(\epsilon)$.
+5 Discussion
+Our work demonstrates that the standalone security of protocols does not guarantee the security under certain types of composition.
+In particular, although quantum WCF protocol can be secure in the standalone scenario, it fails to be composably secure in the abstract cryptography framework.
+Moreover, when we construct new protocols with quantum WCF protocol, it remains secure in sequential composition, but exhibits security loopholes in concurrent composition.
+Similar security loopholes also arise in other two-party cryptographic tasks, such as bit commitment [25, 37].
+For these tasks, though we have hoped for simple and good abstractions for actual protocols in the abstract framework, our work suggests that such abstraction does not exist, which makes security analysis of such protocols notably challenging.
 6 Acknowledgement
@@
 $\mathcal{C}=\bigcup_{t\in\mathcal{C}}\mathcal{T}^{\preceq t}.$ (35)
-A cut is bounded if $\exists t\in\mathcal{T},\mathcal{C}\in\mathcal{T}^{\preceq t}$.
+A cut is bounded if $\exists t\in\mathcal{T},\mathcal{C}\subseteq\mathcal{T}^{\preceq t}$.
 The set of all cuts of $\mathcal{T}$ is denoted by $\mathfrak{C}(\mathcal{T})$.
@@
 $\forall\mathcal{C}\in\bar{\mathfrak{C}}(\mathcal{T}),\forall t\in\mathcal{C},\exists n\in\mathbb{N},\quad t\notin\chi^{n}(\mathcal{C})\,,$ (39)
-Using cuts and causality function, we can formally define causal box.
+Using cuts and causality function, we can formally define the causal box.
 Definition 15 (Causal box).
@@
 where $\mathfrak{T}(\mathcal{H})$ denotes the set of all trace class operators on the Hilbert space $\mathcal{H}$. These maps must be mutually consistent
-$\forall\mathcal{C},\mathcal{D}\in\bar{\mathfrak{C}}(\mathcal{T}),\mathcal{C}\subseteq\mathcal{D}:\Phi^{\mathcal{C}}=\mathrm{tr}_{\mathcal{D}\setminus\mathcal{C}}\circ\Phi^{\mathcal{D}}$ (33)
+$\forall\mathcal{C},\mathcal{D}\in\bar{\mathfrak{C}}(\mathcal{T}),\mathcal{C}\subseteq\mathcal{D}:\Phi^{\mathcal{C}}=\mathrm{tr}_{\mathcal{D}\setminus\mathcal{C}}\circ\Phi^{\mathcal{D}},$ (41)
 and respect causality, i.e., there exists a causality function $\chi:\mathfrak{C}(\mathcal{T})\to\mathfrak{C}(\mathcal{T})$ such that
 $\Phi^{\mathcal{C}}=\Phi^{\mathcal{C}}\circ\mathrm{tr}_{\mathcal{T}\setminus\chi(\mathcal{C})}.$ (42)
-Alternatively, the causal box $\Phi$ can be described in Choi-Jamiolkowski (CJ) representation.
+Alternatively, the causal box $\Phi$ can be described in Choi-Jamiołkowski (CJ) representation.
 To be precise, the Choi operator $R_{\Phi}^{\mathcal{C}}$ for the CPTP map $\Phi^{\mathcal{C}}$ is a sesquilinear semi-definite form on $\mathcal{F}_{Y}^{\mathcal{C}}\otimes\mathcal{F}_{X}^{\chi(\mathcal{C})}$ such that
-$R_{\Phi}^{\mathcal{C}}(\psi_{Y}\otimes\psi_{X},\varphi_{Y}\otimes\varphi_{X})=\langle\psi_{Y}|\Phi^{\mathcal{C}}(|\bar{\psi}_{X}\rangle\langle\bar{\varphi}_{X}|)|\varphi_{Y}\rangle.$ (35)
+$R_{\Phi}^{\mathcal{C}}(\psi_{Y}\otimes\psi_{X},\varphi_{Y}\otimes\varphi_{X})=\langle\psi_{Y}|\Phi^{\mathcal{C}}\left(|\bar{\psi}_{X}\rangle\langle\bar{\varphi}_{X}|\right)|\varphi_{Y}\rangle.$ (43)
 where $|\overline{\psi}\rangle=\sum_{i}|i\rangle\langle\psi|i\rangle$ for a fixed basis $\{|i\rangle\}_{i}$.
@@
 Causal boxes can be connected by parallel composition and loop.
-Definition 16 (Parallel composition of causal box).
+Definition 16 (Parallel composition of causal boxes).
 Let $\Phi=\{\Phi^{\mathcal{C}}\}_{\mathcal{C}\in\bar{\mathfrak{C}}(\mathcal{T})}$ and $\Psi=\{\Psi^{\mathcal{C}}\}_{\mathcal{C}\in\bar{\mathfrak{C}}(\mathcal{T})}$ be two causal boxes.
@@
 The above definition of parallel composition implicitly requires the causal boxes $\Phi$ and $\Psi$ share the same partially ordered set $\mathcal{T}$.
-However, a general causal box may not contain the information of the causal structure of some other causal box, that is, there is no such common set $\mathcal{T}$.
+However, a general causal box may not contain the information of the causal structure of some other causal box.
+Then, a predefined common poset $\mathcal{T}$ may not exist.
 In this case, to make the composition possible, we propose a more general version of parallel composition.
@@
 where $\mathcal{C}_{1}=\mathcal{T}_{1}\cap\mathcal{C},\mathcal{C}_{2}=\mathcal{T}_{2}\cap\mathcal{C}$.
-Lemma 1920 and 21 suggest that Definition 18 is well defined.
+Lemma 1920 and 21 suggest that Definition 18 is well-defined.
 Lemma 19.
@@
 Notice that
-$\mathcal{C}_{1}=\mathcal{C}\cap\mathcal{T}_{1}=\bigcup_{t\in\mathcal{C}}(\mathcal{T}^{\preceq t}\cap\mathcal{T}_{1}).$
+$\mathcal{C}_{1}=\mathcal{C}\cap\mathcal{T}_{1}=\bigcup_{t\in\mathcal{C}}(\mathcal{T}^{\preceq t}\cap\mathcal{T}_{1}).$ (48)
 Assume $t^{\prime}\in\mathcal{T}^{\preceq t}\cap\mathcal{T}_{1}$, then by Definition 17, $\forall x\in\mathcal{T}_{1},x\preceq_{1}t^{\prime}\Rightarrow x\preceq t\Rightarrow x\in\mathcal{T}^{\preceq t}\cap\mathcal{T}_{1}$.
@@
 Assume $\chi_{1}$ and $\chi_{2}$ are causality functions for $\Phi$ and $\Psi$, respectively.
-Construct the function $\chi:\mathfrak{C}(\mathcal{T})\to\mathfrak{C}(\mathcal{T})$ such that $\chi(\mathcal{C})=\chi_{1}(\mathcal{C}\cap\mathcal{T}_{1})\cup\chi_{2}(\mathcal{C}\cap\mathcal{T}_{2})$, then it is easy to verify the conditions in Definition 14.
+Construct the function $\chi:\mathfrak{C}(\mathcal{T})\to\mathfrak{C}(\mathcal{T})$ as $\chi(\mathcal{C})=\chi_{1}(\mathcal{C}\cap\mathcal{T}_{1})\cup\chi_{2}(\mathcal{C}\cap\mathcal{T}_{2})$, then it is easy to verify the conditions in Definition 14.
 ∎
 Lemma 21 (Associativity).
-For boxes $\Phi,\Psi,\Theta$ with compatible posets $\mathcal{T}_{\Phi},\mathcal{T}_{\Psi},\mathcal{T}_{\Theta}$,
+For boxes $\Phi,\Psi,\Theta$ with compatible posets $\mathcal{T}_{\Phi},\mathcal{T}_{\Psi},\mathcal{T}_{\Theta}$, we have
 $(\Phi\|\Psi)\|\Theta=\Phi\|(\Psi\|\Theta).$ (50)
@@
 Let $\Phi=\{\Phi^{\mathcal{C}}:\mathfrak{T}(\mathcal{F}_{AB}^{\mathcal{T}})\to\mathfrak{T}(\mathcal{F}_{CD}^{\mathcal{C}})\}_{\mathcal{C}\in\bar{\mathfrak{C}}(\mathcal{T})}$ be a causal box such that $\dim\mathcal{H}_{B}=\dim\mathcal{H}_{C}$ is its CJ representation.
-Let $\{|k_{C}\rangle\}$ and $\{|l_{C}\rangle\}$ be any orthonormal bases of $\mathcal{F}_{C}^{\mathcal{C}}$ and $\{|k_{B}\rangle\}$ and $\{|l_{B}\rangle\}$ be corresponding bases of $\mathcal{F}_{C}^{\mathcal{C}}$.
+Let $\{|k_{C}\rangle\}$ and $\{|l_{C}\rangle\}$ be any orthonormal bases of $\mathcal{F}_{C}^{\mathcal{C}}$ and $\{|k_{B}\rangle\}$ and $\{|l_{B}\rangle\}$ be corresponding bases of $\mathcal{F}_{B}^{\mathcal{C}}$.
 The new box by looping from wire $B$ to wire $C$, denoted by $\Psi=\Phi^{C\hookrightarrow B}$, is a set of maps $\{\Psi^{\mathcal{C}}:\mathfrak{T}(\mathcal{F}_{A}^{\mathcal{T}})\to\mathfrak{T}(\mathcal{F}_{D}^{\mathcal{C}})\}$ with CJ representation
@@
 where the conjugate $|\overline{\psi}\rangle=\sum_{i}|i_{B}\rangle\langle\psi|i_{B}\rangle$ is taken with respect to the base $|i_{B}\rangle$ of $\mathcal{F}_{B}^{\mathcal{T}}$ on which the CJ representation is defined.
+Note that, to define a loop, it is necessary to specify two bases for each of the connected wire. With different bases for the wires, the result of the loop can be different.
 A.3 Delay box
@@
 $\theta_{1}$ is a delay box from input wire $C^{\prime}$ to output wire $C^{\prime\prime}$ and is derived from $f_{1}$ defined in Equation 31.
-Then, $U_{f_{1}}^{\mathcal{C}}$ is an isometry between $\mathcal{F}_{C^{\prime}}^{\chi_{f_{1}}(\mathcal{C})}$ and $\mathcal{F}_{C^{\prime\prime}}^{\mathcal{C}\cap\mathcal{T}_{\pi^{\prime}}}$
+Then, $U_{f_{1}}^{\mathcal{C}}$ is an isometry between $\mathcal{F}_{C^{\prime}}^{\chi_{f_{1}}(\mathcal{C})}$ and $\mathcal{F}_{C^{\prime\prime}}^{\mathcal{C}\cap\mathcal{T}_{\pi^{\prime}}}$.
 Appendix B Proofs
 B.1 Proof of Proposition 8
-Since the protocol $\pi$ is restricted to the comb model, we can assume that $\pi_{1}$ only outputs at the set $\mathcal{T}_{\pi}^{\text{even}}=\{t_{i}\in\mathcal{T}_{\pi}\mid i\text{ is even}\}$ and $\pi_{2}$ only outputs at set $\mathcal{T}_{\pi}^{\text{odd}}=\{t_{i}\in\mathcal{T}_{\pi}\mid i\text{ is odd}\}$.
-For simplicity, where the cut $\mathcal{C}$ is obvious, denote for $X=B,C,C^{\prime},C^{\prime\prime},E$
+Proposition.
+The box $(\pi_{1}\|\pi_{2}^{\prime})\theta$ has the same output distribution as $\pi_{1}\pi_{2}$.
+Proposition 8 implies that two parties running $\pi_{1}$ and $\pi_{2}^{\prime}$ with an adversary running $\theta$ is equivalent to two parties running $\pi_{1}$ and $\pi_{2}^{\prime}$ between themselves.
+To achieve this, the adversary performs the man-in-the-middle attack, i.e., uses delay boxes to pass messages from $\pi_{1}$ to $\pi_{2}^{\prime}$ and pass responses from $\pi_{2}^{\prime}$ to $\pi_{1}$ subsequently.
+To be precise, notice that the boxes $\pi_{1}$ and $\pi_{2}^{\prime}$ work on two independent posets, i.e., the top part and bottom part in Figure 8(b), respectively.
+The delay boxes $\theta_{1}$ and $\theta_{2}$ are the bridges that connect the two corresponding wire spaces.
+Then the key point is to find proper wire space isomorphisms for the wire connections such that $\pi_{2}^{\prime}$ together with the delay boxes is a perfect emulation of $\pi_{2}$.
+[…]
+$t_{0}$$t^{\prime}_{0}$$t_{2}$$t^{\prime}_{2}$$t_{n}$$t^{\prime}_{n}$$t_{1}$$t^{\prime}_{1}$$t_{3}$$t^{\prime}_{3}$$\mathcal{T}_{(\pi_{1}\|\pi^{\prime}_{2})\theta}:$
+Figure 8: (a) This figure shows the connected box $(\pi_{1}\|\pi_{2}^{\prime})\theta$.
+The wire with a blue notation indicates that the messages on this wire only stay in sets $\mathcal{T}_{\pi}^{\text{even}}$ and $\mathcal{T}_{\pi^{\prime}}^{\text{even}}$, while the red notation corresponds to the sets $\mathcal{T}_{\pi}^{\text{odd}}$ and $\mathcal{T}_{\pi^{\prime}}^{\text{odd}}$.
+(b) This figure explains the partition of the totally ordered set $\mathcal{T}_{(\pi_{1}\|\pi_{2}^{\prime})\theta}$ with a non-standard Hasse diagram.
+Specifically, The blue elements at the top form the set $\mathcal{T}_{\pi}^{\text{even}}$; The blue elements at the bottom form the set $\mathcal{T}_{\pi^{\prime}}^{\text{even}}$; The red elements at the top form the set $\mathcal{T}_{\pi}^{\text{odd}}$; The red elements at the bottom form the set $\mathcal{T}_{\pi^{\prime}}^{\text{odd}}$.
+Define the sets
+$\mathcal{T}_{\pi}^{\text{even}}$$\coloneqq\{t_{i}\in\mathcal{T}_{\pi}\mid i\text{ is even}\},$ $\mathcal{T}_{\pi}^{\text{odd}}$$\coloneqq\{t_{i}\in\mathcal{T}_{\pi}\mid i\text{ is odd}\},$ (55)
+$\mathcal{T}_{\pi^{\prime}}^{\text{even}}$$\coloneqq\{t_{i}\in\mathcal{T}_{\pi^{\prime}}\mid i\text{ is even}\},$ $\mathcal{T}_{\pi^{\prime}}^{\text{odd}}$$\coloneqq\{t_{i}\in\mathcal{T}_{\pi^{\prime}}\mid i\text{ is odd}\}.$ (56)
+Because the protocol $\pi$ is restricted to the comb model, the wires $B,C,C^{\prime},C^{\prime\prime},E$ are only effective on the set $\mathcal{T}_{\pi}^{\text{even}}$ and $\mathcal{T}_{\pi^{\prime}}^{\text{even}}$.
+Similarly, the wires $A,D,D^{\prime},D^{\prime\prime},F$ are only effective on the set $\mathcal{T}_{\pi}^{\text{odd}}$ and $\mathcal{T}_{\pi^{\prime}}^{\text{odd}}$.
+By saying wire $X$ is only effective on $\mathcal{T}^{\prime}$, we mean the state on this wire is always in the space $\mathcal{F}_{X}^{\mathcal{T}^{\prime}}$.
+For convenience, where the cut $\mathcal{C}$ is obvious, denote for $X=B,C,C^{\prime},C^{\prime\prime},E$
 $\hat{X}$ $=\mathcal{F}_{X}^{\mathcal{C}\cap\mathcal{T}_{\pi}^{\text{even}}},$ (57)
@@
 where $\mathcal{T}_{\pi}$ is the poset for $\pi=(\pi_{1},\pi_{2})$ and $\mathcal{T}_{\pi^{\prime}}$ is the partial order for $\pi^{\prime}=(\pi_{2}^{\prime},\pi_{1}^{\prime})$.
-And denote the wire space of $X=A,D,D^{\prime},D^{\prime\prime},F$ by
+Denote the wire space of $X=A,D,D^{\prime},D^{\prime\prime},F$ by
 $\hat{X}$ $=\mathcal{F}_{X}^{\mathcal{C}\cap\mathcal{T}_{\pi}^{\text{odd}}},$ (59)
 $\tilde{X}$ $=\mathcal{F}_{X}^{\mathcal{C}\cap\mathcal{T}_{\pi^{\prime}}^{\text{odd}}}.$ (60)
-For $\mathcal{C}\in\mathfrak{C}(\mathcal{T}_{\pi})$, $\{|i\rangle_{\hat{A}}\}$ is a orthonormal base of space $\mathcal{F}_{A}^{\chi_{\pi}(\mathcal{C})}$, The Choi-Jamiolkowski (CJ) representation of the box $\pi_{1}$ is a sesquilinear form $R_{\pi_{1}}^{\mathcal{C}}$ satisfying
-$R_{\pi_{1}}^{\mathcal{C}}(\psi_{\hat{C}}\otimes\psi_{\hat{E}}\otimes\psi_{\hat{A}},\varphi_{\hat{C}}\otimes\varphi_{\hat{E}}\otimes\varphi_{\hat{A}})=\langle\psi_{\hat{C}}|\langle\psi_{\hat{E}}|\Pi_{1}^{\mathcal{C}}(|\bar{\psi}_{\hat{A}}\rangle\langle\bar{\varphi}_{\hat{A}}|)|\varphi_{\hat{E}}\rangle|\varphi_{\hat{C}}\rangle,$ (50)
+For any wire $X$, we call the isomorphism between $\hat{X}$ and $\tilde{X}$, obtained by substituting $t_{i}$ with $t_{i}^{\prime}$, the natural isomorphism.
+For $\mathcal{C}\in\mathfrak{C}(\mathcal{T}_{\pi})$, $\{|i\rangle_{\hat{A}}\}$ is an orthonormal base of space $\mathcal{F}_{A}^{\chi_{\pi}(\mathcal{C})}$, The Choi-Jamiołkowski (CJ) representation of the box $\pi_{1}$ is a sesquilinear form $R_{\pi_{1}}^{\mathcal{C}}$ satisfying
+$R_{\pi_{1}}^{\mathcal{C}}(\psi_{\hat{C}}\otimes\psi_{\hat{E}}\otimes\psi_{\hat{A}},\varphi_{\hat{C}}\otimes\varphi_{\hat{E}}\otimes\varphi_{\hat{A}})=\langle\psi_{\hat{C}}|\langle\psi_{\hat{E}}|\Pi_{1}^{\mathcal{C}}\left(|\bar{\psi}_{\hat{A}}\rangle\langle\bar{\varphi}_{\hat{A}}|\right)|\varphi_{\hat{E}}\rangle|\varphi_{\hat{C}}\rangle,$ (61)
 where $\Pi_{1}^{\mathcal{C}}$ is the CPTP map for box $\pi_{1}$ on cut $\mathcal{C}$ and the conjugate of $\bar{\psi}_{\hat{A}},\bar{\varphi}_{\hat{A}}$ is taken on basis $\{|i\rangle_{\hat{A}}\}$.
 We only care about the effect of $R_{\pi_{1}}^{\mathcal{C}}$ on space $\hat{A},\hat{C},\hat{E}$ because of the conditions (R1), (R2) required by the comb model.
-For $\mathcal{C}\in\mathcal{T}_{\pi}$, assume that the wire $B,C$ are connected with respect to the bases $\{|k_{\hat{C}}\rangle\},\{|l_{\hat{C}}\rangle\}$ of $\hat{C}$ and the bases $\{|k_{\hat{B}}\rangle\},\{|l_{\hat{B}}\rangle\}$ of $\hat{B}$, the wire $D,A$ are connected with respect to the orthonormal bases $\{|p_{\hat{D}}\rangle\}$, $\{|q_{\hat{D}}\rangle\}$ of $\hat{D}$ and the orthonormal bases $\{|p_{\hat{A}}\rangle\}$, $\{|q_{\hat{A}}\rangle\}$ of $\hat{A}$.
+For $\mathcal{C}\in\mathfrak{C}(\mathcal{T}_{(\pi_{1}\|\pi_{2}^{\prime})\theta})$, assume that the wire $B,C$ are connected with respect to the bases $\{|k_{\hat{C}}\rangle\},\{|l_{\hat{C}}\rangle\}$ of $\hat{C}$ and the bases $\{|k_{\hat{B}}\rangle\},\{|l_{\hat{B}}\rangle\}$ of $\hat{B}$.
+The wire $D,A$ are connected with respect to the orthonormal bases $\{|p_{\hat{D}}\rangle\}$, $\{|q_{\hat{D}}\rangle\}$ of $\hat{D}$ and the orthonormal bases $\{|p_{\hat{A}}\rangle\}$, $\{|q_{\hat{A}}\rangle\}$ of $\hat{A}$.
+These bases are predefined by the protocol $\pi$ itself.
 The CJ representation of $\pi_{1}\pi_{2}$ is
 $R_{\pi_{1}\pi_{2}}^{\mathcal{C}}(\psi_{\hat{E}}\otimes\psi_{\hat{F}},\varphi_{\hat{E}}\otimes\varphi_{\hat{F}})$ (62)
-$=\sum_{k,l,p,q}R_{\pi_{1}}^{\mathcal{C}}(k_{\hat{C}}\otimes\psi_{\hat{E}}\otimes\bar{p}_{A},l_{\hat{C}}\otimes\varphi_{\hat{E}}\otimes\bar{q}_{\hat{A}})R_{\pi_{2}}^{\mathcal{C}}(p_{\hat{D}}\otimes\psi_{\hat{F}}\otimes\bar{k}_{\hat{B}},q_{\hat{D}}\otimes\varphi_{\hat{F}}\otimes\bar{l}_{\hat{B}})$
+$=\sum_{k,l,p,q}R_{\pi_{1}}^{\mathcal{C}}(k_{\hat{C}}\otimes\psi_{\hat{E}}\otimes\bar{p}_{A},l_{\hat{C}}\otimes\varphi_{\hat{E}}\otimes\bar{q}_{\hat{A}})R_{\pi_{2}}^{\mathcal{C}}(p_{\hat{D}}\otimes\psi_{\hat{F}}\otimes\bar{k}_{\hat{B}},q_{\hat{D}}\otimes\varphi_{\hat{F}}\otimes\bar{l}_{\hat{B}}).$
 For $\mathcal{C}^{\prime}\in\mathfrak{C}(\mathcal{T}_{\theta_{1}})$, $\{|i\rangle_{\hat{C}^{\prime}}\}$ is a base of space $\mathcal{F}_{C^{\prime}}^{\chi_{\theta_{1}}(\mathcal{C}^{\prime})}$, and $\{|i\rangle_{\tilde{C}^{\prime\prime}}\}$ is a base of space $\mathcal{F}_{C^{\prime\prime}}^{f_{1}(\chi_{\theta_{1}}(\mathcal{C}^{\prime}))}$ such that $U_{f_{1}}|i\rangle_{\hat{C}}=|i\rangle_{\tilde{C}^{\prime\prime}}$  [^2] This is possible because $\mathcal{F}_{C^{\prime}}^{\chi_{\theta_{1}}(\mathcal{C}^{\prime})}\simeq\mathcal{F}_{C^{\prime\prime}}^{f_{1}(\chi_{\theta_{1}}(\mathcal{C}^{\prime}))}$..
@@
 where $\bar{\alpha}_{\tilde{C}^{\prime\prime}}\in\mathcal{F}_{C^{\prime\prime}}^{\mathcal{C}^{\prime}}$ and $\bar{\alpha}_{\hat{C}^{\prime}}\in\mathcal{F}_{C^{\prime}}^{\chi(\mathcal{C}^{\prime})}$ have the same coefficients in the bases of $\{|i\rangle_{\tilde{C}^{\prime\prime}}\}$ and $\{|i\rangle_{\hat{C}^{\prime}}\}$.
-Since the state on space $\hat{C}^{\prime\prime}$ is always a vacuum state, we omit the arguments $\psi_{\hat{C}^{\prime\prime}},\varphi_{\hat{C}^{\prime\prime}}$ of $R_{\theta_{1}^{\mathcal{C}^{\prime}}}$ and the second term of Equation (52) in the remaining part of the paper.
-For all $\mathcal{C}\in\bar{\mathfrak{C}}(\mathcal{T}_{\pi_{1}\theta_{1}})$, wire $C$ and $C^{\prime}$ are connected with respect the bases $\{|k_{\hat{C}^{\prime}}\rangle\}$ and $\{|l_{\hat{C}^{\prime}}\rangle\}$ such that $|k_{\hat{C}^{\prime}}\rangle\cong|k_{\hat{C}}\rangle,|l_{\hat{C}^{\prime}}\rangle\cong|l_{\hat{C}}\rangle$.
+Since the state in space $\hat{C}^{\prime\prime}$ is always a vacuum state, we omit the arguments $\psi_{\hat{C}^{\prime\prime}},\varphi_{\hat{C}^{\prime\prime}}$ of $R_{\theta_{1}^{\mathcal{C}^{\prime}}}$ and the second term of Equation 63 in the remaining part of the paper.
+For all $\mathcal{C}\in\bar{\mathfrak{C}}(\mathcal{T}_{\pi_{1}\theta_{1}})$, wire $C$ and $C^{\prime}$ are connected with respect to the bases $\{|k_{\hat{C}^{\prime}}\rangle\}$ and $\{|l_{\hat{C}^{\prime}}\rangle\}$.
+Then we have the isomorphisms $|k_{\hat{C}^{\prime}}\rangle\cong|k_{\hat{C}}\rangle,|l_{\hat{C}^{\prime}}\rangle\cong|l_{\hat{C}}\rangle$  [^3] Since any pair of bases $\{|k_{\hat{C}^{\prime}}\rangle\}$, $\{|k_{\hat{C}}\rangle\}$ defines a specific isomorphism between $\hat{C}^{\prime}$ and $\hat{C}$, we use the notation $|k_{\hat{C}^{\prime}}\rangle\cong|k_{\hat{C}}\rangle$ to denote this isomorphism..
 the CJ representation of the composition box $\bar{\pi}_{1}=\pi_{1}\theta_{1}$ is,
@@
 where $|\gamma_{\hat{C}}\rangle=\left(\sum_{k}|k_{\hat{C}}\rangle\langle k_{\tilde{C}^{\prime\prime}}|\right)|\gamma_{\tilde{C}^{\prime\prime}}\rangle$ and $|\eta_{\hat{C}}\rangle=\left(\sum_{l}|l_{\hat{C}}\rangle\langle l_{\tilde{C}^{\prime\prime}}|\right)|\eta_{\tilde{C}^{\prime\prime}}\rangle$.
-Equation (53) follows from Definition 18, Equation (54) follows from Equation (52) and Equation (55) follows from the sesquilinearity.
-Similarly, for all $\mathcal{C}\in\bar{\mathfrak{C}}(\mathcal{T}_{\pi_{2}^{\prime}\theta_{2}})$, wire $D$ and $D^{\prime}$ are connected with respect the bases $\{|p_{\tilde{D}^{\prime}}\rangle\}$ and $\{|q_{\tilde{D}^{\prime}}\rangle\}$ such that $|p_{\tilde{D}^{\prime}}\rangle\cong|p_{\tilde{D}}\rangle,|q_{\tilde{D}^{\prime}}\rangle\cong|q_{\tilde{D}}\rangle$.
+Equation 64 follows from Definition 18, Equation 65 follows from Equation 63, and Equation 66 follows from the sesquilinearity.
+Similarly, for all $\mathcal{C}\in\bar{\mathfrak{C}}(\mathcal{T}_{\pi_{2}^{\prime}\theta_{2}})$, wire $D$ and $D^{\prime}$ are connected with respect to bases $\{|p_{\tilde{D}^{\prime}}\rangle\}$ and $\{|q_{\tilde{D}^{\prime}}\rangle\}$.
 The CJ representation of the box $\bar{\pi}^{\prime}_{2}=\pi^{\prime}_{2}\theta_{2}$ is
@@
 where $|\lambda_{\tilde{D}}\rangle=\left(\sum_{p}|p_{\tilde{D}}\rangle\langle p_{\hat{D}^{\prime\prime}}|\right)|\lambda_{\hat{D}^{\prime\prime}}\rangle$ and $|\omega_{\tilde{D}}\rangle=\left(\sum_{q}|q_{\tilde{D}}\rangle\langle q_{\hat{D}^{\prime\prime}}|\right)|\omega_{\hat{D}^{\prime\prime}}\rangle$.
-Notice that, for all $\mathcal{C}\in\bar{\mathfrak{C}}(\mathcal{T}_{(\pi_{1}\|\pi^{\prime}_{2})})$, we have the relation chain of the base vectors
+Notice that, for all $\mathcal{C}\in\bar{\mathfrak{C}}(\mathcal{T}_{(\pi_{1}\|\pi^{\prime}_{2})})$, we have the isomorphism chain between the wires:
 $|k_{\tilde{B}}\rangle\cong|k_{\hat{B}}\rangle\cong|k_{\hat{C}}\rangle\cong|k_{\hat{C}^{\prime}}\rangle\cong|k_{\tilde{C}^{\prime\prime}}\rangle,$ (70)
@@
 $|q_{\hat{A}}\rangle\cong|q_{\tilde{A}}\rangle\cong|q_{\tilde{D}}\rangle\cong|q_{\tilde{D}^{\prime}}\rangle\cong|q_{\hat{D}^{\prime\prime}}\rangle.$
-The first column of $\cong$ follows naturally by simply substituting $t_{i}\in\mathcal{T}_{\pi}$ with $t^{\prime}_{i}\in\mathcal{T}_{\pi^{\prime}}$, The second column is defined by the connection of $\pi_{1},\pi_{2}$.
-The third column is defined by the connection inside $\bar{\pi}_{1}$ and $\bar{\pi}^{\prime}_{2}$.
-The fourth column follows from the effect of delay boxes $\theta_{1}$ and $\theta_{2}$.
-Then for $\mathcal{C}=\mathcal{T}_{(\pi_{1}\|\pi_{2}^{\prime})\theta}$, we choose the bases $\{|k\rangle_{\tilde{C}^{\prime\prime}}\},\{|l\rangle_{\tilde{C}^{\prime\prime}}\},\{|k\rangle_{\hat{B}}\},\{|l\rangle_{\hat{B}}\}$ for the connection of $C^{\prime\prime},B$ and the bases $\{|p_{\hat{D}^{\prime\prime}}\rangle\},\{|q_{\hat{D}^{\prime\prime}}\rangle\},\{|p_{\tilde{A}}\rangle\},\{|q_{\tilde{A}}\rangle\}$ for the connection of $D^{\prime\prime},A$.
+The isomorphisms $\cong$ in the first column are natural isomorphisms.
+The isomorphisms in the second column are predefined by the connection of $\pi_{1}\pi_{2}$.
+The isomorphisms in the third column can be chosen freely; that is, they are neither predefined nor natural.
+The isomorphisms in the fourth column follows from the definition of delay boxes $\theta_{1}$ and $\theta_{2}$.
+Therefore, the isomorphisms in the third column induce two pairs of bases for $\tilde{B},\tilde{C}^{\prime\prime}$ and two pairs of bases for $\hat{A},\hat{D}^{\prime\prime}$, which are precisely the bases we will use for $B,C^{\prime\prime}$ and $A,D^{\prime\prime}$ connections.
+Then, for $\mathcal{C}=\mathcal{T}_{(\pi_{1}\|\pi_{2}^{\prime})\theta}$, we choose the bases $\{|k\rangle_{\tilde{C}^{\prime\prime}}\},\{|l\rangle_{\tilde{C}^{\prime\prime}}\},\{|k\rangle_{\hat{B}}\},\{|l\rangle_{\hat{B}}\}$ for the connection of $C^{\prime\prime},B$ and the bases $\{|p_{\hat{D}^{\prime\prime}}\rangle\},\{|q_{\hat{D}^{\prime\prime}}\rangle\},\{|p_{\tilde{A}}\rangle\},\{|q_{\tilde{A}}\rangle\}$ for the connection of $D^{\prime\prime},A$.
 The CJ representation of the final box $\bar{\pi}_{1}\bar{\pi}^{\prime}_{2}$ is
@@
 where $\mathcal{C}_{1}^{\prime}=\mathcal{C}\cap\mathcal{T}_{\pi}$ and $\mathcal{C}_{2}^{\prime}=\mathcal{C}\cap\mathcal{T}_{\pi^{\prime}}$.
-Since $R_{\pi_{2}}$ and $R_{\pi^{\prime}_{2}}$ have the same effect on their respective spaces, there exist isomorphisms from $\hat{B},\hat{F},\hat{D}$ to $\tilde{B},\tilde{F},\tilde{D}$ such that
+$R_{\pi_{2}}$ and $R_{\pi^{\prime}_{2}}$ should have the same effect on their respective wire spaces.
+Rigorously speaking, for natural isomorphisms
+$|p_{\tilde{D}}\rangle\cong|p_{\hat{D}}\rangle,|q_{\tilde{D}}\rangle\cong|q_{\hat{D}}\rangle,$ (74)
+$|k_{\tilde{B}}\rangle\cong|k_{\hat{B}}\rangle,|l_{\tilde{B}}\rangle\cong|l_{\hat{B}}\rangle,$
+$|\psi_{\tilde{F}}\rangle\cong|\psi_{\hat{F}}\rangle,|\varphi_{\tilde{F}}\rangle\cong|\varphi_{\hat{F}}\rangle,$
+we have
 $R_{\pi^{\prime}_{2}}^{\mathcal{T}_{\pi^{\prime}}}(p_{\tilde{D}}\otimes\psi_{\tilde{F}}\otimes\bar{k}_{\tilde{B}},q_{\tilde{D}}\otimes\varphi_{\tilde{F}}\otimes\bar{l}_{\tilde{B}})=R_{\pi_{2}}^{\mathcal{T}_{\pi}}(p_{\hat{D}}\otimes\psi_{\hat{F}}\otimes\bar{k}_{\hat{B}},q_{\hat{D}}\otimes\varphi_{\hat{F}}\otimes\bar{l}_{\hat{B}}).$ (75)
-Note that the isomorphisms here are exactly the ones used in the first column of Equation (59). Then we have
-$R_{\bar{\pi}_{1}\bar{\pi}^{\prime}_{2}}^{\mathcal{T}_{(\pi_{1}\|\pi_{2})\theta}}(\psi_{\hat{E}}\otimes\psi_{\tilde{F}},\varphi_{\hat{E}}\otimes\varphi_{\tilde{F}})=R_{\pi_{1}\pi_{2}}^{\mathcal{T}_{\pi}}(\psi_{\hat{E}}\otimes\psi_{\hat{F}},\varphi_{\hat{E}}\otimes\varphi_{\hat{F}})$ (64)
+Substituting Equation 75 into Equation 73, we have
+$R_{\bar{\pi}_{1}\bar{\pi}^{\prime}_{2}}^{\mathcal{T}_{(\pi_{1}\|\pi_{2})\theta}}(\psi_{\hat{E}}\otimes\psi_{\tilde{F}},\varphi_{\hat{E}}\otimes\varphi_{\tilde{F}})=R_{\pi_{1}\pi_{2}}^{\mathcal{T}_{\pi}}(\psi_{\hat{E}}\otimes\psi_{\hat{F}},\varphi_{\hat{E}}\otimes\varphi_{\hat{F}}),$ (76)
+which concludes the proof.
 B.2 Proof of Theorem 10
+Theorem.
+Let $\pi$ be a $\mathrm{WCF}(z,\epsilon)$ protocol and $\eta$ be any protocol with only classical output at the outer interface.
+$\xi=\pi\|\eta$ is their parallel composition with the ordered set $\mathcal{T}$.
+Then $\pi$ is globally secure in the protocol if there exists a partition $\mathcal{T}_{1},\mathcal{T}_{2},\mathcal{T}_{3}$ of $\mathcal{T}$ such that
+  1. (C1)
+    $\forall t_{1}\in\mathcal{T}_{1},t_{2}\in\mathcal{T}_{2},t_{3}\in\mathcal{T}_{3}$, $t_{1}\prec t_{2}\prec t_{3}$;
+  2. (C2)
+    $\eta$ only accepts input and produces output on set $\mathcal{T}_{1}\cup\mathcal{T}_{3}$, while $\pi$ only accepts input and produces output on set $\mathcal{T}_{2}$.
+In this proof, we focus on the first condition in Definition 9, then the second condition follows similarly.
+Let $\alpha$ be the adversarial box connected to the parallel composition $\pi_{A}\|\eta_{A}$, as shown in Figure 9(a).
+For convenience, $c_{A},K_{A}$ in Definition 9 are relabeled as $E,K$, respectively.
+The idea of this proof is by contradiction, explained as follows.
+Our target is to bound the conditional probability $\Pr[E=1|K=k]$.
+Due to condition (C1) and (C2), the interaction of $\alpha$ divides into three phases (see Figure 9(b)).
+Observe that the only way that $K$ can influence $E$ is through the information $Q_{1}$ passed from the first phase $\alpha_{1}$ to the second phase $\alpha_{2}$.
+If the adversarial box $\alpha$ can bias $\Pr[E=1|K=k]$ beyond $\frac{1}{2}+\epsilon$, then the information on $Q_{1}$ conditioned on $K=k$, together with the box $\alpha_{2}$, becomes a successful attacking strategy for $\pi_{A}$ alone.
 We first simplify the adversarial box and prove this theorem by contradiction.
-Let $\alpha$ be a box connected with the parallel composition $\pi_{A}\|\eta_{A}$, as shown in Figure 8(a).
-For convenience, $c_{A},K_{A}$ are relabeled as $E,K$.
 Let $\mathcal{T}$ be the ordered set of $\pi_{A}\|\eta_{A}$.
@@
 $\mathcal{T}_{2}$$\mathcal{T}_{1}$$\mathcal{T}:$$\mathcal{T}_{3}$
-Figure 8: (a) The basic connection.
+Figure 9: (a) The basic connection of box $(\pi_{A}\|\eta_{A})\alpha$.
 (b) An equivalent form of the box.
-The wire $X,Y,V,W$ are split into $X_{1},X_{2},Y_{1},Y_{2},V_{1},V_{2},W_{1},W_{2}$.
-(c) Causal order of the box $(\pi_{A}\|\eta_{A})\alpha$.
+The wires $X,Y,V,W$ are split into $(X_{1},X_{2}),(Y_{1},Y_{2}),(V_{1},V_{2}),(W_{1},W_{2})$, respectively.
+(c) Causal order of box $(\pi_{A}\|\eta_{A})\alpha$.
+The sets $\mathcal{T}_{1},\mathcal{T}_{2},\mathcal{T}_{3}$ as a whole are strictly ordered.
 According to the condition (C1) and (C2), the wire $(V_{1},W_{1})$, $(B,D)$ and $(V_{2},W_{2})$ are effective only on $\mathcal{T}_{1}$, $\mathcal{T}_{2}$ and $\mathcal{T}_{3}$, respectively.
@@
 Since we are only interested in the final output, we set $\mathcal{C}=\mathcal{T}$.
-By [35, Lemma 9], the map $\Phi_{\alpha}^{\mathcal{T}}$ can be decomposed into three maps $\Phi_{\alpha_{1}}^{\mathcal{T}}:V_{1}\mapsto W_{1}Q_{1}$, $\Phi_{\alpha_{2}}^{\mathcal{T}}:BQ_{1}\mapsto DQ_{2}$ and $\Phi_{\alpha_{3}}^{\mathcal{T}}:V_{2}Q_{2}\mapsto W_{2}$ that satisfy
+By [36, Lemma 9], the map $\Phi_{\alpha}^{\mathcal{T}}$ can be decomposed into three maps $\Phi_{\alpha_{1}}^{\mathcal{T}}:V_{1}\mapsto W_{1}Q_{1}$, $\Phi_{\alpha_{2}}^{\mathcal{T}}:BQ_{1}\mapsto DQ_{2}$ and $\Phi_{\alpha_{3}}^{\mathcal{T}}:V_{2}Q_{2}\mapsto W_{2}$ that satisfy
 $\Phi_{\alpha}^{\mathcal{T}}=(\mathrm{id}_{W_{1}}\otimes\mathrm{id}_{D}\otimes\Phi_{\alpha_{3}}^{\mathcal{T}})\circ(\mathrm{id}_{W_{1}}\otimes\Phi_{\alpha_{2}}^{\mathcal{T}}\otimes\mathrm{id}_{V_{2}})\circ(\Phi_{\alpha_{1}}^{\mathcal{T}}\otimes\mathrm{id}_{B}\otimes\mathrm{id}_{V_{2}}).$ (77)
@@
 Similarly, the box $\eta_{A}$ is split into $\eta_{A1}$ and $\eta_{A2}$.
+The decomposition is illustrated in Figure 9(b).
 Let $\Phi_{\eta_{A}\alpha_{1}}^{\mathcal{T}}$ and $\Phi_{\pi_{A}\alpha_{2}}^{\mathcal{T}}$ be the map of connected boxes $\eta_{A}\alpha_{1}$ and $\pi_{A}\alpha_{2}$.
 Denote $\bar{Q}=\mathcal{F}_{Q}^{\mathcal{T}}$, then the output of $\Phi_{\eta_{A}\alpha_{1}}^{\mathcal{T}}$ is a classical-quantum state $\rho_{\bar{Q}K}=\sum_{k}P_{K}(k)|k\rangle\langle k|\otimes\rho_{\bar{Q}}^{k}$.
-$\Phi_{\pi_{A}\alpha_{2}}^{\mathcal{T}}$ is a map from $\mathfrak{T}(\bar{Q})$ to the classical bit $E$, which can be represented as a binary POVM measurement $\{M_{0},M_{1}\}$.
+$\Phi_{\pi_{A}\alpha_{2}}^{\mathcal{T}}$ is a map from $\mathfrak{T}(\bar{Q})$ to the classical bit $E$, which can be represented as a binary POVM $\{M_{0},M_{1}\}$.
 Then the conditional probability is $P_{E|K}(e|k)=\mathrm{tr}(M_{e}\rho_{\bar{Q}}^{k})$.
+Now we are ready to prove the theorem by contradiction.
 Assume $\exists k\in\mathbb{N},\Pr[E=1|K=k]\geq z+\epsilon$, then $\mathrm{tr}(M_{e}\rho_{\bar{Q}}^{k})\geq z+\epsilon$.
-Consider a box $\gamma$ that outputs state $\rho_{\bar{Q}}^{z}$ and has no input.
+Consider a box $\gamma$ that outputs state $\rho_{\bar{Q}}^{k}$ and has no input.
 The box $\pi_{A}\alpha_{2}\gamma$ will output $E=1$ with probability larger than $z+\epsilon$, which contradicts the stand-alone security.
 Therefore, the first condition in Definition 9 is satisfied.
-The second condition follows similarly.
 B.3 Proof of Corollary 11
-We denote the fractional part of $z$ as a binary string $b_{1}^{n}$.
-We denote the set $I=\{i\in\mathbb{N}\mid b_{i}=1\}$ as a string $a_{1}^{|I|}$ and the set $I^{\prime}=\{i\in\mathbb{N}\mid b_{i}=0\}$ as a string ${a^{\prime}}_{1}^{|I^{\prime}|}$ in ascending order.
+Corollary.
+Let $z\in[0,\frac{1}{2}]$. The unbalanced WCF protocol shown in Figure 7 is a $\mathrm{WCF}(z,\epsilon^{\prime})$ protocol with $\epsilon^{\prime}=2\epsilon+o(\epsilon)$.
+The intuition for the proof of Corollary 11 is as follows.
+The probability of one party ultimately winning is a continuous function of the probabilities of winning in each round.
+Thus, a small bias in each round necessarily leads to a small overall bias.
+Denote the fractional part of $z$ as a binary string $b_{1}^{n}$ (with finite precision), that is, $z=0.b_{1}b_{2}\dots b_{n}$.
+Define the sets $I\coloneqq\{i\in[n]\mid b_{i}=0\}$ and $I^{\prime}\coloneqq\{i\in[n]\mid b_{i}=1\}$.
+The elements of $I$ and $I^{\prime}$ form the sequences $(a_{1},\dots,a_{|I|})$ and $(a^{\prime}_{1},\dots,a^{\prime}_{|I^{\prime}|})$, respectively, arranged in ascending order.
 The $i$th WCF protocol $\pi_{i}$ as well as the other WCF protocols $\pi_{1}\|...\|\pi_{i-1}$ and $\pi_{i+1}\|...\|\pi_{n}$ output classical bits and follow strict temporal orders.
-Therefore, we obtain global security of $\pi_{i}$ according to Theorem 10
-$\Pr\left[C_{i}=0|C_{1}^{i-1}=c_{1}^{i-1}\right]$ $\leq\frac{1}{2}+\epsilon,$ (66)
-$\Pr\left[C_{i}=1|C_{1}^{i-1}=c_{1}^{i-1}\right]$ $\leq\frac{1}{2}+\epsilon.$ (67)
-We first argue that a cheating party needs to maximize the winning probability of each $\pi_{i}$ in order to maximize the final winning probability.
-At each step $\pi_{i}$, if $b_{i}=1$, Alice wins with certainty if $C_{i}=0$ while wins with a probability smaller than $1$ if $C_{i}=1$; if $b_{i}=0$, Alice wins with a non-zero probability if $C_{i}=0$ while loses with certainty if $C_{i}=1$.
-Therefore, a cheating Alice strives to win at each step $\pi_{i}$ in order to finally win with the highest probability.
-It is the same for Bob.
-In conclusion, considering global security in Inequality (67), Alice’s winning probability is no larger than the case when
-$\Pr\left[C_{i}=0|C_{1}^{i-1}=b_{1}^{i-1}\right]=\frac{1}{2}+\epsilon,$ (68)
-and Bob’s winning probability is no larger than the case when
-$\Pr\left[C_{i}=1|C_{1}^{i-1}=b_{1}^{i-1}\right]=\frac{1}{2}+\epsilon.$ (69)
+Therefore, for $i=2,3,\dots,n$ and any bit sequence $c_{1}^{i-1}$, we have global security of $\pi_{i}$ according to Theorem 10:
+$\Pr\left[C_{A,i}=1|C_{A,1}^{i-1}=c_{1}^{i-1}\right]$ $\leq\frac{1}{2}+\epsilon,$ (78)
+$\Pr\left[C_{B,i}=0|C_{B,1}^{i-1}=c_{1}^{i-1}\right]$ $\leq\frac{1}{2}+\epsilon,$ (79)
+where $C_{B(A),1}^{i}=C_{B(A),1}C_{B(A),2}\dots C_{B(A),i}$.
+We first argue that a cheating party needs to maximize the winning probability of each game $\pi_{i}$ in order to maximize the final winning probability.
+At each step $\pi_{i}$, if $b_{i}=0$, Bob wins the final game with certainty if $C_{A,i}=1$ and wins the final game with a probability smaller than $1$ if $C_{A,i}=0$; if $b_{i}=1$, Bob wins with non-zero probability if $C_{A,i}=0$ while loses with certainty if $C_{A,i}=1$.
+Therefore, a cheating Bob strives to win at each round of $\pi_{i}$ in order to finally win with the highest probability.
+The similar argument holds for cheating Alice.
+In conclusion, Bob’s winning probability is no larger than the case when
+$\forall i\in[n],\Pr\left[C_{A,i}=1|C_{A,1}^{i-1}=b_{1}^{i-1}\right]=\frac{1}{2}+\epsilon,$ (80)
+and Alice’s winning probability is no larger than the case when
+$\forall i\in[n],\Pr\left[C_{B,i}=0|C_{B,1}^{i-1}=b_{1}^{i-1}\right]=\frac{1}{2}+\epsilon.$ (81)
 We will use the above probabilities to compute the upper bound on Alice and Bob’s optimal winning probability.
-For a cheating Alice,
-Pr[Alice $\Pr[\text{Alice wins}]$ $=\sum_{i=1}^{|I|}\Pr\left[C_{1}^{a_{i}}=(b_{1}^{a_{i}-1},0)\right]$ (70)
-$=\sum_{i=1}^{|I|}\Pr\left[C_{a_{i}}=0|C_{1}^{a_{i}-1}=b_{1}^{a_{i}-1}\right]\prod_{j=1}^{a_{i}-1}\Pr\left[C_{j}=b_{j}|C_{1}^{j-1}=b_{1}^{j-1}\right]$ (71)
+According to the program of $\tau_{A}$ in Figure 7, Bob can win the final game only if the loop breaks at the $a$th round where $b_{a}=0$, or equivalently, $a\in I$.
+Then, his winning probability can be derived as
+Pr[Bob $\Pr[\text{Bob wins}]$ $=\sum_{a\in I}\Pr[\text{Loop breaks at $a$th round and $C_{A,a}=1$}]$ (82)
+$=\sum_{i=1}^{|I|}\Pr\left[C_{A,1}^{a_{i}}=(b_{1}^{a_{i}-1},1)\right]$ (83)
+$=\sum_{i=1}^{|I|}\Pr\left[C_{A,a_{i}}=1|C_{A,1}^{a_{i}-1}=b_{1}^{a_{i}-1}\right]\cdot\prod_{j=1}^{a_{i}-1}\Pr\left[C_{A,j}=b_{j}|C_{A,1}^{j-1}=b_{1}^{j-1}\right]$ (84)
+$=\sum_{i=1}^{|I|}\Pr\left[C_{A,a_{i}}=1|C_{A,1}^{a_{i}-1}=b_{1}^{a_{i}-1}\right]\cdot\prod_{j\in\{{a}_{k}\}_{k=1}^{i-1}}\Pr\left[C_{A,j}=0|C_{A,1}^{j-1}=b_{1}^{j-1}\right]\cdot$
+$\hphantom{=}\prod_{j^{\prime}\in[a_{i}-1]\backslash\{{a}_{k}\}_{k=1}^{i-1}}\Pr\left[C_{A,j^{\prime}}=0|C_{A,1}^{j^{\prime}-1}=b_{1}^{j^{\prime}-1}\right]$ (85)
 $\leq\sum_{i=1}^{|I|}\left(\frac{1}{2}+\epsilon\right)^{a_{i}-i+1}\left(\frac{1}{2}-\epsilon\right)^{i-1},$ (86)
-where $C_{1}^{0}=b_{1}^{0}$ is a certain event and Inequality (72) follows from Inequality (67). Similarly, for a cheating Bob, we obtain
-Pr[Bob $\Pr[\text{Bob wins}]$ $=\sum_{i=1}^{|I^{\prime}|}\Pr\left[C_{1}^{a^{\prime}_{i}}=(b_{1}^{a^{\prime}_{i}-1},1)\right]$ (73)
-$=\sum_{i=1}^{|I^{\prime}|}\Pr\left[C_{a^{\prime}_{i}}=1|C_{1}^{a^{\prime}_{i}-1}=b_{1}^{a^{\prime}_{i}-1}\right]\prod_{j=1}^{a^{\prime}_{i}-1}\Pr\left[C_{j}=b_{j}|C_{1}^{j-1}=b_{1}^{j-1}\right]$ (74)
+where $C_{A,1}^{0}=b_{1}^{0}$ is defined as a certain event and Inequality (86) follows from Inequality (78) and (80). Similarly, for cheating Alice, we obtain
+Pr[Alice $\Pr[\text{Alice wins}]$ $=\sum_{i=1}^{|I^{\prime}|}\Pr\left[C_{B,1}^{a^{\prime}_{i}}=(b_{1}^{a^{\prime}_{i}-1},0)\right]$ (87)
+$=\sum_{i=1}^{|I^{\prime}|}\Pr\left[C_{B,a^{\prime}_{i}}=0|C_{B,i}^{a^{\prime}_{i}-1}=b_{1}^{a^{\prime}_{i}-1}\right]\prod_{j=1}^{a^{\prime}_{i}-1}\Pr\left[C_{B,j}=b_{j}|C_{B,1}^{j-1}=b_{1}^{j-1}\right]$ (88)
 $\leq\sum_{i=1}^{|I^{\prime}|}\left(\frac{1}{2}+\epsilon\right)^{a^{\prime}_{i}-i+1}\left(\frac{1}{2}-\epsilon\right)^{i-1},$ (89)
-where we again apply Inequality (67) in (75). Then the bias is
-$\epsilon^{\prime}_{A}$ $=|\Pr[\text{Alice wins}]-z|\leq\sum_{i=1}^{|I|}\left(\frac{1}{2}+\epsilon\right)^{a_{i}-i+1}\left(\frac{1}{2}-\epsilon\right)^{i-1}-\frac{1}{2^{a_{i}}}\leq\sum_{i=1}^{|I|}\frac{2(a_{i}-2i+2)}{2^{a_{i}}}\epsilon+o(\epsilon).$ (76)
-We have omitted the absolute value because the winning probability of a cheating Alice is larger than that of an honest Alice, Pr[Alice $\Pr[\text{Alice wins}]\geq z$. Note that
+Then the bias achieved by Bob is
+$\epsilon^{\prime}_{B}$ $=|\Pr[\text{Bob wins}]-(1-z)|$ (90)
+$\leq\sum_{i=1}^{|I|}\left(\frac{1}{2}+\epsilon\right)^{a_{i}-i+1}\left(\frac{1}{2}-\epsilon\right)^{i-1}-\frac{1}{2^{a_{i}}}$ (91)
+$\leq\sum_{i=1}^{|I|}\frac{2(a_{i}-2i+2)}{2^{a_{i}}}\epsilon+o(\epsilon).$ (92)
+We have omitted the absolute value because the winning probability of a cheating Bob is always larger than that of an honest Bob, Pr[Bob $\Pr[\text{Bob wins}]\geq 1-z$. Note that
 $\sum_{n=i}^{\infty}\frac{n}{2^{n}}=\frac{i+1}{2^{i-1}}.$ (93)
 When $|I|=1$,
-$\epsilon^{\prime}_{A}\leq\frac{2a_{1}}{2^{a_{1}}}\epsilon+o(\epsilon)\leq\epsilon+o(\epsilon).$ (78)
+$\epsilon^{\prime}_{B}\leq\frac{2a_{1}}{2^{a_{1}}}\epsilon+o(\epsilon)\leq\epsilon+o(\epsilon).$ (94)
 When $|I|\geq 2$,
-$\epsilon^{\prime}_{A}\leq\frac{2a_{1}}{2^{a_{1}}}\epsilon+\frac{1}{2}\sum_{i=2}^{|I|}\frac{(a_{i}-2)}{2^{a_{i}-2}}\epsilon+o(\epsilon)\leq\frac{2a_{1}}{2^{a_{1}}}\epsilon+\frac{1}{2}\sum_{i=2}^{\infty}\frac{i-2}{2^{i-2}}\epsilon+o(\epsilon)\leq 2\epsilon+o(\epsilon).$ (79)
-In conclusion, for Alice,
+$\epsilon^{\prime}_{B}\leq\frac{2a_{1}}{2^{a_{1}}}\epsilon+\frac{1}{2}\sum_{i=2}^{|I|}\frac{(a_{i}-2)}{2^{a_{i}-2}}\epsilon+o(\epsilon)\leq\frac{2a_{1}}{2^{a_{1}}}\epsilon+\frac{1}{2}\sum_{i=2}^{\infty}\frac{i-2}{2^{i-2}}\epsilon+o(\epsilon)\leq 2\epsilon+o(\epsilon).$ (95)
+In conclusion, for cheating Bob,
+$\epsilon^{\prime}_{B}=|\Pr[\text{Bob wins}]-(1-z)|\leq 2\epsilon+o(\epsilon).$ (96)
+Similarly, for cheating Alice,
 $\epsilon^{\prime}_{A}=|\Pr[\text{Alice wins}]-z|\leq 2\epsilon+o(\epsilon).$ (97)
-Similarly for Bob,
-$\epsilon^{\prime}_{B}=|\Pr[\text{Bob wins}]-z|\leq 2\epsilon+o(\epsilon).$ (81)
 The final bias is
-$\epsilon^{\prime}=\max\{\epsilon^{\prime}_{A},\epsilon^{\prime}_{B}\}\leq 2\epsilon+o(\epsilon)$ (82)
+$\epsilon^{\prime}=\max\{\epsilon^{\prime}_{A},\epsilon^{\prime}_{B}\}\leq 2\epsilon+o(\epsilon).$ (98)